### v1

# Quality control in sublinear time: a case study via random graphs

Abstract
Many algorithms are designed to work well on average over inputs. When running such an algorithm on an arbitrary input, we must ask: Can we trust the algorithm on this input? We identify a new class of algorithmic problems addressing this, which we call “Quality Control Problems.” These problems are specified by a (positive, real-valued) “quality function” $\rho$ and a distribution $\mathcal{D}$ such that, with high probability, a sample drawn from $\mathcal{D}$ is “high quality,” meaning its $\rho$-value is near $1$. The goal is to accept inputs $x\sim\mathcal{D}$ and reject potentially adversarially generated inputs $x$ with $\rho(x)$ far from $1$. The objective of quality control is thus weaker than either component problem: testing for $\rho(x)\approx 1$” or testing if $x\sim\mathcal{D}$, and offers the possibility of more efficient algorithms.
[…]
1 Introduction
In this work, we formalize a new class of problems — which we call “quality control” problems — that aim to bridge the gap between algorithms designed for random inputs and adversarially generated examples. We then study a specific class of quality control problems for some basic graph-theoretic properties on random graphs that show the (sublinear) algorithmic possibilities in this space. We start by giving an illustrative example for this class of problems.
Suppose you are in a bookstore trying to find a book to your taste. We assume that given a book, if you read it fully, you would know if you like it or not. However, you do not have the time to read the whole book in the store (not to mention that the storekeepers may not be too happy either if you read books fully and then return them to the shelves claiming you don’t like them). You do have time to scan the book before you determine whether to buy it or not. You would not like to have a false positive (buying the book if it is not to your taste). You might not mind a false negative, though: if you reject this book, there are plenty of other books aimed at people like you. So, as long as the scanning procedure typically accepts books aimed at people like you, you should be happy — you’ll find something to buy and read. Our goal is to take advantage of the abundance of books satisfying our taste to make fast asymmetric decisions. We refer to this class of problems as quality control problems, which we formalize in the next section.
Of course, the example above is just for illustration purposes (and we do not give an efficient algorithm for filtering books). The primary motivation for our work is that there are many settings where it is easier — either complexity-theoretically or analytically — to design algorithms for instances drawn according to some distribution $\mathcal{D}$. Such algorithms notoriously perform poorly on adversarially generated examples. For example, while finding a satisfying assignment for a $k$-SAT formula is NP-complete [Coo71], there exist polynomial-time algorithms for finding a satisfying assignment with high probability over uniformly random $k$-SAT instances (in some parameter regimes) [CR92]. Other examples appear in Sum-of-Squares algorithms for planted cliques (e.g. [MPW15]) and throughout learning theory. Counts of motifs (i.e., patterns) in random graphs also display this property: for example, $\binom{n}{3}p^{3}$ is an estimate for the number of triangles in an Erdős–Rényi graph with parameters $n,p$ with high probability, but is not a good estimate for many specific graphs.
Will an algorithm that performs well on average-case instances work well on a specific instance? Quality control problems aim to address this. The goal is to construct an algorithm for certifying the safety of using an algorithm defined to work well over distribution $\mathcal{D}$ on a specific input, based on the “quality” (a positive, real-valued function $\rho$) of the input. It is impossible to test if a single input was drawn from $\mathcal{D}$. For quality control, one option is to estimate $\rho$ on the input; does the weakened objective allow for more efficient algorithms than this? In the context of our illustrative example of choosing books, quality control corresponds to our scanning procedure. We were fine rejecting a book we might have liked, as long as we don’t reject such books too often. On the other hand, we want to trust that we will not choose books we will not enjoy, since choosing a bad book is costly.
1.1 Quality Control problems
A generic quality control problem is given by a distribution $\mathcal{D}$ supported, say, on $\{0,1\}^{n}$, and a function $\rho:\{0,1\}^{n}\to\mathbb{R}_{\geq 0}$ that takes a value in $1\pm\varepsilon$ with high probability on the distribution $\mathcal{D}$ and some $\epsilon>0$. Our goal is to design a quality control algorithm $A$ that takes as input $x\in\{0,1\}^{n}$ and outputs an accept/reject verdict in time $t(n)$. The completeness property that we would like is that $A(x)$ accepts with high probability when $x\sim D$. The soundness condition we would like is that, for all $x$ such that $|\rho(x)-1|>\epsilon$, $A(x)$ rejects with high probability. We define quality control algorithms formally below.
Definition 1.0 (Quality control algorithms).
For function $\rho:\mathcal{X}\to\mathbb{R}_{\geq 0}$, small constant $\varepsilon>0$, and a distribution $\mathcal{D}:\mathcal{X}\to\mathbb{R}_{\geq 0}$ such that $\mathbb{P}_{x\sim\mathcal{D}}\left[\left|\rho(x)-1\right|\leq\varepsilon\right]\geq 9/10$, a $(\mathcal{D},\rho)$-quality control algorithm $A$ must satisfy the following:
  1. 1.
    $\mathbb{P}_{x\sim\mathcal{D};R}\left[A(x)=\textsc{Accept}\right]\geq 2/3$, where the probability is taken over distribution $\mathcal{D}$ and the randomness $R$ used by algorithm $A$.
  2. 2.
    For all $x$ satisfying $\left|\rho(x)-1\right|>\varepsilon$, $\mathbb{P}_{R}\left[A(x)=\textsc{Reject}\right]\geq 2/3$.
Let us first compare the related problems of testing if an input $x$ was drawn from $\mathcal{D}$ and approximating $\rho(x)$. First, testing if $x\sim\mathcal{D}$ is impossible, since we are only receiving one input from the distribution. Second, while an approximation algorithm for $\rho(x)$ can indeed by made into a quality control algorithm, we want to construct algorithms that are more efficient than those for approximating/certifying $\rho$ on worst-case inputs.
In the context of sublinear algorithms, we assume that the algorithm $A$ is given oracle access to $x$ (so the computation may be denoted $A^{x}$) and restrict the algorithm to make $q(n)$ queries to $x$. Thus a quality control problem is specified by a pair $(\mathcal{D},\rho)$ (we assume $\epsilon$ is some fixed tiny constant), and the performance of the algorithm is captured by the query complexity $q=q(n)$ and runtime $t=t(n)$.
To see how this captures our motivating setting, a book is a string $x\in\{0,1\}^{n}$. Our taste is captured by the function $\rho$ with the ideal book having $\rho(x)=1$, and our preference reduces as $\rho(x)$ gets further away from 1. The distribution $\mathcal{D}$ captures books aimed to satisfy our taste. The soundness condition ensures that we won’t buy a book that is not to our taste (with high probability). The completeness condition ensures we will find a book to our taste if a decent fraction of the books in the store are drawn from distribution $\mathcal{D}$. (So the bookstore does not have to exclusively carry books to our taste; it simply needs to carry a decent fraction of such books.)
Quality control algorithms, in our opinion, give the right set of constraints that, on the one hand, allow the algorithms to take advantage of the distribution $\mathcal{D}$ to gain algorithmic efficiency, while, on the other hand, do not fail catastrophically on adversarially generated examples. In particular, one significant concern with algorithms designed for specific distributions is that they typically assume vast amounts of independence among features. Such independence is notoriously hard to test, let alone being impossible to verify for a single instance. Quality control algorithms give us a path around such concerns by focusing on the validity of the conclusion rather than the assumptions used to design the algorithm.
Quality control algorithms do already exist in the literature, but to our knowledge, this is the first time the problem class has been formally introduced. In Section 1.3, we describe some previous instances of quality control problems that have been studied in the literature and also explain the relationship to the field of average-case complexity. In this paper, we show the viability of sublinear-time quality control algorithms for basic classes of graph-theoretic parameters relative to basic classes of random graphs, notably Erdős–Rényi graphs $\mathcal{G}_{n,p}$ — where every pair of vertices has an edge between them with probability $p$ independently of all other edges. (As we explain later, these algorithms work significantly faster than proven lower bounds for estimating the same parameters in the worst-case setting, while satisfying our soundness requirement.) We elaborate on our problems next.
1.2 Counting motifs in random graphs: problems and our results
Our testbed for quality control algorithms is problems that estimate basic counts of constant-sized subgraphs (i.e., motifs) in the Erdős–Rényi model of random graphs $\mathcal{G}_{n,p}$ – in which a graph on $n$ vertices is generated by including each possible edge independently with probability $p$. Most of our results are best captured by setting $\epsilon$ to some fixed constant (say $0.1$), fixing $\rho$, and then letting $p\to 0$, and then letting $n\to\infty$ much faster than, say, polynomially in $1/p$. This captures “sparse” and “dense” settings of $\mathcal{G}_{n,p}$. We aim for query and time complexities independent of $n$ that grow as slowly in $1/p$ as possible.
Motif counting is a natural case study for random graphs since motif statistics give structural information about the graph. In the dense setting, where $p\in(0,1)$, motif counts matching that expected from $\mathcal{G}_{n,p}$ can imply quasirandom properties [CGW89, SY10] or convergence of a graph sequence to a certain limiting object [LS06]. Quasirandomness is a graph property with many applications to theoretical computer science, notably due to its relation to graph expansion. Second, the counts of motifs in graphs parallels the role of moments of a distribution in various algorithmic contexts – e.g. [BCL11] – providing useful information about the underlying graph, such as parameters of network models that fit one’s data. Third, sublinear approximate motif counting is well-studied in the setting of arbitrary/general graphs (as we will describe in Section 1.3), and therefore, the superpolynomial improvement to the complexity in our setting can be readily compared.
k-clique counting
We consider the quality parameter corresponding to counting the number of $k$-cliques in a random graph $G\sim\mathcal{G}_{n,p}$. This problem is the technical core of this paper where we show that quality control algorithms offer significant benefits over worst-case algorithms. In this setting, we prove tight bounds on the query and runtime complexity of quality control. Let $\mu_{k}=\mu_{k}(n,p)$ denote the expected number of $k$-cliques in $\mathcal{G}_{n,p}$. For every $p>0$ and $k\in\mathbb{N}$, if $n$ is sufficiently large as a function of $p$ and $k$, then it is well-known that the number of $k$-cliques for $G\sim\mathcal{G}_{n,p}$ concentrates around $(1\pm o(1))\mu_{k}$. We let our parameter $\rho_{k}(G)$ be the number of $k$-cliques in $G$ divided by $\mu_{k}$. Since $\rho_{k}(G)$ for $G\sim\mathcal{G}_{n,p}$ concentrates around $1$, this leads to the question: what is the complexity of a sublinear time algorithm solving $(\mathcal{G}_{n,p},\rho_{k})$-quality control? We consider this question in the setting of sublinear graph access: the algorithm has access to the graph $G$ in the form of adjacency matrix queries. That is, the algorithm queries a pair $u,v\in G$, and receives the response $1$ if $(u,v)$ is an edge in $G$ and 0 otherwise.
It is trivial to produce a $(1\pm\varepsilon)$-approximation of the count of $k$-cliques that succeeds with probability $2/3$ over $G\sim\mathcal{G}_{n,p}$ (for $n\geq p^{-ck}$ for some constant $c$) by simply outputting $\mu_{k}$, but this does not verify whether $\mu_{k}$ is indeed a $(1\pm\varepsilon)$-approximation of the number of $k$-cliques on our specific instance. On the other hand, it is too costly to compute/certify the number of $k$-cliques for every worst-case graph; we are fine rejecting graphs such that $\mu_{k}$ is not a good estimate. Algorithms for testing whether a worst-case graph has around $\mu_{k}$ $k$-cliques require roughly $\Theta\left(1/p^{\binom{k}{2}}\right)$ queries [^1] This follows from subroutines in [ERS20b] that have query complexity and runtime of roughly $n/(\mu_{k})^{1/k}$ $m^{k/2}/\mu_{k}$., with even stronger query access to the underlying graph [ERS20b]. Assuming that the average degree and arboricity of the input graph correspond to $\mathcal{G}_{n,p}$ – towards the goal of applying [ERS20a] – yields a $\Theta\left(1/p^{\binom{k}{2}-k+1}\right)$ worst-case query complexity (omitting $\text{polylog}(n)$ factors), in a stronger query access model. The exponent on $1/p$ still depends quadratically on $k$.
We circumvent the $\Theta(k^{2})$ bound on the exponent of $1/p$ that arises from sublinear algorithms for estimating $k$-cliques in worst-case graphs. Our result in this setting is an asymptotically tight bound of $\Theta(k)$ on the exponent of $1/p$ for the quality control problem. Specifically we show:
Theorem 1.1.
There exist constants $c_{1},c_{2},c_{3}$ such that for every constant $k$, the quality control problem $(\mathcal{G}_{n,p},\rho_{k})$ corresponding to estimating the number of $k$-cliques in a graph $G$ supposedly drawn according to $\mathcal{G}_{n,p}$ is solvable in $O(p^{-c_{1}k})$ queries and time given adjacency matrix access to $G$, provided $n\geq p^{-c_{2}k}$. Furthermore no algorithm solves this problem with $o(p^{-c_{2}k})$ queries.
This theorem is formalized and proved as Theorem 3.1 in Section 3. To illustrate this result, we note that there exist constants $k_{0},c_{0}$ such that for all $k\geq k_{0}$ and $c\geq c_{0}$, when $n\approx p^{-ck}$, our algorithm from Theorem 1.1 uses $o(n)$ queries and time, while the algorithm from [ERS20a] requires $\Omega(n^{2})$ queries and time.
One ingredient in our proof of Theorem 1.1 is a non-trivial polynomial time quality control algorithm for estimating the number of $k$-cliques in $\mathcal{G}_{n,p}$ in the standard (non-sublinear time) setting.
Theorem 1.2.
There exist constants $c_{4}$ and $c_{5}$ such that for every constant $k$, the quality control problem $(\mathcal{G}_{n,p},\rho_{k})$ corresponding to estimating the number of $k$-cliques in a graph $G$ supposedly drawn according to $\mathcal{G}_{n,p}$ can be solved in $O(n^{c_{4}})$ time provided $n\geq p^{-c_{5}k}$.
Note that the requirement on $n$ is roughly comparable to the requirement that the number of $k$-cliques in a random graph drawn from $\mathcal{G}_{n,p}$ is concentrated. Our result shows that, in this setting, one does not need a brute-force $n^{k}$ time algorithm to estimate the number of $k$-cliques — one can get a fixed polynomial-time algorithm to do so by taking advantage of the allowance to reject graphs that would arise with low probability in $\mathcal{G}_{n,p}$.
Special case: Triangle Counting
We consider the simpler parameter of quality control of the count of triangles compared to $\mathcal{G}_{n,p}$. We prove tight bounds on the query and runtime complexity of quality control. Let $\mu_{\Delta}$ be the expected number of triangles in $\mathcal{G}_{n,p}$, $C_{\Delta}(G)$ be the number of triangles in $G$, and $\rho_{\Delta}(G)=C_{\Delta}(G)/\mu_{\Delta}$.
We consider $(\mathcal{G}_{n,p},\rho_{\Delta})$-quality control in the sublinear graph access setting where the algorithm now has query access to the graph $G$ in the form of adjacency matrix, adjacency list, and degree queries. It is well-known that testing if the number of triangles in a worst-case graph is approximately $\mu_{\Delta}$ uses $\Theta(p^{-3})$ queries (up to polylogarithmic factors) [ELRS17]. We show that the quality control problem offers a quadratic speedup over the worst-case setting, and this is tight.
Theorem 1.3.
The quality control problem $(\mathcal{G}_{n,p},\rho_{\Delta})$ corresponding to estimating the number of triangles in a graph $G$ supposedly drawn according to $\mathcal{G}_{n,p}$ is solvable in $\widetilde{O}(1/p)$ queries and time given adjacency list, adjacency matrix, and degree query access to $G$, provided $n\geq c_{6}\cdot p^{-1}$ for some constant $c_{6}$. Furthermore no algorithm solves this problem with $o(1/p)$ queries.
This theorem is formalized and proved as Theorem A.1 in Appendix A.
General Motif Counting
Our results and techniques regarding quality control of the count of $k$-cliques also extend to counting the number of copies of any fixed graph $H$ in some larger graph $G$. Specifically, let $C_{H}(G)$ denote the number of labeled induced copies of the graph $H$ in a graph $G$ (see Section 2.1 for a formal definition), and let $\mu_{H}:=\mathbb{E}_{G\sim\mathcal{G}_{n,p}}[C_{H}(G)]$ denote the expected number of copies of $H$ in a graph drawn from $\mathcal{G}_{n,p}$. Finally let $\rho_{H}(G):=C_{H}(G)/\mu_{H}$. We consider the quality control problem $(\mathcal{G}_{n,p},\rho_{H})$. Our extension of Theorem 1.1 identifies the exact exponent of $1/p$ in the query/time complexity of quality control for every graph $H$ (up to a fixed universal constant factor). The parameter we identify is the maximum degree of any vertex in $H$, denoted $\Delta(H)$. Specifically we show:
Theorem 1.4.
There exist constants $c_{7},c_{8},c_{9}$ such that for every graph $H$ on $k$ vertices, the quality control problem $(\mathcal{G}_{n,p},\rho_{H})$ corresponding to estimating the number of copies of $H$ in a graph $G$ supposedly drawn according to $\mathcal{G}_{n,p}$ is solvable in $O(p^{-c_{7}\Delta(H)})$ queries and running time given adjacency matrix access to $G$, provided $n\geq p^{-c_{8}k}$. Furthermore no algorithm solves this problem with $o(p^{-c_{9}\Delta(H)})$ queries.
This theorem is formalized and proved as Theorem 4.1 in Section 4.
We note that $\Delta(H)$ is not the most natural candidate parameter associated with counting the number of copies of $H$ in a graph. In particular, concentration results rely on the arboricity of a graph, which is usually smaller. $\Delta(H)$ is comparable in regular graphs, but then is much larger in a star.
To give some examples of motifs, this result implies that the complexity of quality control of cycles in $\mathcal{G}_{n,p}$ is $1/p^{\Theta(1)}$. This is much smaller than the $1/p^{O(k)}$ bound from sublinear approximation algorithms [AKK19] for large enough $n$ with respect to $p,k$. Next, the complexity of quality control for the star graph on $k$ vertices is $1/p^{\Theta(k)}$. This is incomparable to known approximation results [ABG+16], which give query/time complexity $\widetilde{\Theta}(n^{1-1/k})$.
Our results and techniques for quality control of motif counts more generally apply to quality control of “Sum-NC0 graph parameters” compared to $\mathcal{G}_{n,p}$; see Appendix B. This provides further justification that quality control of motif counts is a natural class of problems to study.
Beyond Erdős–Rényi Graphs
We extend the study of quality control of motif counts to a broad family of distributions over graphs that have concentration of the counts of small motifs. To illustrate that our results and techniques apply to other natural random graph distributions, we prove the following.
Theorem 1.5.
Consider $(\mathcal{D}_{n},\rho_{k})$-quality control, for $\rho_{k}(G)=C_{k}(G)/\mathbb{E}_{G^{\prime}\sim\mathcal{D}_{n}}\left[C_{k}(G^{\prime})\right]$. Define parameters $n,p$ where $n\geq p^{-c_{10}k}$ for some constant $c_{10}$. For the following settings of $\mathcal{D}_{n}$, the query complexity of $(\mathcal{D}_{n},\rho_{k})$-quality control is $1/p^{O(k)}$:
  1. 1.
    $\mathcal{D}_{n}$ is a Stochastic Block Model (as defined, e.g. in [HLL83]) with $n$ vertices and minimum in-community/ between-community probability at least $p$.
  2. 2.
    $\mathcal{D}_{n}$ is the uniform distribution over $d$-regular graphs on $n$ vertices, where $d=np$.
This theorem is formalized and proved as Section 5 in Section 5. It follows from a more general result we prove — Theorem 5.1 — which applies broadly to other distributions over random graphs that have concentrated counts of small motifs. Section 5 defines a notion called the density increment of a graph, which measures the ratio between the frequency of a motif and the frequency of its subgraphs on one fewer vertex. The inverse of the density increment asymptotically upper-bounds the query complexity of quality control of motif counting for distributions over random graphs with concentration properties.
1.3 Comparison to related works
In this section, we compare our models and results to some of the past literature.
Average Case Complexity
In computational complexity, the field of average case complexity (see, for instance, the work of Levin [Lev86] or surveys by Goldreich [Gol11] or Bogdanov and Trevisan [BT06]) is very close to our notions. Here, again, a problem may be described by a pair $(\mathcal{D},\rho)$ with $\rho$ describing some function that needs to be computed on distribution $\mathcal{D}$. Here one can consider two notions of what it means for an algorithm $A$ to compute a function $\rho$ on distribution $\mathcal{D}$: A strong definition may require that $A$ is always correct, but the distribution of the runtime is nice (e.g., has low-expected value, or is mostly fast) when inputs are drawn from $\mathcal{D}$. A weak requirement may be that $A$ is always fast, but only correct on most inputs drawn from $\mathcal{D}$. When translated to our setting (of $\rho$ being a real-valued parameter that we want to test for closeness to $1$), a quality control algorithm for $(\mathcal{D},\rho)$ is implied by a strong average case algorithm; and implies a weak average case algorithm. In Appendix C, we prove that a quality control algorithm for $(\mathcal{D},\rho)$ also implies a strong average case algorithm with the same runtime, under stronger assumptions about the concentration of $\mathcal{D}$ around its expected value.
Given this proximity to the notions of average case complexity, one may anticipate seeing instances of quality control problems arising naturally in the field. Indeed, one notable class of such problems is “refutations of random 3SAT” problems. Here a seminal definition of Feige [Fei02] requires that a refutation algorithm for 3SAT problem on distribution $\mathcal{D}$ should only accept unsatisfiable 3SAT inputs, while accepting most inputs drawn from $\mathcal{D}$. This corresponds exactly to a quality control problem with the parameter $\rho_{\mathrm{SAT}}(\phi)$ being $1$ if a formula $\phi$ is unsatisfiable and 0 otherwise. Additionally, the problem of certifying properties of a random object has appeared in various contexts throughout theoretical computer science. Notably, natural proofs [RR97] examine such a problem (within the context of ruling out avenues towards super-polynomial circuit lower bounds). Additionally, as noted by [Tre06], certifying properties of random instances is a challenge and a crucial step in explicit constructions from random objects. The problem of certifying properties of random objects in these various settings can be framed as a quality control problem.
Another class of problems that have been studied in the context of average case complexity are the “planted solution problems” — notably planted SAT or planted clique problems. Broadly, in this class of problems, there is a natural distribution $\mathcal{D}_{0}$ and then a planted distribution $\mathcal{D}_{1}$ that is far in total variation distance from $\mathcal{D}_{0}$ but has many locally similar features — and the challenge is to distinguish the two classes. This problem has been previously studied when the algorithm has access to the entire graph (e.g., [AKS98]), as well as in the sublinear access setting with query access to the adjacency matrix of the graph [RS20, HMP24]. If one were to consider the quality control problem $(\mathcal{D}_{1},\rho)$ where $\rho(x)=1$ if $x$ has a planted solution, then a solution to the quality control problem also solves the associated planted solution problem in both the sublinear and global frameworks. Indeed, our results imply that there is an algorithm with query complexity and runtime $p^{-O(\Delta(H))}$ that distinguishes between $\mathcal{G}_{n,p}$ and $\mathcal{G}_{n,p}$ with an adversarially placed structure contributing at least $\varepsilon\mu_{H}$ labeled induced copies of a constant-sized subgraph $H$, for $\mu_{H}:=\mathbb{E}_{G\sim\mathcal{G}_{n,p}}[C_{H}(G)]$.
One key difference between our motivation and those in average case complexity is the typical focus: Our primary focus is on sublinear algorithms, whereas traditional work in average case complexity focused on polynomial time algorithms. (In the case of clique counting, see, e.g., [BBB19] for counting cliques over $\mathcal{G}_{n,p}$ and [GR18] for worst-case to average-case reductions.)
Sublinear Algorithms (for Motif Counting)
By now, there is also a vast body of work on sublinear algorithms (see, for instance, surveys by Rubinfeld and Shapira  [RS11] or Ron [Ron09]). Approximately counting and sampling motifs in “worst-case” graphs with sublinear algorithms have been thoroughly studied – e.g. [FGP20, BER21, ELRR25, ABG+16] – and optimal bounds have been proven in various settings [AKK19, ELRS17, ERS20a, ERR22].
While the quality control framework of course builds on notions already developed in those fields, we stress two main differences in our setting: Sublinear time algorithms are required to work for worst-case inputs, while we intend to take advantage of the distribution of inputs — leading to a weaker completeness requirement in our case. (Indeed this is best evidenced in the $k$-clique counting results — where it is known that checking whether the number of $k$-cliques is roughly $\binom{n}{k}p^{\binom{k}{2}}$ via approximate counting requires $p^{-\Omega(k^{2})}$ queries/time in the sublinear setting [ERS20a], while our Theorem 1.1 achieves a query complexity and running time of $p^{-O(k)}$.) The soundness requirements in the two models are also not identical: In sublinear algorithms it is common to place a distance measure on inputs and allow algorithms to produce outputs that would be consistent with outputs on a small perturbation of the input. Our soundness requirement is different in that it is built into the (real-valued) parameter we aim to compute and effectively we are asking for a close approximation of this parameter.
Of course, despite the differences, many results developed in the literature on sublinear algorithms turn out to be useful for us (and sometimes they are even designed for our notion of soundness). In particular our triangle-counting result (Theorem A.1 in Appendix A) is obtained by combining known results in sublinear algorithms, specifically for estimating arboricity [EMR22] and for counting triangles in bounded arboricity graphs [ERS20a]. This ability to combine algorithms is a highlight and feature of our quality control definitions: An algorithm for $(\mathcal{D},\rho)$-quality control can run an algorithm for some task $(\mathcal{D},\tau)$ (same distribution but unrelated parameter) as a filter without hurting either completeness or soundness! In the triangle counting setting, we do exactly this to filter out graphs whose arboricity is not roughly $pn$, and this allows us to take advantage of fast sublinear algorithms for bounded arboricity graphs.
Distribution Testing and Testable Learning
The concern of “Does the input really come from distribution $\mathcal{D}$?” has been approached in several other ways in the literature. One natural class of problems motivated by this concern are the distribution testing problems (see, e.g., [Can22]) where the goal is to determine, given access to samples from a distribution $\mathcal{D}$ if $\mathcal{D}\in\mathcal{D}_{0}$ or if $\mathcal{D}$ is far from $\mathcal{D}_{0}$. However, these results require many samples from $\mathcal{D}$ to achieve their goals. Indeed, if we had multiple samples, testing if the distribution over graphs we are receiving is close to the Erdős–Rényi graph distribution would require at least exponentially many samples, in the size of the graph. Each sample itself is a very large graph, which could require significant computation to read, so this is computationally intractable for several reasons. Our goal diverges from distribution testing because we aim to do something with a sublinear access to single sample.
The idea that one can sidestep the large sample lower bounds required to test distributions, and directly get to validating the conclusion, arose recently in work of Rubinfeld and Vasilyan [RV23], in the notion they call testable learning. They consider the setting of a learning algorithm that works with a distribution $\mathcal{D}$ with the assumption that $\mathcal{D}\in\mathcal{D}_{0}$. They show that it is possible to test that the output of the learner can be trusted with lower sample complexity than required to test the distribution alone, such that if the learned hypothesis is incorrect then the tester will reject, while if $\mathcal{D}\in\mathcal{D}_{0}$ the tester will accept. While these results still work in the setting of multiple samples from $\mathcal{D}$, our model is inspired by their model — and the emphasis on testing the validity of the computation rather than the correctness of a distributional assumption.
Quasirandomness
One line of research in the combinatorics literature that aims to give “deterministic” definitions of randomness is the study of quasirandom graphs (see, e.g., the survey of Krivelevich and Sudakov [KS06]). The study of graph pseudorandomness — beginning with Thomason [Tho87] and first formally defined as “quasirandomness” by Chung, Graham, and Wilson [CGW89] — identifies equivalent deterministic structural properties possessed by graphs from $\mathcal{G}_{n,p}$ with high probability. While often studied for constant $p$ and $\mathcal{G}_{n,p}$, some work has focused on sparse quasirandom graphs ($p=o(1)$) [CG02, Zha23] and quasirandom graphs with other (non-$\mathcal{G}_{n,p}$) degree sequences [GG08].
One way to view these works in our language is that they identify specific quality parameters $\rho$ that concentrate around $1$ for random graphs drawn from some distribution $\mathcal{D}$. Many of the results can then be described as implications showing, e.g., graphs with parameter $\rho\approx 1$ also have some parameter $\rho^{\prime}\approx 1$ — thus giving in our language reductions among quality control problems. In some cases, they also give efficient (polynomial-time) algorithms to compute some non-trivial parameters. Thus, these results form a useful toolkit for the design of quality control algorithms, and indeed, they play a significant role in our results.
Applications
Beyond applications to problems in theoretical computer science — such as average-case complexity (see Appendix C) and planted solution problems (see the “Average Case Complexity” section above) — quality control has potential applications to applied algorithmic settings. The counts, locations, and presence of different motifs can be used to analyze and classify the structure of real-world networks, including technological and social networks [DT14, DGP17, CGGL17, LHW+18, DGP19], biological networks [SOMMA02, KLWP11, WLY14, PM20], brain networks [SK04, BNCL17], and ecological networks [BKRS15, SCB+19, SSAR19], among many others. On the theoretical side of the analysis of the structure of real-world networks, for tractability, simplified models of random processes that generate graphs are often considered; see, e.g. [VDH24], which discusses a variety of random graph models of real-world networks. Along with Erdős–Rényi graphs, different real-world networks can be modeled by stochastic block models, the Chung-Lu graph model, and the preferential attachment model, just to name a few. However, on the practical side of work on real-world networks, the network comes from data and it may not be clear if the theoretical assumptions align with the real-world scenario. Are theoretical algorithms safe to use on the real-world data? Quality control provides an intermediate step for verifying whether a theoretical algorithm is truly safe to use on an input. This parallels the framework of testable learning defined by Rubinfeld and Vasilyan [RV23], which focuses on testing the assumptions of learning algorithms.
1.4 Techniques and proof overviews
Our technical results are obtained by a combination of tools from probability theory (notably “concentration of sampled value of functions with high-probability bounded differences” [Com24]), tools from quasirandomness (notably the notion of jumbledness of graphs [Tho87] and its implications for sparse quasirandomness and efficient motif counting [CFZ14]) and finally by the following simple but powerful insight:
A quality control algorithm $A$ for a problem $(\mathcal{D},\rho)$ can always run a quality control algorithm $A^{\prime}$ for an unrelated parameter $\tau$ on the same distribution $\mathcal{D}$ and filter out (reject) instances that are rejected by $A^{\prime}$.
In particular running $A^{\prime}$ as a filter preserves soundness simply by definition and does not hurt completeness since it only rejects instances from $\mathcal{D}$ with probability $o(1)$. (We note that standard definitions in sublinear algorithms and average complexity do not readily admit such filtering, making it harder to compose algorithms designed to satisfy those definitions.) Composability is also useful when analyzing the marginals of $\mathcal{D}$. Suppose that $\mathcal{D}$ is supported on $\mathcal{X}$, and $S\subseteq\mathcal{X}$. If the marginal of $\mathcal{D}$ on $S\mathcal{D}|_{S}$ — has a concentrated parameter $\tau$, for a large proportion of $S$ of a certain size, we can run $(\mathcal{D}|_{S},\tau)$-quality control as a filter as well.
Of course, in order to be useful, $\tau$ ought to be related to $\rho$ in some ways. The key to our upper bound results turns out to be in identifying the right set of related parameters $\tau$ that make quality control for $\rho$ easier. In the following, we describe these related parameters and then show how we get our results from these.
Clique counting: Upper Bounding Query Complexity
Our algorithm for quality control for the problem $(\mathcal{G}_{n,p},\rho_{k})$ (i.e., testing if number of $k$-cliques in $G$ is roughly what it should be for a random graph from $\mathcal{G}_{n,p}$) is the following: Given a graph $G$, we sample a random (multi-set) $S$ of $s\approx p^{-O(k)}$ vertices of $G$ and count the number of $\ell$-cliques in the induced subgraph of $G$ on $S$, for every $\ell\in\{1,\ldots,k\}$. If all these numbers are consistent with the expectations for $\mathcal{G}_{n,p}$, then we accept; else, we reject. (So in this case the related parameters are $\rho_{\ell}$ for $\ell<k$.)
We prove that this algorithm performs quality control for $(\mathcal{G}_{n,p},\rho_{\ell})$ for all $\ell\leq k$. We use the fact that the algorithm is a quality control algorithm for the count of $\ell$-cliques to improve its query complexity for quality control of the count of $(\ell+1)$-cliques.
Particularly, quality control on $\ell^{\prime}$-cliques, $\ell^{\prime}\leq k$, allows us to say the following: if, for some $\ell\leq k-1$, all of the $\ell^{\prime}$-clique counts for $\ell^{\prime}\leq\ell$ are consistent with the expectations for $\mathcal{G}_{n,p}$, then a global property of the graph which we call exponentially robust quasirandomness must hold. Exponentially robust quasirandomness, defined below, captures the property that most (multi-)sets of size $s$ have clique counts that correspond to the expected number of cliques in $\mathcal{G}_{s,p}$. This strong property is possessed by $\mathcal{G}_{n,p}$, can be efficiently verified with high probability with a single sampled multiset of size $s$, and allows for more efficient counting of $(\ell+1)$-cliques.
Typical definitions of quasirandomness involve some strong property of graphs that requires that some local condition holds with probability one (i.e., for all choices of some variables). A notable example is jumbledness [Tho87], which requires that some property holds for all pairs (or triples) of vertices. As a result, these properties are not efficiently testable. Exponentially robust quasirandomness is a relaxation of quasirandomness that holds “with high probability,” is efficiently testable, and is strong enough to give us what we need with respect to motif counts.
Therefore, we see that the filtering/composability property of quality control allows us to argue that accepted graphs have a strong property (exponentially robust quasirandomness) that, particularly, holds for Erdős–Rényi graphs.
In the following, let $a\propto b$ stand for “$a$ is proportional to bb”, which means that, according to some natural and consistent scaling $c$, $a=b\cdot c$.
Definition 1.5 (Exponentially robustly quasirandom (informal)).
We say that a graph $G=([n],E)$ is $(\varepsilon,\alpha,s_{0})$-exponentially robustly quasirandom with respect to the count of $\ell$-cliques if for every $s\geq s_{0}$ we have that
$\Pr_{X_{1},\ldots,X_{s}\sim_{\mathrm{i.i.d.}}[n]}\left[\text{\# $\ell$-cliques in $\{X_{1},\ldots,X_{s}\}$}\propto(1\pm\varepsilon)\binom{n}{\ell}p^{\binom{\ell}{2}}\right]\geq 1-\exp(-\alpha s).$
The formal definition is given in Section 3.1.
In the proof of correctness, the property that exponentially robust quasirandomness on $\ell$-cliques for increasing $\ell$ can be checked by our algorithm is given by the following two lemmas.
First, consider the following lemma for $\ell\in\{3,4,\dots,k\}$. For a multiset $S$, let $C_{\ell}(S)$ be the number of $\ell$-cliques in $S$ (defined formally in Section 3.1).
Lemma 1.6 (Inductive lemma (informal)).
Consider any parameter $s>0$. There exist parameters $\alpha_{\ell-1},\alpha_{\ell}$ and $\varepsilon_{\ell-1},\varepsilon_{\ell}$ such that the following holds. Suppose that the graph $G$ is $(\varepsilon_{\ell-1},\alpha_{\ell-1},s)$-exponentially robustly quasirandom with respect to the count of $(\ell-1)$-cliques. Define tester $T_{\ell}$ that samples (multi-set) $S$ with $s\approx p^{-O(\ell)}$ vertices of $G$ and accepts if and only if the $\ell$-clique count in $S$ is consistent with the expectations for $\mathcal{G}_{n,p}$.
Then, with probability at least $1-\exp(-\alpha_{\ell}s)$, $T_{\ell}$ accepts graphs that are $(\varepsilon_{\ell-1},\alpha_{\ell},s)$-exponentially robustly quasirandom with respect to the count of $\ell$-cliques and rejects graphs that are not $(\varepsilon_{\ell},\alpha_{\ell},s)$-exponentially robustly quasirandom with respect to the count of $\ell$-cliques.
The inductive lemma follows naturally from the following “concentration lemma.”
Lemma 1.7 (Concentration lemma (informal)).
Consider any parameter $s>0$. There exist parameters $\alpha_{\ell-1},\alpha_{\ell}$ and $\varepsilon_{\ell-1},\varepsilon_{\ell}$ such that the following holds. Suppose that the graph $G$ is $(\varepsilon_{\ell-1},\alpha_{\ell-1},s)$-exponentially robustly quasirandom with respect to the count of $(\ell-1)$-cliques.
Then,
$\mathbb{P}_{S\sim[n]^{s}}\left(C_{\ell}(S)\propto(1\pm\varepsilon_{\ell-1})\cdot C_{\ell}(G)\right)\geq 1-\exp(-\alpha_{\ell}s).$
These lemmas are formalized and proved as Lemma 3.2 and Lemma 3.3, respectively, in Section 3.1.
The proof of the concentration lemma is the technical core of the proof of Theorem 1.1. It follows from a repeated application of the following concentration inequality for functions satisfying a “high-probability” bounded differences property:
Theorem 1.8 ([Com24] (Modified, simplifed, informal)).
Consider a function $f:\{0,1\}^{N}\to\mathbb{R}$ and a subset of the domain $\mathcal{Y}\subseteq\{0,1\}^{N}$. Suppose that there exists a $c\geq 0$ such that for all $\vec{x},\vec{y}\in\mathcal{Y}$,
$\left|f(\vec{x})-f(\vec{y})\right|\leq c\cdot\#(i:x_{i}\neq y_{i}).$
Let $X_{1},X_{2},\dots,X_{N}$ be 0/1-valued independent random variables; denote $X=(X_{1},X_{2},\dots,X_{N})$. Let $q=\mathbb{P}\left(X\not\in\mathcal{Y}\right)$. Then, for any $\delta\geq q\cdot\max(f)$:
$\mathbb{P}\left(\left|f(X)-\mathbb{E}(f(X))\right|\geq\delta\right)\lessapprox q+\exp\left(-\delta^{2}/(2Nc^{2})\right).$ (1)
The formal statement is given as Section 2.4 in Section 2.
Below, let $\mu_{\ell}(s)$ be the expected number of $\ell$-cliques in a multiset of size $s$ of vertices, when the underlying graph is drawn from $\mathcal{G}_{n,p}$. To prove the concentration lemma using Theorem 1.8, we let the random variable $X_{i}$ be the $i$th vertex sampled (with repetition) by the algorithm for the multiset $S$ of size $s$. These are independent. Let $\mathcal{Y}$ be the subset of size-$s$ multisets with $(\ell-1)$-clique counts proportional to the expected amount from $\mathcal{G}_{n,p}$. By the exponentially robust quasirandomness assumption on the count of $(\ell-1)$-cliques, and the setting of $s$, $q$ is very small. We set, for a constant $a>0$, $\delta=\frac{\varepsilon}{a}\cdot\mu_{\ell}(s)$. We prove that the “bounded difference” $c$ satisfies $c\leq\mu_{\ell-1}(s)$, and so the right-hand side of Equation 1 will approximately be (omitting constants);
$q+\exp\left(-\delta^{2}/(2Nc^{2})\right)\approx\exp\left(-\varepsilon^{2}\mu_{\ell}(s)^{2}/(s\cdot\mu_{\ell-1}(s)^{2})\right).$
Plugging in the values of $\mu_{\ell-1}(s),\mu_{\ell}(s)$ yields the concentration lemma, where $\alpha_{\ell}$ depends on $\mu_{\ell}(s)^{2}/\mu_{\ell-1}(s)^{2}$ and $s$. More concretely, $\alpha_{\ell}=p^{O(\ell)}$ in the concentration lemma.
Ultimately, to obtain a low error probability for the algorithm, we need to set $s\approx\Omega(1/\alpha_{k})$, corresponding to $s\approx 1/p^{-O(k)}$.
The inductive lemma then follows easily from the concentration lemma. This is because, given the assumption about exponentially robust quasirandomness for $(\ell-1)$-cliques, the concentration lemma tells us that the number of $\ell$-cliques in one sampled multiset $S$ will reveal whether the graph’s global count of $\ell$-cliques is consistent with the expected number from $\mathcal{G}_{n,p}$. Moreover, if it is consistent, the concentration lemma (with a bit of manipulation of expressions) also implies exponentially robust quasirandomness of $\ell$-cliques, for some parameters.
Observe that the inductive and concentration lemmas already imply that the stated algorithm is a $(\mathcal{G}_{n,p},\rho_{k})$-quality control algorithm. The algorithm stated performs $T_{\ell}$ from the inductive lemma for all $\ell\in[k-1]$. If the graph is not $(\varepsilon_{\ell},\alpha_{\ell},s)$-exponentially robustly quasirandom with respect to $\ell$-cliques for any $\ell\in[k-1]$, it will reject, by the inductive lemma. If the graph is $(\varepsilon_{k-1},\alpha_{k-1},s)$-exponentially robustly quasirandom with respect to (k1)(k-1)-cliques, then the concentration lemma implies that the algorithm will detect whether the number of $k$-cliques corresponds to $\mathcal{G}_{n,p}$ or not. We prove that $G\sim\mathcal{G}_{n,p}$ is $(\varepsilon_{\ell-1},\alpha_{\ell},s)$-exponentially robustly quasirandom for all $\ell\in[k]$ with high probability, and so the algorithm accepts input graphs drawn from $\mathcal{G}_{n,p}$ with high probability.
The formal proof can be found in Section 3.1.
Clique counting: Upper Bounding Time
As stated so far, the algorithm uses $p^{-O(k)}$ queries, for all adjacencies in $S$. However, counting the number of $\ell$-cliques, for $\ell\in[k]$, in a worst case graph could take time $p^{-O(k^{2})}$. We again take advantage of the composability of quality control problems to improve the runtime to $p^{-O(k)}$. We now define a quality parameter $\rho_{\mathcal{J}}$ that corresponds to the jumbledness of a graph. Graph jumbledness [Tho87] – one of the earliest forms of graph pseudorandomness studied – aims to capture a property that is held by Erdős–Rényi random graphs. It captures the property that the number of edges between any vertex subsets $X,Y$ is approximately $p|X||Y|$ in $\mathcal{G}_{n,p}$ (see Section 2.2 for a formal definition).
Let $\rho_{\mathcal{J}}(G)=1$ if the graph is jumbled (for some specified parameters; see Section 3.3), and $\rho_{\mathcal{J}}(G)=0$ otherwise. Therefore, $(\mathcal{G}_{s,p},\rho_{\mathcal{J}})$-quality control should accept graphs from $\mathcal{G}_{s,p},$ (which are jumbled with high probability), and reject graphs that are not jumbled, with high probability.
For a graph on $s$ vertices, verifying jumbledness takes time $O(s^{3})$. Additionally, counting $k$-cliques on a jumbled graph (for certain jumbledness parameters; see Section 2.2 and Section 3.3) takes time $O(s^{c})$ for some universal constant $c$. Therefore, performing $(\mathcal{G}_{s,p},\rho_{\mathcal{J}})$-quality control on the subgraph of the vertices sampled will improve the runtime to $1/p^{O(k)}$.
We therefore consider the following updated algorithm: Given a graph $G$, sample a random (multi-set) $S$ of $s\approx p^{-O(k)}$ vertices of $G$. Verify if the subgraph on $S$ is jumbled. If it is not, reject. Else, using an efficient algorithm for jumbled graphs, count the number of $\ell$-cliques in the induced subgraph of $G$ on $S$, for every $\ell\in\{1,\ldots,k\}$. If all these numbers are consistent with the expectations for $\mathcal{G}_{n,p}$, then we accept; else, we reject.
The formal proof can be found in Section 3.3.
Triangle Counting
Our results so far already imply that quality control of the triangle count compared to $\mathcal{G}_{n,p}$ takes $1/p^{O(1)}$ adjacency matrix queries and runtime. For the case of triangles, we identify the exact constant in the exponent, proving a $\widetilde{O}(1/p)$ upper bound on the query complexity and runtime and a $\Omega(1/p)$ lower bound. The algorithm will rely on results from the literature on sublinear approximation of graph parameters [EMR22, ERS20a], and will correspondingly utilize more expansive sublinear access to the graph: query access to adjacency lists, the adjacency matrix, and degree queries.
Suppose that we want to apply an algorithm for estimating the number of triangles in a worst-case graph to the quality control setting. Such an algorithm would approximate the number of triangles, and accept the graph if the number corresponds approximately to the expected amount in $\mathcal{G}_{n,p}$ and reject otherwise. This is a plain $(\mathcal{G}_{n,p},\rho_{\Delta})$-quality control problem, without composing the problem with any other quality control questions. Without any other knowledge of the underlying graph, this takes $\Theta(p^{-3})$ queries (ignoring polylogarithmic factors) [ELRS17]. If we first perform $(\mathcal{G}_{n,p},\rho_{e})$-quality control, where $\rho_{e}(G)=e(G)/\mathbb{E}_{G^{\prime}\sim G}\left[e(G^{\prime})\right]$, then subsequent $(\mathcal{G}_{n,p},\rho_{\Delta})$-quality control on the accepted graphs takes $\Theta(p^{-3/2})$ queries (ignoring polylogarithmic factors) [ELRS17]. To get to $\widetilde{O}(1/p)$, we first perform $(\mathcal{G}_{n,p},\rho_{e})$-quality control and $(\mathcal{G}_{n,p},\rho_{\alpha})$-quality control, where $\alpha(G)$ is the arboricity of the graph $G$ (see Section 2.1) and $\rho_{\alpha}(G)=\alpha(G)/\mathbb{E}_{G^{\prime}\sim G}\left[\alpha(G^{\prime})\right]$. On graphs accepted by both of the quality control algorithms, perform $(\mathcal{G}_{n,p},\rho_{\Delta})$-quality control. Using algorithms from [ERS20a] for the approximation of $k$-clique counts in graphs with a bounded arboricity, we can reduce the bounds to $\widetilde{O}(1/p)$ queries and runtime.
For the lower bound on $(\mathcal{G}_{n,p},\rho_{\Delta})$-quality control, we construct two distributions over graphs: a YES distribution, which is forced to be $\mathcal{G}_{n,p}$ since the quality control problem has this as the target distribution, and a NO distribution, which must have a number of triangles out of the range $(1\pm\varepsilon)\mathbb{E}_{G^{\prime}\sim G}\left[C_{\Delta}(G^{\prime})\right]$, where $C_{\Delta}(G)$ is the number of triangles in graph $G$. Since we want the NO distribution to look indistinguishable from $\mathcal{G}_{n,p}$ except for in a small region, we plant a clique of size $\ell=O(pn)$ in a graph drawn from $\mathcal{G}_{n,p}$. Distinguishing a graph from itself with a planted clique has previously been studied [ERS20a, RS20], but differs either because the underlying graph is not $\mathcal{G}_{n,p}$ [ERS20a] or the analysis is different [RS20].
The graph created by planting an $\ell$-clique in $\mathcal{G}_{n,p}$ has more than $(1\pm\varepsilon)\mathbb{E}_{G^{\prime}\sim G}\left[C_{\Delta}(G)\right]$ triangles, and a query must be made inside the planted clique to identify that the graph does not look like $\mathcal{G}_{n,p}$ and is from the NO distribution. Since the input is randomized, it will take $\Omega(1/p)$ queries until a query is made involving a vertex/edge inside the planted clique.
The formal proofs can be found in Appendix A. In Section A.3, we explain why this approach would likely yield a $1/p^{O(k^{2})}$ query complexity and runtime, for larger $k$.
Motif counting: Upper Bounding Query Complexity and Time
We extend the principles of the proof of quality control of clique counts to general motifs. First, we can define exponentially robust quasirandomness naturally for general induced labeled motifs. We prove inductive and concentration lemmas for the case of motifs, as well. The changes are the following:
  1. 1.
    For the inductive lemma, we now aim to test the exponentially robust quasirandomness of the count of a labeled induced motif $H$ on $\ell$ vertices. The assumption we make now is that for all $H^{\prime}\subseteq H$ such that $|H^{\prime}|=\ell-1$, the graph is $(\varepsilon_{\ell-1},\alpha_{\ell-1},s)$-exponentially robustly quasirandom, for some parameters $\varepsilon_{\ell-1},\alpha_{\ell-1},s$.
  2. 2.
    The concentration lemma makes a similar new assumption about the count of all $H^{\prime}\subseteq H$ such that $|H^{\prime}|=\ell-1$. We compare the count of $H$ in $S$ (see Section 5) to the global count in the concentration lemma.
For motif $H$, let $\mu_{H}(s)$ be the expected number of labeled induced copies of $H$ in a multiset of size $s$ of vertices chosen from $G\sim\mathcal{G}_{n,p}$. The size $s$ of the multiset sampled is chosen so that $\mu_{H}(s)^{2}/(s\cdot\mu_{H^{\prime}}(s)^{2})$ is large for all $H^{\prime}\subset H,|H^{\prime}|=\ell-1$. Therefore, we need to look at $H^{\prime}$ maximizing $\mu_{H^{\prime}}(s)^{2}$. Since $p\leq 1/2$, this corresponds to the $H^{\prime}$ obtained by removing the maximum degree vertex from $H$. The ratio $\mu_{H}(s)^{2}/(s\cdot\mu_{H^{\prime}}(s)^{2})$ will therefore depend on $p^{O(\Delta(H))}$, where $\Delta(H)$ is the maximum vertex degree in $H$. From this observation, we will find that we need $s\approx p^{-O(\Delta(H))}$.
The soundness of the algorithm is argued over any worst-case graph. The completeness argument only relies on the fact that $\mathcal{G}_{n,p}$ has concentration of counts of motifs of size at most $\ell$. Therefore, our quality control results additionally extend to other distributions over random graphs that possess good concentration of small motif counts.
The formal proofs can be found in Section 4 (motifs over $\mathcal{G}_{n,p}$) and Section 5 (motifs over general random graph distributions).
Motif Counting: Lower Bound
As in the lower bound for quality control of the triangle count, the YES distribution is $\mathcal{G}_{n,p}$. For the NO distribution, plant a structure in a graph drawn from $\mathcal{G}_{n,p}$ so that the number of induced labeled copies of a motif $H$ is greater than $(1\pm\varepsilon)\mathbb{E}_{G^{\prime}\sim G}\left[C_{H}(G^{\prime})\right]$, where $C_{H}(G)$ is the number of copies of $H$ in a graph $G$. We cannot just plant a large $\ell$-clique in $G$ this time, however, since this may not add enough induced copies of $H$. Instead, we plant a graph drawn from $\mathcal{G}_{\ell,1/2}$ for $\ell=O(np^{\Delta(H)})$, and connect each of these $\ell$ vertices to other vertices in the graph, with edge probability $1/2$. That is, the NO distribution is a stochastic block model with two communities; Community 1 has size $n$-$\ell$ and in-community edge probability $p$, Community 2 has size $\ell$ and in-community edge probability $1/2$, and the between-community edge probability is $1/2$.
Given this construction, it takes $\Omega(p^{-\Delta(H)})$ queries to distinguish the YES and NO distributions. The choice of $\ell=O(np^{\Delta(H)})$ arises because the vertices in the planted/denser region play the role of the maximum-degree vertex in new $H$ copies we count (and copies of $H^{\prime}=H$ minus this vertex are counted from existing copies in the $\mathcal{G}_{n,p}$ part of the graph).
The formal proof can be found in Section 4.2.
1.5 Open problems and future directions
Our work opens up a variety of future research directions. We begin with open problems that are more directly related to quality control of motif counts for $\mathcal{G}_{n,p}$, and then turn to broader questions in quality control.
Quality control of motif counts
First, our work has characterized $(\mathcal{G}_{n,p},\rho_{H})$-quality control by proving that there exist universal constants $c_{1},c_{2}$ such that the query complexity $q(n)$ and runtime $t(n)$ are $1/p^{c_{1}\Delta(H)}\leq q(n),t(n)\leq 1/p^{c_{2}\Delta(H)}$. Identifying/proving the precise universal constants in the exponent remains open.
Second, when $p=o\left(n^{-1/(ck)}\right)$ for a constant $c$, quality control remains unstudied. For the problem of counting motifs compared to $\mathcal{G}_{n,p}$, our completeness guarantee required that $G\sim\mathcal{G}_{n,p}$ satisfies $\rho_{k}\approx 1$ with high probability (where $\rho_{k}$ is the quality function for the count of $k$-cliques), which is true when $n=\Omega\left(p^{-ck}\right)$. It seems natural that such a property needs to hold, given our formulation of quality control. Is there a different natural notion of $(D,\rho)$-quality control when the distribution $D$ we are comparing against does not have concentration of the value of $\rho$?
Third, $\mathcal{G}_{n,p}$ has many other graph parameters that are well-concentrated. Understanding quality control for these graph parameters over $\mathcal{G}_{n,p}$ is a compelling future direction. For example, understanding quality control for the minimum vertex cover, diameter, and chromatic number are suggested future directions.
Fourth, when sublinear algorithms for counting motifs are studied, it is often natural to study sampling motifs as well [BER21, ELRR25, ERR22]. While sampling may not fit as naturally with counting in the quality control framework, what can we say about efficient algorithms for sampling motifs in graphs accepted by the algorithm? The notion of exponentially robust quasirandomness suggests that sampling a random multiset of a certain size $s$ and choosing a uniform random $k$-clique in the multiset could suffice to produce an almost-uniform sample of a motif in the graph. Such an algorithm would have an efficient query complexity, but understanding and improving its runtime would require further consideration.
Finally, there is still more to understand about the runtime for quality control over other random graph distributions. While we have proven a query complexity and runtime bound for $(\mathcal{D}_{n},\rho_{H})$-quality control for general random graph distributions $\mathcal{D}_{n}$, the runtime requires further study because we are not able to extend the techniques from $\mathcal{G}_{n,p}$ as directly. The query complexity argument was extendable to general graph distributions because it had only relied on $\mathcal{G}_{n,p}$’s property of having concentration of the counts of small subgraphs. However, the runtime argument for $\mathcal{G}_{n,p}$ required testing whether the input graph is jumbled (Section 2.2), which holds for $G\sim\mathcal{G}_{n,p}$ with high probability but not necessarily for graphs drawn from other random graph distributions one might like to consider. For example, in Theorem 1.5, while the query complexity is $1/p^{O(k)}$ for quality control of the $k$-clique count, the runtime still would scale with $1/p^{O(k^{2})}$ for other distributions over graphs.
Other directions in quality control
In this paper, we performed a case study of our newly-identified class of quality control problems by studying the quality control of motif counts over Erdős–Rényi graphs. However, quality control need not be studied only over graphs, nor does it need to be studied in the sublinear access model only. The question of whether an algorithm that works well on average over a distribution can safely be applied to an input is relevant for a variety of different inputs. For example, one could imagine performing $(D,\rho)$ quality control where $D=\mathcal{N}(0,I_{n})$ (and so the input is a single vector drawn from this high-dimensional Gaussian distribution) and $\rho_{k}$ is the $k$th norm (for some reasonable definition) of a vector $v\in\mathbb{R}^{n}$. What can be said about quality control in such settings beyond graphs? Additionally, global algorithms also naturally fit into the quality control model. Our proofs relied on a local-to-global property of exponentially robust quasirandomness (local motif counts reveal information about global counts); see Theorem 1.2 for the global algorithm we considered. Imaginably, global algorithms can also contribute to the understanding of quality control in sublinear access models for other families of input objects and problems.
Additionally, understanding the algorithmic implications of quality control – both from a theoretical and practical perspective – has many open directions. As in the testable learning paradigm of [RV23], in some settings, testing whether an algorithm is safe to use on an input provides improved guarantees. In our context, algorithms that assume the graph is from $\mathcal{G}_{n,p}$, but really only use motif counts, would be safe to lift from average-case graphs to worst-case graphs (which may be deemed safe to use the algorithm on, or otherwise rejected), given the results in this paper. There is much unexplored regarding understanding the ways in which quality control can be used as a component of a broader algorithmic pipeline.
1.6 Paper outline
The rest of the paper is organized as follows. First, in Section 2, we introduce preliminary definitions and results related to graphs and sublinear access models, pseudorandom graphs, properties of Erdős–Rényi graphs, and bounded difference concentration inequalities. In Section 3.1, we study quality control of $k$-clique counts compared to Erdős–Rényi graphs. Section 3.1 and Section 3.3 incorporate the main technical elements for the query complexity and runtime proofs, which will be applied and extended in later sections for motif counts, other random graph models, and general graph parameters. Section 4 extends the results to quality control of counts of general motifs compared to Erdős–Rényi graphs. The main technical elements of the lower bound proofs in the paper can be found in Section 4.2. We recommend that a reader primarily interested in the core technical ideas of our paper focus on Section 3.1, Section 3.3, and Section 4.2. Section 5 looks at quality control for motif counts compared to other random graph models with concentration. This section includes the most general forms of our results for quality control of motif counts and the technical lemmas.
In Appendix A, we focus on the case of counting triangles compared to Erdős–Rényi graphs, relying on results from the literature on sublinear approximate counting to strengthen the quality control result for triangles. Next, in Sections B and C of the Appendix, we provide two alternative perspectives on quality control of motifs. First, Appendix B further extends the results for quality control of graph parameters expressible as “sum-NC0” circuits (see Appendix B). Second, in Appendix C, we describe the interplay between average-case complexity and quality control and show how our quality control algorithm implies an algorithm for approximately counting motifs with average-case runtime $1/p^{O(\Delta(H))}$ where $\Delta(H)$ is the maximum vertex degree of a motif $H$. Lastly, Appendix D provides proofs of concentration of subgraph counts for the Erdős–Rényi graph model and another model used for our lower bound proofs.
2 Preliminaries
In this section, we provide a background on relevant notation, definitions, and results related to graphs, sublinear access models, pseudorandom graphs, properties of Erdős–Rényi graphs, and concentration inequalities. We also prove some elementary results. The preliminary results that we prove in this paper are Section 2.2, Section 2.3, Section 2.3, Section 2.3, Corollary Section 2.3, Section 2.3, Section 2.3, and Section 2.4.
A reader familiar with graph theory and sublinear access models can skip Section 2.1. Section 2.2 introduces the notion of jumbled graphs from [Tho87], highlighting known result proving that jumbledness is efficiently verifiable (Section 2.2) and implies efficient algorithms for approximate motif counting (Section 2.2). Section 2.3 covers various properties we will utilize about Erdős–Rényi graphs: motif counts are concentrated around their expectations, arboricity is bounded with high probability, and the graph is jumbled with high probability. Section 2.4 highlights McDiarmid’s inequality [M+89] and an extension [Com24] that we will crucially rely on.
2.1 Graphs and sublinear access models
We begin by introducing some notation. For a graph $G$, let $V(G)$ be the set of vertices of $G$, and $v(G)=|V(G)|$ be the number of vertices. Let $E(G)$ be the edges, and $e(G)=|E(G)|$ be the number of edges. Let $\alpha(G)$ be the arboricity of the graph $G$, which is defined as the minimal number of forests that are needed to cover the edges of $G$.
For a graph $G$, an induced subgraph $H\subseteq G$ is the subgraph that contains exactly the edges of $G$ whose endpoints are both in $V(H)$.
We will consider algorithms that have query access to the input graph $G$. This means that the algorithm does not have full access to the input graph, and instead must perform queries to access information about the underlying structure of the graph. The main object of interest in this paper is the query complexity of the algorithm, which is the maximum number of queries that the algorithm must make on any input graph in order to produce its output.
We consider algorithms under the following query access models (see Chapters 8 and 9 of [Gol17] on testing graph properties):
  1. 1.
    Adjacency list queries: The algorithm makes queries of the form $(u,i)$ for $u\in V(G)$ and $i\in[n-1]$. If $u$ has at least $i$ neighbors, it obtains as a response the $i$-th neighbor of vertex $u$, according to some arbitrary but consistent ordering of the neighbors of $u$. Else, it obtains as a response $\bot$.
  2. 2.
    Adjacency matrix queries: The algorithm makes queries of the form $(u,v)$ for $u,v\in V(G)$. It obtains as a response a $1$ if $(u,v)\in E(G)$ and 0 otherwise.
When the adjacency lists of vertices are sorted (according to some arbitrary, consistent ordering of the vertices), adjacency list access is so powerful that it can simulate adjacency matrix access. Particularly, a query to the adjacency matrix can be answered with $O(\log n)$ queries to a vertex’s sorted adjacency list. Most work on sublinear algorithms for approximating motif counts on graphs (e.g., [AKK19, ELRS17, ERS20a]) assumes access to both the adjacency matrix and adjacency lists (and sometimes even more, such as access to random edge samples). On the other hand, literature on hypothesis testing for planted solution problems in sublinear access models typically assumes adjacency matrix access only (e.g., [RS20, HMP24]).
We will primarily study quality control algorithms with adjacency matrix queries. However, we will also study algorithms with adjacency list queries in certain settings (triangle counting); leveraging previous results that use adjacency list queries allows us to obtain improved and optimal query complexity in this setting.
We now define some relevant terminology related to graphs.
Definition 2.0 (Graph parameter).
Let $\mathcal{G}(n)$ be the family of graphs on $n$ vertices. A graph parameter is a function $F:\mathcal{G}(n)\to\mathbb{R}$ that is invariant under graph isomorphism. That is, for any graph $G\in\mathcal{G}(n)$, and any graph $G^{\prime}$ resulting from permuting the vertices of $G$, $F(G)=F(G^{\prime})$.
Definition 2.0 (Motif).
A motif of size $k$ on a graph $G$ is a labeled induced subgraph on $k$ vertices of $G$. Typically, motifs refer to small graphs that frequently match size-$k$ induced subgraphs of a random graph model being considered.
Definition 2.0 (Motif count).
For a motif $H$ and a graph $G$, let $C_{H}(G)$ be the number of labeled, induced copies of $H$ in $G$. Equivalently, $C_{H}(G)$ equals the number of injective maps $\psi:V(H)\to V(G)$ such that for all $(u,v)\in E(H)$, $(\psi(u),\psi(v))\in E(G)$, and for all $(w,x)\not\in E(H)$, $(\psi(w),\psi(x))\not\in E(G)$.
Definition 2.0 (Arboricity).
The arboricity $\alpha(G)$ of a graph $G$ is the minimal number of forests that the graph’s edges can be partitioned into.
The arboricity of a graph measures the graph’s density, in the sense that, in a graph of arboricity at most $\alpha$, every subgraph has average degree at most $O(\alpha)$.
2.2 Pseudorandom graphs
One of the main characterizations of graphs that we leverage is that of jumbledness. Initially defined by Thomason [Tho87] and extended to more general notions of graph pseudorandomness by subsequent papers – notably, [CGW89] – jumbledness characterizes some respect in which the edges in the graph are uniform/regular. The property of being jumbled can be checked efficiently in the size of the graph and implies efficient algorithms for subgraph counting.
Definition 2.0 (Jumbled graph ([Tho87] as presented in [KS06])).
A graph $G=(V,E)$ is $(p,\beta)$-jumbled if for every pair of vertex subsets $X,Y\subseteq V$ the following holds:
[…]
Above, $e(X,Y)$ counts unordered edges between $X$ and $Y$. When $X=Y$, $e(X,Y)$ thus equals the number of edges in the subgraph on $X$.
Let us understand the values of $(p,\beta)$ for different graphs. First, Erdős–Rényi graphs $\mathcal{G}_{n,p}$ are known to be almost surely $(p,O(\sqrt{np}))$-jumbled [KS06]. In general, the definition above implies $\beta\geq 0$. $\beta=0$ when the graph $G=(V,E)$ is a complete bipartite graph and $p=1$, or $G$ is the empty graph and $p=0$. On the other hand, the value of $\beta$ could be high ($\Theta(np)$), setting $p$ to be the average degree of the graph) if there exists a very sparse or dense region in the graph, compared to the average degree of the graph.
Setting $\beta\geq n$ is trivial, because for any graph $e(X,Y)\leq|X||Y|$, and so $\beta\geq n\geq\sqrt{|X||Y|}$ yields an inequality that is always true/provides no meaningful information. Indeed, the property of being $(p,\Theta(np))$-jumbled does not reveal much about the graph, as linear-sized very sparse/dense subgraphs can still exist. On the other hand, $(p,o(np))$-jumbledness in the dense case (constant $p$) implies strong properties about the underlying graph [CGW89]. In general, graph sequences that satisfy these strong properties are called quasirandom (or pseudorandom) graphs. See the survey by Krivelevich and Sudakov on pseudorandom graphs [KS06] for further information.
Thomason [Tho87] gives efficiently verifiable sufficient conditions for jumbledness.
Proposition 2.0 ([Tho87] as in [KS06]).
Let $G=(V,E)$ be a graph with $|V|=n$ such that every vertex has degree at least $np$, for $p\in(0,1)$. If every pair of vertices in $G$ has at most $np^{2}+\ell$ common neighbors, then $G$ is $(p,\sqrt{(p+\ell)n})$-jumbled.
The proof of the proposition above is not readily found online. We prove a similar statement, given below.
Proposition 2.0.
Let $G=(V,E)$ where $|V|=n$. Consider a parameter $\delta>0$. If every vertex in $G$ has degree in $(n-1)(p\pm\delta)$, and every pair of distinct vertices in $G$ has $(n-2)(p^{2}\pm\delta)$ common neighbors, then $G$ is $(p,n\sqrt{3\delta}+\sqrt{n(\delta+p)})$-jumbled.
Proof.
[…]
$\leq\sum_{x\in V(G)}\left(\text{deg}(x,Y)-p|Y|\right)^{2}=\sum_{x\in V(G)}\left(\text{deg}(x,Y)^{2}-2p|Y|\text{deg}(x,Y)+p^{2}|Y|^{2}\right).$
Using $\sum_{x\in V(G)}\text{deg}(x,Y)^{2}=\sum_{y,y^{\prime}\in Y}\text{codeg}(y,y^{\prime})$ (where $\text{codeg}(y,y)=\text{deg}(y)$), and also that $\sum_{x\in V(G)}\text{deg}(x,Y)=\sum_{y\in Y}\text{deg}(y)$, the above expression equals:
$\sum_{y,y^{\prime}\in Y}\text{codeg}(y,y^{\prime})-2p|Y|\sum_{y\in Y}\text{deg}(y)+p^{2}|Y|^{2}n.$
By the assumptions about vertex degrees and codegrees (shared neighbors) in the lemma, this is:
$\leq|Y|^{2}(n-2)p^{2}\pm|Y|^{2}\delta(n-2)+|Y|(n-1)p\pm|Y|\delta(n-1)-2p|Y|^{2}(n-1)p\pm 2p|Y|^{2}\delta(n-1)+p^{2}|Y|^{2}n.$
Removing terms that cancel each other out, we are left with:
$\leq|Y|^{2}\delta(n-2)+2p|Y|^{2}\delta(n-1)+|Y|(\delta+p)(n-1)\leq 3|Y|^{2}\delta n+|Y|(\delta+p)n.$
Therefore, we have so far found that:
[…]
Equivalently,
[…]
$\sqrt{3|Y|\delta n+(\delta+p)n}\cdot\sqrt{|X||Y|}\leq\sqrt{3\delta n^{2}+(\delta+p)n}\cdot\sqrt{|X||Y|}\leq\left(n\sqrt{3\delta}+\sqrt{n(\delta+p)}\right)\sqrt{|X||Y|}.$
Therefore, $G$ is $(p,n\sqrt{3\delta}+\sqrt{n(\delta+p)})$-jumbled. ∎
Importantly for our goals of testing/counting subgraphs, approximate subgraph counts are efficiently computed (with respect to the size of the graph, not with respect to sublinear access models) in jumbled graphs. This efficient computability result is given in [CFZ14] as an extension of [DLR95]. Before giving the theorem, we must introduce some relevant notation. First, for a graph $H$, let $L(H)$ be the graph whose vertices are the edges of $H$, and two vertices in $L(H)$ are connected by an edge if the corresponding edges in $H$ share an endpoint. Let $\text{max-deg}(L(H))$ denote the maximum degree of any vertex in $L(H)$, and let $D(L(H))$ denote the degeneracy of $L(H)$. Then define $\sigma(H)$ as follows:
$\sigma(H)=\min\left\{\frac{\text{max-deg}(L(H))+4}{2},\frac{D(L(H))+6}{2}\right\}.$ (2)
When $H$ is a $k$-clique, $\sigma(H)=k$, as noted in [CFZ14]. For general motifs, $\sigma(H)\leq\Delta(H)+2$ [CFZ14].
We are now ready to state the result.
Proposition 2.0 (Proposition 9.8 in [CFZ14]).
Consider a graph $H$ on $h$ vertices such that $\sigma(H)\leq j$. Let $\varepsilon>0$. Then, there is an absolute constant $c$, and a constant $C=C(\varepsilon,h)=\exp(\mathrm{poly}(h/\varepsilon))$, such that the following is satisfied.
Suppose that $G$ is a graph on $n$ vertices that is a (spanning subgraph of a) $(p,\beta)$-jumbled graph, where $\beta\leq p^{j}n/C$. Then the number of labeled copies of $H$ in $G$ can be computed up to a $\varepsilon h^{h}\cdot p^{e(H)}\binom{n}{h}$ error, in running time $Cn^{c}$.
The original proposition in [CFZ14] is stated with an error of $\varepsilon p^{e(H)}n^{h}$, which we transform into the stated error above by noting that $n^{h}\leq h^{h}\binom{n}{h}$.
This proposition hinges on the algorithm and analysis of [DLR95], which applies for the dense setting ($p$ is a constant). The algorithm of [DLR95] relies on a variation of the Szemerédi regularity lemma, and [CFZ14] shows that the jumbledness assumption can be used instead of regularity to transfer the algorithm to the setting of sparse $p$.
At a high level, we will use jumbledness and its efficient verification and algorithmic implications for efficient approximate subgraph counting as follows. First, we will reduce the global quality control problem on a graph to the problem of approximately counting motifs in a small subgraph $G_{S}$ on around $s$ vertices. We first efficiently (i.e., in $\text{poly}(s)$ time) test for jumbledness, rejecting graphs that are not sufficiently jumbled. If the graph is jumbled, we then use the algorithm from Section 2.2. As we will observe in Section 2.3, Erdős–Rényi graphs satisfy a jumbledness property on subgraphs of size around $s$, and so for quality control, it is acceptable to reject graphs with subgraphs that are not jumbled.
2.3 Properties of Erdős–Rényi graphs
Proposition 2.0.
Let $H$ be an induced motif on $k$ vertices. Consider parameters $n,p$ with $p=\Omega\left(n^{-1/m_{H}}\cdot\varepsilon^{-2}\right)$, where $m_{H}:=\max_{F\subseteq H}e(F)/v(F)$.
For $G\sim\mathcal{G}_{n,p}$, with probability at least $9/10$, the number of labeled induced copies of $H$ in $G$ is in the range $(1\pm\frac{\varepsilon}{100})\binom{n}{v(H)}v(H)!\cdot p^{e(H)}(1-p)^{\binom{v(H)}{2}-e(H)}$.
[Jan90, JOR04] have proven such a result previously – with more refined guarantees – for the case of non-induced subgraphs. We provide a proof of the result for induced subgraphs in Appendix D. We note that the choice of $1/100$ in the expression above is chosen to best fit our setting, and can be changed to another constant by modifying the constant hidden in the lower bound on $p$.
From this proposition, we can derive the following consequence for the concentration of clique counts for $G\sim\mathcal{G}_{n,p}$.
Proposition 2.0.
For $G\sim\mathcal{G}_{n,p}$ and $p=\Omega\left(n^{-2/(k-1)}\cdot\varepsilon^{-2}\right)$, the number of $k$-cliques in $G$ is in the range $(1\pm\frac{\varepsilon}{100})\binom{n}{k}k!\cdot p^{\binom{k}{2}}$ with high constant probability (say, $9/10$).
Next, [GPS18] proves the following regarding the arboricity $\alpha(G)$ of $G\sim\mathcal{G}_{n,p}$ (which indeed is a precise statement that holds asymptotically almost surely, while we provide a weaker version of their statement below for simplicity).
Proposition 2.0 ([GPS18]).
For all $p=\Omega\left(\frac{1}{n}\right)$, there exists an $N$ such that for all $n\geq N$, $\mathbb{P}_{G\sim\mathcal{G}_{n,p}}\left(\alpha(G)\leq 2\frac{e(G)}{n-1}+1\right)\geq 9/10$.
Combining this with Section 2.3 (for the concentration of the number of edges in $\mathcal{G}_{n,p}$), we obtain the following corollary.
Corollary 2.0.
For $p=\Omega\left(\varepsilon^{-2}\cdot 1/n\right)$, there exists an $N$ such that for all $n\geq N$,
$\mathbb{P}_{G\sim\mathcal{G}_{n,p}}\left(\alpha(G)\leq 2np\right)\geq 4/5.$
When we focus on the algorithmic complexity of quality control for $k$-cliques matching $\mathcal{G}_{n,p}$ in Section 3.3, we will utilize the notion of graph jumbledness (Section 2.2) to obtain an efficient runtime. Our algorithm reduces the global quality control problem to a local problem over a subgraph of size $s$. Then, over $\mathcal{G}_{n,p}$ these subgraphs are jumbled; jumbledness is both efficiently checked Section 2.2 and implies efficient approximate subgraph counting algorithms Section 2.2. With this application in mind, we prove various statements regarding the jumbledness of $G\sim\mathcal{G}_{n,p}$ and its subgraphs.
Proposition 2.0.
When $p\geq\left(\frac{1200C^{4}\log n}{n}\right)^{1/(4k-1)}$, for $\delta=\frac{p^{2k}}{12C^{2}}$, with probability at least 11/n1-1/n, $G\sim\mathcal{G}_{n,p}$ satisfies the following:
  1. 1.
    Every vertex in $G$ has degree in $(n-1)(p\pm\delta)$, and
  2. 2.
    Every pair of vertices in $G$ has $(n-2)(p^{2}\pm\delta)$ common neighbors.
Proof.
We first prove (1). Consider a graph $G\sim\mathcal{G}_{n,p}$, and any vertex $u\in G$. Let us analyze the degree distribution of this vertex. Let $D$ be the random variable corresponding to the degree of vertex $u$. First, the expected degree of vertex $u$ is $E\left(D\right)=(n-1)p$, and the degree distribution of $u$ is a binomial distribution with parameters n1n-1 and $p$. Applying Chernoff bounds,
$\mathbb{P}\left(\left|D-(n-1)p\right|\geq\delta(n-1)\right)\leq 2\exp\left(-(n-1)p\cdot\frac{\delta^{2}}{3p^{2}}\right)=2\exp\left(\frac{-(n-1)p^{(4k-1)}}{432C^{4}}\right).$
When $p\geq\left(\frac{1200C^{4}\log n}{n}\right)^{1/(4k-1)}$, this probability is at most $n^{-2.5}$. By a union bound over all vertices in the graph, we find that with probability at least 11/(2n)1-1/(2n), every vertex in $G\sim\mathcal{G}_{n,p}$ has degree in $(n-1)(p\pm\delta)$.
Next we prove (2) by a similar argument. Consider a graph $G\sim\mathcal{G}_{n,p}$ and any pair of vertices $u,v\in G$. Let’s analyze the distribution of the number of common neighbors (i.e. codegree) of $u$ and $v$. Let $N$ be the random variable corresponding to the number of common neighbors of $u$ and $v$. The expected value of $N$ is $(n-2)p^{2}$, and its distribution is binomial with parameters n2n-2 and $p^{2}$. By Chernoff bounds, we find:
$\mathbb{P}\left(\left|N-(n-2)p^{2}\right|\geq\delta(n-2)\right)\leq 2\exp\left(-(n-2)p^{2}\cdot\frac{\delta^{2}}{3p^{4}}\right)=2\exp\left(\frac{-(n-2)p^{4k-2}}{432C^{4}}\right).$
When $p\geq\left(\frac{1200C^{4}\log n}{n}\right)^{1/(4k-1)}$, the probability is then at most $n^{-3.5}$. By a union bound over pairs of vertices, with probability at least 11/(2n)1-1/(2n), every pair of vertices in $G$ has $(n-2)(p^{2}\pm\delta)$ common neighbors. Therefore, we obtain the proposition. ∎
We observe that, by combining Section 2.2 with Section 2.3, we immediately obtain the following.
Corollary 2.0.
For $p\geq\left(\frac{1200C^{4}\log n}{n}\right)^{1/(4k-1)}$, with probability at least 11/n1-1/n, $G\sim\mathcal{G}_{n,p}$ is $(p,\beta)$-jumbled for $\beta\leq p^{k}n/C$.
Proof.
By Section 2.3, with probability at least 11/n1-1/n, for $\delta=\frac{p^{2k}}{12C^{2}}$, every vertex in $G$ has degree in $(n-1)(p\pm\delta)$, and every pair of vertices in $G$ has $(n-2)(p^{2}\pm\delta)$ common neighbors. By Section 2.2, since $n\sqrt{3\delta}+\sqrt{n(\delta+p)}\leq p^{k}n/C$, this implies that with probability at least 11/n1-1/n, $G\sim\mathcal{G}_{n,p}$ is $(p,\beta)$-jumbled for $\beta\leq p^{k}n/C$. ∎
Given that, for any $1\leq t\leq n$, the marginal of a $t$-vertex subgraph of $G\sim\mathcal{G}_{n,p}$ is distributed according to $G_{t,p}$, we can use the previous proposition to prove the following. For a graph $G$ and subset $S\subseteq[n]$, let $G_{S}$ be the induced subgraph on the vertices in $S$.
Proposition 2.0.
Let $p\geq\left(\frac{1200C^{4}\log n}{n}\right)^{1/(4k-1)}$ and $s\geq 1200C^{4}/p^{5k}$. Let $\delta=\frac{p^{2k}}{12C^{2}}$. With probability at least $9/10$, for $G\sim\mathcal{G}_{n,p}$, at least a $9/10$-fraction of subgraphs $S$ of size $s$ satisfy the following:
  1. 1.
    Every vertex in $G_{S}$ has degree in $(s-1)(p\pm\delta)$ in $G_{S}$, and
  2. 2.
    Every pair of vertices in $G_{S}$ has $(s-2)(p^{2}\pm\delta)$ common neighbors in $G_{S}$.
Proof.
Let us call $G_{S}$ “good” if it satisfies conditions (1) and (2) above.
For $G\sim\mathcal{G}_{n,p}$, let us first analyze the expected number of $S\subseteq[n],|S|=s$ such that $G_{S}$ is not good.
$\mathbb{E}_{G\sim\mathcal{G}_{n,p}}\left(\sum_{S\subseteq[n]:|S|=s}1(G_{S}\text{ is not good})\right)=\sum_{S\subseteq[n]:|S|=s}\mathbb{P}(G_{S}\text{ is not good}).$
Next, observe that the marginal of the graph on $S$ ($G_{S}$) is distributed according to $G_{s,p}$. Also observe that for $s\geq 1200C^{4}/p^{5k}$, $p\geq\left(\frac{1200C^{4}\log s}{s}\right)^{1/(4k-1)}$. Therefore, we can apply Section 2.3 to $G_{S}$. We find that, with probability at most $1/s$, $G_{S}$ is not good.
Therefore, the expected number of $S\subseteq[n],|S|=s$ such that $G_{S}$ is not good is at most $\frac{1}{s}\cdot\binom{n}{s}\leq\frac{1}{100}\binom{n}{s}$ for $s\geq 100$.
For a graph $G$, let $t(G)$ be the number of $S\subseteq[n],|S|=s$ such that $G_{S}$ is not good. By Markov’s inequality,
$\mathbb{P}\left(t(G)\geq\frac{1}{10}\cdot\binom{n}{s}\right)\leq\frac{\mathbb{E}(t(G))}{\frac{1}{10}\cdot\binom{n}{s}}\leq\frac{\frac{1}{100}\cdot\binom{n}{s}}{\frac{1}{10}\cdot\binom{n}{s}}\leq\frac{1}{10}.$
Therefore, at least a $9/10$-fraction of subgraphs of size $s$ are good. ∎
We note the following corollary of this proposition, for ease of use.
Corollary 2.0.
Let $p\geq\left(\frac{1200C^{4}\log n}{n}\right)^{1/(4k-1)}$, $s\geq 1200C^{4}/p^{5k}$, and $\delta=\frac{p^{2k}}{12C^{2}}$. Consider $G\sim\mathcal{G}_{n,p}$, and choose a random subgraph $G_{S}$ of $G$ on $s$ vertices. Then, with probability at least $4/5$, the subgraph on these $s$ vertices satisfies the following:
  1. 1.
    Every vertex in $G_{S}$ has degree in $(s-1)(p\pm\delta)$ in $G_{S}$, and
  2. 2.
    Every pair of vertices in $G_{S}$ has $(s-2)(p^{2}\pm\delta)$ common neighbors in $G_{S}$.
Proof.
This corollary follows by taking a union bound over the events in Section 2.3 that $G\sim\mathcal{G}_{n,p}$ does not satisfy the property that at least a $9/10$-fraction of size-$s$ subgraphs are good and that a non-good subgraph is chosen at random. ∎
2.4 Concentration inequalities
Theorem 2.1 (McDiarmid’s Inequality [M+89]).
Consider a function $f:\mathcal{X}_{1}\times\mathcal{X}_{2}\times\dots\times\mathcal{X}_{t}\to\mathbb{R}$ such that, for all $i\in[t]$ the bounded differences property is satisfied: For all $(x_{1},x_{2},\dots,x_{t})\in\mathcal{X}_{1}\times\mathcal{X}_{2}\times\dots\times\mathcal{X}_{t}$, $i\in[t]$ and $y_{i}\in\mathcal{X}_{i}$,
$\sup_{y_{i}\in\mathcal{X}_{i}}\left|f(x_{1},\dots,x_{i-1},x_{i},x_{i+1},\dots,x_{t})-f(x_{1},\dots,x_{i-1},y_{i},x_{i+1},\dots,x_{t})\right|\leq c_{i}.$
Then, for independent random variables $X_{1},X_{2},\dots,X_{t}$ where $X_{i}\in\mathcal{X}_{i}$ for $i\in[t]$, and for any $\delta>0$:
$\mathbb{P}\left(\left|f(X_{1},X_{2},\dots,X_{t})-\mathbb{E}\left(f(X_{1},X_{2},\dots,X_{t})\right)\right|\geq\delta\right)\leq 2\exp\left(-\frac{2\delta^{2}}{\sum_{i=1}^{t}c_{i}^{2}}\right).$
Theorem 2.2 (McDiarmid’s Inequality when differences are bounded whp [Com24]).
Consider a function $f:\mathcal{X}_{1}\times\mathcal{X}_{2}\times\dots\times\mathcal{X}_{t}\to\mathbb{R}$ and a subset of the domain $\mathcal{Y}\subseteq\mathcal{X}_{1}\times\mathcal{X}_{2}\times\dots\times\mathcal{X}_{t}$ such that the following bounded differences property holds over elements of the subset $\mathcal{Y}$:
For all $(x_{1},x_{2},\dots,x_{t})\in\mathcal{Y}$ and $(y_{1},y_{2},\dots,y_{t})\in\mathcal{Y}$,
$\left|f(x_{1},x_{2},\dots,x_{t})-f(y_{1},y_{2},\dots,y_{t})\right|\leq\sum_{i:x_{i}\neq y_{i}}c_{i}.$
Let $X_{1},X_{2},\dots,X_{t}$ be independent random variables where $X_{i}\in\mathcal{X}_{i}$ for $i\in[t]$. Let $q:=1-\mathbb{P}\left((X_{1},X_{2},\dots,X_{t})\in\mathcal{Y}\right)$.
Then, for any $\delta>0$:
$\mathbb{P}\left(\left|f(X_{1},X_{2},\dots,X_{t})-\mathbb{E}\left(f(X_{1},X_{2},\dots,X_{t})|(X_{1},X_{2},\dots,X_{t})\in\mathcal{Y}\right)\right|\geq\delta\right)$
$\leq 2q+2\exp\left(-\frac{2\max\left(0,\delta-q\sum_{i=1}^{t}c_{i}\right)^{2}}{\sum_{i=1}^{t}c_{i}^{2}}\right).$
We will utilize the following corollary of Theorem 2.2, which compares $f(X_{1},X_{2},\dots,X_{t})$ to its expectation instead of the expectation conditioned on $(X_{1},X_{2},\dots,X_{t})\in\mathcal{Y}$.
Corollary 2.2.
Consider a bounded, nonnegative function $f:\mathcal{X}_{1}\times\mathcal{X}_{2}\times\dots\times\mathcal{X}_{t}\to\mathbb{R}$ and a subset of the domain $\mathcal{Y}\subseteq\mathcal{X}_{1}\times\mathcal{X}_{2}\times\dots\times\mathcal{X}_{t}$ such that the following bounded differences property holds over elements of the subset $\mathcal{Y}$: For all $(x_{1},x_{2},\dots,x_{t})\in\mathcal{Y}$ and $(y_{1},y_{2},\dots,y_{t})\in\mathcal{Y}$,
$\left|f(x_{1},x_{2},\dots,x_{t})-f(y_{1},y_{2},\dots,y_{t})\right|\leq\sum_{i:x_{i}\neq y_{i}}c_{i}.$
Let $X_{1},X_{2},\dots,X_{t}$ be independent random variables where $X_{i}\in\mathcal{X}_{i}$ for $i\in[t]$. Let $q:=1-\mathbb{P}\left((X_{1},X_{2},\dots,X_{t})\in\mathcal{Y}\right)$. Consider any $\delta>0$. If $q\leq\min\{\frac{\delta}{2\max(f)},\frac{\delta}{4\mathbb{E}(f(X_{1},X_{2},\dots,X_{t}))},\frac{1}{2}\}$, then
$\mathbb{P}\left(\left|f(X_{1},X_{2},\dots,X_{t})-\mathbb{E}\left(f(X_{1},X_{2},\dots,X_{t})\right)\right|\geq\delta\right)\leq 2q+2\exp\left(-\frac{2\max\left(0,\delta/2-q\sum_{i=1}^{t}c_{i}\right)^{2}}{\sum_{i=1}^{t}c_{i}^{2}}\right).$
Proof.
For simplicity of notation, let us denote the vector of random variables $(X_{1},X_{2},\dots,X_{t})$ as $\vec{X}$.
We first prove that when $q\leq\min\{\frac{\delta}{2\max(f)},\frac{\delta}{4\mathbb{E}(f(X_{1},X_{2},\dots,X_{t}))},\frac{1}{2}\}$, we have:
$\left|\mathbb{E}\left(f(\vec{X}|\vec{X}\in\mathcal{Y})\right)-\mathbb{E}\left(f(\vec{X})\right)\right|\leq\delta/2.$ (3)
To prove this, we begin by expanding the expectation of $f$ into conditional expectations:
$\mathbb{E}\left(f(\vec{X})\right)=\mathbb{E}\left(f(\vec{X})|\vec{X}\in\mathcal{Y}\right)\cdot\Pr\left(\vec{X}\in\mathcal{Y}\right)+\mathbb{E}\left(f(\vec{X})|\vec{X}\not\in\mathcal{Y}\right)\cdot\Pr\left(\vec{X}\not\in\mathcal{Y}\right).$
Therefore, we have:
$\mathbb{E}\left(f(\vec{X})|\vec{X}\in\mathcal{Y}\right)=\frac{\mathbb{E}\left(f(\vec{X})\right)-\mathbb{E}\left(f(\vec{X})|\vec{X}\not\in\mathcal{Y}\right)\cdot\Pr\left(\vec{X}\not\in\mathcal{Y}\right)}{\Pr\left(\vec{X}\in\mathcal{Y}\right)}=\frac{\mathbb{E}\left(f(\vec{X})\right)-\mathbb{E}\left(f(\vec{X})|\vec{X}\not\in\mathcal{Y}\right)\cdot q}{1-q}.$
We first upper-bound $\mathbb{E}\left(f(\vec{X})|\vec{X}\in\mathcal{Y}\right)$. From the expression above, this is upper-bounded by $\frac{\mathbb{E}\left(f(\vec{X})\right)}{1-q}$. Since $q\leq 1/2$, this is at most $(1+2q)\mathbb{E}(f(\vec{X})).$ Next, since $q\leq\frac{\delta}{4\mathbb{E}(f(\vec{X}))}$, we have:
$\mathbb{E}\left(f(\vec{X})|\vec{X}\in\mathcal{Y}\right)\leq(1+2q)\mathbb{E}(f(\vec{X}))\leq\mathbb{E}(f(\vec{X}))+\frac{\delta}{2}.$
We now lower-bound $\mathbb{E}\left(f(\vec{X})|\vec{X}\in\mathcal{Y}\right)$:
$\mathbb{E}\left(f(\vec{X})|\vec{X}\in\mathcal{Y}\right)\geq\mathbb{E}\left(f(\vec{X})\right)-\mathbb{E}\left(f(\vec{X})|\vec{X}\not\in\mathcal{Y}\right)\cdot q\geq\mathbb{E}\left(f(\vec{X})\right)-\max(f)\cdot q.$
Since $q\leq\frac{\delta}{2\max(f)}$, this expression is at least $\mathbb{E}\left(f(\vec{X})\right)-\frac{\delta}{2}$.
Therefore, we have shown that Equation 3 holds.
Next, observe that $\left|f(\vec{X})-\mathbb{E}\left(f(\vec{X})\right)\right|\geq\delta$ implies that either
$\left|f(\vec{X})-\mathbb{E}\left(f(\vec{X})|\vec{X}\in\mathcal{Y}\right)\right|\geq\delta/2\text{ or }\left|\mathbb{E}\left(f(\vec{X})|\vec{X}\in\mathcal{Y}\right)-\mathbb{E}\left(f(\vec{X})\right)\right|\geq\delta/2.$
Since we know that the second of these inequalities does not hold, $\left|f(\vec{X})-\mathbb{E}\left(f(\vec{X})\right)\right|\geq\delta$ implies $\left|f(\vec{X})-\mathbb{E}\left(f(\vec{X})|\vec{X}\in\mathcal{Y}\right)\right|\geq\delta/2$. Therefore,
$\mathbb{P}\left(\left|f(\vec{X})-\mathbb{E}\left(f(\vec{X})\right)\right|\geq\delta\right)\leq\mathbb{P}\left(\left|f(\vec{X})-\mathbb{E}\left(f(\vec{X})|\vec{X}\in\mathcal{Y}\right)\right|\geq\delta/2\right).$
Applying Theorem 2.2 to the probability on the right-hand side yields the corollary. ∎
3 Quality control of k-clique count
In this section, we prove Theorem 1.1, restated below. Let $C_{k}(G)$ be the number of cliques in a graph $G$. Let $\rho_{k}:=C_{k}(G)/\mathbb{E}_{G^{\prime}\sim\mathcal{G}_{n,p}}\left[C_{k}(G^{\prime})\right]$.
Theorem 3.1 (Theorem 1.1).
There exist constants $c_{1},c_{2},c_{3}$ such that for every constant $k$ and $p\geq\Omega\left(n^{-2/(k-1)}\right)$, the quality control problem $(\mathcal{G}_{n,p},\rho_{k})$ corresponding to estimating the number of $k$-cliques in a graph $G$ supposedly drawn according to $\mathcal{G}_{n,p}$ satisfies the following:
  1. 1.
    The problem is solvable in $O(p^{-c_{1}k})$ adjacency matrix queries to $G$.
  2. 2.
    When $p=\Omega(\log(n)/n)^{1/(4k-1)}$, the problem is additionally solvable in $O(p^{-c_{3}k})$ time.
  3. 3.
    Furthermore, no algorithm solves this problem with $o(p^{-c_{2}k})$ queries.
We typically assume $\varepsilon$ is a constant, and therefore omit the explicit dependence on $\varepsilon$ in the theorem statements. However, the upper bound on the query complexity with the dependence on $\varepsilon$ is $O(p^{-c_{1}k}\cdot\varepsilon^{-4})$, and the runtime is $O(p^{-c_{3}k})\cdot\exp(\text{poly}(k/\varepsilon))$.
For $k$-cliques, we are interested in the asymptotic exponent of $1/p$ in the query complexity and runtime. However, for triangles, we furthermore achieve a tight bound on the exponent, obtaining a $\Omega(1/p)$ lower bound and $\widetilde{O}(1/p)$ upper bound on the query complexity and runtime (with adjacency list, adjacency matrix, and degree queries). See Theorem A.1 in Appendix A. It remains an interesting open question to identify the precise exponent in the case of $k$-cliques, for growing $k$.
3.1 Query complexity upper bound
In this section, we prove Theorem 3.1 Part (1).
For $\ell\in[k]$, define $s_{\ell}$ as:
$s_{\ell}=\frac{\ell^{2}\cdot 8192}{p^{2\ell-2}\cdot\varepsilon^{2}}\cdot\left(\frac{3}{2}\right)^{2(k-\ell)}\cdot\ln\left(\frac{(6\ell)^{\ell}}{p^{\binom{\ell+2}{2}}\cdot\varepsilon}\right).$ (4)
Define sample size $s_{*}$ as:
$s_{*}=s_{k}.$ (5)
Additionally, for $s\in\mathbb{N}$ and $\ell\in[k]$, let
$A_{s,\ell}=\binom{s}{\ell}\cdot\ell!\cdot\frac{1}{n^{\ell}}$ (6)
and
$E_{s,\ell}=\binom{n}{\ell}\cdot\ell!\cdot p^{\binom{\ell}{2}}\cdot A_{s,\ell}.$ (7)
$E_{s,\ell}$ is the expected number of labeled $\ell$-cliques in a multiset of vertices (where each element is drawn uniformly from the vertex set) when the graph is drawn from $\mathcal{G}_{n,p}$ (see Claim 3.1.1). Particularly, since we are considering the number of $\ell$-cliques in a multiset, if a vertex appears $c$ times in the multiset, the $\ell$-cliques it is a part of are counted $c$ times.
We consider the following algorithm for quality control of the $k$-clique count compared to $\mathcal{G}_{n,p}$.
Algorithm Clique-Quality:
On inputs $G$, $n,p,\varepsilon$, $\mathcal{G}_{n,p}$, and $k$
Sample $s_{*}$ vertices, each uniformly at random (with repetition). Let $S$ be the corresponding multiset of vertices. For $\ell$ from $2$ to $k$, count the number of $\ell$-cliques from $S$ in $G$ (i.e., $C_{\ell}(S)$ from Section 3.1). If, for any $\ell$, $C_{\ell}(S)$ is not in $(1\pm\varepsilon/2)E_{s_{*},\ell}$, reject. Otherwise, accept.
We now define exponential quasirandomness, a property of graphs that implies concentration of the number of $\ell$-cliques around the expected amount in $G_{s,p}$, in most (multi-sets) $S$ of $\approx s$ vertices. We first set up some notation.
In what follows, let $C_{\ell}(G)$ be the number of labeled $\ell$-cliques in $G$. For a multiset $S$ of size $s$, define $C_{\ell}(S)$ as follows.
Definition 3.1.
Let $S$ be a multiset of size $s$. $C_{\ell}(S)$ is defined to be the number of labeled $\ell$-cliques in $S$. Since $S$ is a multiset, $\ell$-cliques are counted as follows: if a vertex appears $c$ times in the multiset, the $\ell$-cliques it is a part of are counted $c$ times.
Definition 3.1 (Exponentially robustly quasirandom (see Section 5.1 for general definition)).
We say that a graph $G=([n],E)$ is $(\varepsilon,\alpha,s_{0})$-exponentially robustly quasirandom with respect to the count of $\ell$-cliques if for every $s\geq s_{0}$ we have that
$\Pr_{X_{1},\ldots,X_{s}\sim_{\mathrm{i.i.d.}}[n]}\left[C_{\ell}(\{X_{1},\ldots,X_{s}\})\not\in(1\pm\varepsilon)\binom{n}{\ell}p^{\binom{\ell}{2}}\cdot\binom{s}{\ell}\frac{\ell!}{n^{\ell}}\right]\leq 4^{\ell+1}\cdot\exp(-\alpha s).$
In what follows, we will specify an $\varepsilon,\alpha$ pair for each $\ell$, setting $\alpha_{\ell}=O(p^{2(\ell-1)})$.
Definition 3.1.
For $2\leq\ell\leq k-1$, define
$\varepsilon_{\ell}=\varepsilon\cdot\left(\frac{2}{3}\right)^{k-\ell-1}.$
Define $\alpha_{2}=\varepsilon^{2}p^{2}\cdot\left(\frac{2}{3}\right)^{2(k-3)}/128$, and for $\ell\in\{3,4,\dots,k\}$ define
$\alpha_{\ell}=\frac{p^{2\ell-2}\cdot\varepsilon_{\ell-1}^{2}}{4096\cdot\ell^{2}}.$
Observe that, for $s_{B}\geq s_{A}$, $(\varepsilon,\alpha,s_{A})$-exponentially robust quasirandomness with respect to the count of $\ell$-cliques implies $(\varepsilon,\alpha,s_{B})$-exponentially robust quasirandomness.
The key ingredient in the analysis of the algorithm is the following inductive lemma (given qualitatively in Lemma 1.6), which states that a graph $G$ is $(\varepsilon_{\ell-1},\alpha_{\ell-1},s_{\ell})$-exponentially robustly quasirandom with respect to $(\ell-1)$-clique counts, then it can be tested for being $(\varepsilon_{\ell},\alpha_{\ell},s_{\ell})$-exponentially robustly quasirandom with respect to $\ell$-clique counts in $\widetilde{O}\left(p^{-4\ell+4}\right)$ queries to the adjacency matrix.
Lemma 3.2 (Inductive Lemma).
Suppose that the graph $G$ is $(\varepsilon_{\ell-1},\alpha_{\ell-1},s_{\ell})$-exponentially robustly quasirandom with respect to the count of $(\ell-1)$-cliques.
Define the tester $T_{\ell}$ that samples $s_{\ell}$ (as specified in Equation 4) random vertices of $G$ (with replacement) and accepts if and only if the $\ell$-clique count in the induced subgraph on the sampled vertices is within $(1\pm\varepsilon_{\ell-1})\binom{n}{\ell}\ell!\cdot p^{\binom{\ell}{2}}\cdot\binom{s_{\ell}}{\ell}\frac{\ell!}{n^{\ell}}$. This algorithm uses $\widetilde{O}\left(p^{-4\ell+4}\varepsilon^{-4}\right)$ adjacency matrix queries.
With probability at least $1-4^{\ell+1}\exp(-\alpha_{\ell}s_{\ell})$, the algorithm accepts graphs that are $(\varepsilon_{\ell-1},\alpha_{\ell},s_{\ell})$-exponentially robustly quasirandom with respect to the $\ell$-clique count and rejects graphs that are not $(\varepsilon_{\ell},\alpha_{\ell},s_{\ell})$-exponentially robustly quasirandom with respect to the $\ell$-clique count.
The induction lemma is (“easily”) proven via the following concentration lemma (given qualitatively in Lemma 1.7):
Lemma 3.3 (Concentration Lemma (see Lemma 5.3 for general statement)).
Suppose $G$ is $(\gamma,\beta,s)$-exponentially robustly quasirandom with respect to the count of $(\ell-1)$-cliques. Suppose $s$ satisfies $\exp(-\beta s)\leq p^{\binom{\ell+2}{2}}$.
Then:
$\Pr_{S}\left(\left|C_{\ell}(S)-C_{\ell}(G)\cdot A_{s,\ell}\right|\geq\frac{\gamma}{4}\binom{n}{\ell}p^{\binom{\ell}{2}}\cdot\binom{s}{\ell}\frac{\ell!}{n^{\ell}}\right)$
$\leq 2\cdot 4^{\ell}\exp(-\beta s)+2\exp\left(-\gamma^{2}s\cdot\frac{p^{2\ell-2}}{2048}\right).$
For our application to checking $k$-clique counts, we will repeatedly apply this concentration lemma with the parameters $\varepsilon_{\ell}$ and $\alpha_{\ell}$ from Section 3.1. Therefore, for ease of use, we provide the concentration lemma with the specific parameters as a corollary.
Lemma 3.4 (Concentration Lemma, instantiated).
For $2\leq\ell\leq k-1$, consider the definitions of $\varepsilon_{\ell}$ and $\alpha_{\ell}$ from Section 3.1 and the definition of $s_{\ell}$ from Equation 4.
Suppose $G$ is $(\varepsilon_{\ell-1},\alpha_{\ell-1},s_{\ell})$-exponentially robustly quasirandom with respect to $(\ell-1)$-clique counts. Then:
$\Pr_{S}\left(\left|C_{\ell}(S)-C_{\ell}(G)\cdot A_{s_{\ell},\ell}\right|\geq\frac{\varepsilon_{\ell-1}}{4}\binom{n}{\ell}p^{\binom{\ell}{2}}\cdot\binom{s_{\ell}}{\ell}\frac{\ell!}{n^{\ell}}\right)\leq 4^{\ell+1}\exp(-\alpha_{\ell}s_{\ell}).$
To see how to prove Lemma 3.2 from Lemma 3.3, observe that the concentration lemma tells us that if a graph is exponentially robustly quasirandom with respect to $(\ell-1)$-cliques, then with high probability a sampled multiset $S$ of size at least $s_{\ell}$ will have $C_{\ell}(S)$ that is proportional to $C_{\ell}(G)$. We can therefore use the number of $\ell$-cliques in $S$ to determine if $C_{\ell}(G)$ corresponds to the $\mathcal{G}_{n,p}$ expected amount or not. If it does, then the concentration lemma exponential quasirandomness for $\ell$-cliques. If it does not, we can efficiently reject the input graph.
3.1.1 Proof of Concentration Lemma
In this section, we prove Lemma 3.3.
We first prove the following two claims. Let $C_{\ell}(G)$ be the number of $\ell$-cliques in the graph $G$. For a multiset $S$, define $C_{\ell}(S)$ to be the number of labeled $\ell$-cliques in $S$, with repetitions for repeated vertices, as defined in Section 3.1.
Claim 3.4.
Let $S$ be a multiset generated by sampling $s$ vertices from the graph, with replacement. Then $\mathbb{E}_{S}\left[C_{\ell}(S)\right]=C_{\ell}(G)\cdot A_{s,\ell}.$
Proof.
Let us first write the number of $\ell$-cliques in $S$ (denoted as $C_{\ell}(S)$) as a sum of indicator random variables. For ordered set $T$ and ordered $\ell$-clique $C$, let $T\equiv C$ mean that the ordered set of vertices of $T$ is exactly that of $C$.
$C_{\ell}(S)=\sum_{C\in C_{\ell}(G)}\sum_{T\in S^{\ell}}1(T\equiv C).$
Fix an ordered tuple of $\ell$ vertices in $S$, and some labeled $\ell$-clique $C\in C_{\ell}(G)$. The probability that $T\equiv C$ is $1/n^{\ell}$, since $T$ is created by sampling $\ell$ vertices, each time uniformly at random (with repetitions).
By linearity of expectation, $\mathbb{E}\left(C_{\ell}(S)\right)=C_{\ell}(G)\cdot\binom{s}{\ell}\ell!/n^{\ell}.$
In the following claim, let $E_{s,\ell-1}$ be the expected number of $(\ell-1)$-cliques in a size-$s$ multiset of vertices in a graph drawn from $\mathcal{G}_{n,p}$, as defined in Equation 7.
Claim 3.4.
Consider any two sets $S,S^{\prime}$ such that the number of $(\ell-1)$-cliques in each of $S$ and $S^{\prime}$ is in the range $C_{\ell-1}(S),C_{\ell-1}(S^{\prime})\in(1\pm\varepsilon)E_{s,\ell-1}$. Let $f_{\ell}(S)$ be the number of labeled $\ell$-cliques in $S$ (and similarly for $f_{\ell}(S^{\prime})$). Then,
$\left|f_{\ell}(S)-f_{\ell}(S^{\prime})\right|\leq|S\triangle S^{\prime}|\cdot\ell\cdot(1+\varepsilon)E_{s,\ell-1}.$
Proof.
Let $C_{\ell}(v,S)$ be the number of $\ell$-cliques in $S$ that involve vertex $v\in S$ (and define such a quantity for $S^{\prime}$ similarly). The only $\ell$-cliques that contribute to the difference $\left|f_{\ell}(S)-f_{\ell}(S^{\prime})\right|$ are those involving vertices $u\in S\setminus S^{\prime}$ or vertices $v\in S^{\prime}\setminus S$. For any vertex $u\in S$ (respectively $S^{\prime}$), the number of $\ell$-cliques it is involved in is upper-bounded by $\ell\cdot f_{\ell-1}(S)$, which is the number of $(\ell-1)$-cliques in $S$, times $\ell$ (since we are counting labeled copies). This yields the following:
$\left|f_{\ell}(S)-f_{\ell}(S^{\prime})\right|\leq\sum_{u\in S\setminus S^{\prime}}C_{\ell}(u,S)+\sum_{v\in S^{\prime}\setminus S}C_{\ell}(v,S^{\prime})$
$\leq|S\setminus S^{\prime}|\cdot\ell\cdot f_{\ell-1}(S)+|S^{\prime}\setminus S|\cdot\ell\cdot f_{\ell-1}(S^{\prime})\leq|S\triangle S^{\prime}|\cdot\ell\cdot(1+\varepsilon)E_{s,\ell-1}.$
Before proving Lemma 3.3, we make a brief observation about the values of $\alpha_{\ell}$ for $1\leq\ell\leq k$.
Observation 3.4.
For $p\leq 1/2$, for all $\ell\in\mathbb{N}$, $\alpha_{\ell+1}\leq\alpha_{\ell}$.
Proof of Lemma 3.3.
Suppose that $G$ is $(\gamma,\beta,s)$-exponentially robustly quasirandom with respect to $(\ell-1)$-clique counts. Let $X_{1},\dots,X_{s}$ be the random variables corresponding to which vertices are selected to be in $S$ (i.e., $X_{i}$ is the $i$-th vertex added to $S$, with repetitions). Define $f_{\ell-1}:[n]^{s}\to\mathbb{R}$ to be the number of $(\ell-1)$-cliques in the multiset. Let $E_{s,\ell-1}$ and $E_{s,\ell}$ be the expected number of $(\ell-1)$-cliques and $\ell$-cliques, respectively, in a size-$s$ multiset for a graph drawn from $\mathcal{G}_{n,p}$, as defined in Equation 7. By Section 3.1 of exponential quasirandomness, for every $s$ we have that
$\Pr\left[f_{\ell-1}(X_{1},X_{2},\dots,X_{s})\not\in(1\pm\gamma)E_{s,\ell-1}\right]\leq 4^{\ell}\exp(-\beta s).$
We will apply McDiarmid’s Inequality when differences are bounded with high probability (Section 2.4) to the function $f_{\ell}:[n]^{s}\to\mathbb{R}$, which is the number of $\ell$-cliques in the multiset.
Let us set up the relevant definitions for the inequality. First, let the domain be $n]^{s}$, and define $\mathcal{Y}\subseteq[n]^{s}$ to be the multisets of the domain such that the number of $(\ell-1)$-cliques in $\mathcal{Y}$ is in $(1\pm\gamma)E_{s,\ell-1}$. Then, observe that $q:=1-\mathbb{P}\left((X_{1},X_{2},\dots,X_{s})\in\mathcal{Y}\right)\leq 4^{\ell}\exp(-\beta s)$.
We argue about bounded differences of $f_{\ell}$ on $\mathcal{Y}$. For each $i\in[s]$, define $c_{i}=\ell\cdot(1+\gamma)E_{s,\ell-1}$. By Claim 3.1.1, for all $S=(x_{1},x_{2},\dots,x_{s})\in\mathcal{Y}$ and $S^{\prime}=(y_{1},y_{2},\dots,y_{s})\in\mathcal{Y}$,
$\left|f_{\ell}(S)-f_{\ell}(S^{\prime})\right|\leq|S\triangle S^{\prime}|\cdot\ell\cdot(1+\gamma)E_{s,\ell-1}=\sum_{i:x_{i}\neq y_{i}}c_{i},$
so the bounded difference property holds with this setting of $c_{i}$’s.
We want to apply Section 2.4 with error bound $\delta=\gamma E_{s,\ell}/4$. To do so, we need to verify that
$q\leq\min\left\{\frac{\delta}{2\max(f)},\frac{\delta}{4\mathbb{E}(f(X_{1},X_{2},\dots,X_{t}))},\frac{1}{2}\right\}.$
Using $\delta=\gamma E_{s,\ell}/4$, and since $\max(f)\leq\binom{s}{\ell}\ell$! and $\mathbb{E}(f(X_{1},X_{2},\dots,X_{t}))\leq\binom{s}{\ell}\ell$!, it suffices to show that:
$q\leq\min\left\{\frac{\gamma E_{s,\ell}}{8\binom{s}{\ell}\ell!},\frac{\gamma E_{s,\ell}}{16\binom{s}{\ell}\ell!},\frac{1}{2}\right\}=\min\left\{\frac{\gamma}{16}\cdot\binom{n}{\ell}p^{\binom{\ell}{2}}\cdot\frac{1}{n^{\ell}},\frac{1}{2}\right\}.$ (8)
For our setting of $q$, we now show this holds. First, observe that, from Observation 3.1.1, $q\leq 4^{\ell}\exp\left(-\beta s\right)\leq 4^{\ell}\exp\left(-\alpha_{k-1}s\right)$. Next, since
$s\geq\frac{1}{\alpha_{k-1}}\cdot\ln\left(\frac{(3k)^{k}}{p^{\binom{k+2}{2}}\cdot\varepsilon\cdot 2^{k}}\right),$
we have that $\exp\left(-\alpha_{k-1}s\right)<p^{\binom{k+2}{2}}\cdot\varepsilon\cdot 2^{k}/(3k)^{k}$ and Equation 8 holds.
We are ready to apply Section 2.4, which implies that:
$\mathbb{P}\left(\left|f_{\ell}(X_{1},X_{2},\dots,X_{s})-\mathbb{E}\left(f_{\ell}(X_{1},X_{2},\dots,X_{s})\right)\right|\geq\gamma E_{s,\ell}/4\right)$
$\leq 2q+2\exp\left(-\frac{2\max\left(0,\gamma E_{s,\ell}/4-q\sum_{i=1}^{s}c_{i}\right)^{2}}{\sum_{i=1}^{s}c_{i}^{2}}\right).$ (9)
Let us analyze the second term. First, $q\sum_{i=1}^{s}c_{i}\leq\exp(-\alpha_{\ell-1}s)\cdot s\cdot(1+\gamma)E_{s,\ell-1}$. This is significantly smaller than $\gamma E_{s,\ell}/8$, and so $2\max\left(0,\gamma E_{s,\ell}/4-q\sum_{i=1}^{s}c_{i}\right)^{2}\geq 2\gamma^{2}E_{s,\ell}^{2}/64=\gamma^{2}E_{s,\ell}^{2}/32$.
Therefore, (also rounding $(1+\gamma)E_{s,\ell}$ up to $2E_{s,\ell}$ in the denominator of the expression) Equation 9 is upper-bounded by:
$\leq 2q+2\exp\left(-\frac{\gamma^{2}E_{s,\ell}^{2}}{32\cdot s\cdot(2\ell)^{2}E_{s,\ell-1}^{2}}\right).$
By definition of $E_{s,\ell-1}$ and $E_{s,\ell}$ (Equation 7), the expression above is at most:
$\leq 2q+2\exp\left(-\frac{\gamma^{2}}{128\cdot\ell^{2}\cdot s}\cdot\left((n-\ell+1)\cdot p^{\ell-1}\cdot\frac{s-\ell+1}{n}\right)^{2}\right).$
For large enough $n$, and by definition of $s$ (Equation 5), we can use that $(n-\ell+1)/n\geq 1/2$ and $(s-\ell+1)\geq s/2$ to obtain that the expression is bounded above by:
$\leq 2q+2\exp\left(-\frac{\gamma^{2}}{128\cdot\ell^{2}\cdot s}\cdot\frac{p^{2\ell-2}s^{2}}{16}\right).$
By definition of $q$, this is:
$\leq 2\cdot 4^{\ell}\exp(-\beta s)+2\exp\left(-\gamma^{2}s\cdot\frac{p^{2\ell-2}}{2048\cdot\ell^{2}}\right).$
Proof of Lemma 3.4 given Lemma 3.3.
Consider the definitions of $\varepsilon_{\ell},\alpha_{\ell}$ from Section 3.1, and the definition of $s_{\ell}$ from Equation 4. Suppose $G$ is $(\varepsilon_{\ell-1},\alpha_{\ell-1},s_{\ell})$-exponentially robustly quasirandom with respect to $(\ell-1)$-clique counts. Observe that $s_{\ell}$ satisfies $\exp(-\alpha_{\ell-1}s_{\ell})\leq p^{\binom{\ell+2}{2}}$. The above tells us that
$\Pr_{S,|S|=s_{\ell}}\left(\left|C_{\ell}(S)-C_{\ell}(G)\cdot A_{s_{\ell},\ell}\right|\geq\frac{\varepsilon_{\ell-1}}{4}E_{s_{\ell},\ell}\right)$
$\leq 2\cdot 4^{\ell}\exp(-\alpha_{\ell-1}s_{\ell})+2\exp\left(-\varepsilon_{\ell-1}^{2}s_{\ell}\cdot\frac{p^{2\ell-2}}{2048\cdot\ell^{2}}\right).$
By the definition of $\alpha_{\ell}$ (Section 3.1) and Observation 3.1.1, this is:
$\leq 2\cdot 4^{\ell}\exp(-\alpha_{\ell-1}s_{\ell})+2\exp(-2\alpha_{\ell}s_{\ell})\leq(2\cdot 4^{\ell}+2)\exp(-2\alpha_{\ell}s_{\ell})\leq 4^{\ell+1}\exp(-\alpha_{\ell}s_{\ell}).$
3.1.2 Proof of Inductive Lemma given Concentration Lemma
In this section, we prove the inductive lemma (Lemma 3.2) from the concentration lemma (Lemma 3.3). Towards doing so, we first present some corollaries of Lemma 3.3.
The following corollary proves that if $G$ is $(\varepsilon_{\ell-1},\alpha_{\ell-1},s_{\ell})$-exponentially robustly quasirandom with respect to $(\ell-1)$-clique counts, a size-$s_{\ell}$ multiset chosen uniformly at random will, with very high probability, reveal what the global number of $\ell$-cliques in $G$ is and whether $G$ is $(\varepsilon_{\ell},\alpha_{\ell},s_{\ell})$-exponentially robustly quasirandom with respect to $\ell$-clique counts.
Corollary 3.4 (Corollary of Lemma 3.4).
Consider $2\leq\ell\leq k-1$. If $G$ is $(\varepsilon_{\ell-1},\alpha_{\ell-1},s_{\ell})$-exponentially robustly quasirandom with respect to $(\ell-1)$-clique counts, then the following hold:
  1. 1.
    Suppose that $G$ is additionally $(\varepsilon_{\ell-1},\alpha_{\ell},s_{\ell})$-exponentially robustly quasirandom with respect to $\ell$-clique counts. Sample a multiset $S$ by selecting $s_{\ell}$ vertices uniformly at random (with repetition). Then with probability at least $1-4^{\ell+1}\exp(-\alpha_{\ell}s_{\ell})$ (over the choice of $S$), $C_{\ell}(S)\in(1\pm\varepsilon_{\ell-1})E_{s_{\ell},\ell}$.
    This holds for any choice of parameters $\varepsilon_{\ell-1},\alpha_{\ell},s_{\ell}$ and is not specific to their definitions in Section 3.1 and Equation 4.
  2. 2.
    Sample a multiset $S$ by selecting $s_{\ell}$ vertices uniformly at random (with repetition). If $C_{\ell}(S)\in(1\pm\varepsilon_{\ell-1})E_{s_{\ell},\ell}$, then with probability at least $1-4^{\ell+1}\exp(-\alpha_{\ell}s_{\ell})$ (over the choice of $S$), $C_{\ell}(G)\in(1\pm\frac{5}{4}\varepsilon_{\ell-1})\binom{n}{\ell}\ell!\cdot p^{\binom{\ell}{2}}$ and, furthermore, $G$ is $(\varepsilon_{\ell},\alpha_{\ell},s_{\ell})$-exponentially robustly quasirandom with respect to $\ell$-clique counts.
Proof of Section 3.1.2.
Suppose that $G$ is $(\varepsilon_{\ell-1},\alpha_{\ell-1},s_{\ell})$-exponentially robustly quasirandom with respect to $(\ell-1)$-clique counts. Then, by Lemma 3.4,
$\Pr_{S}\left(\left|C_{\ell}(S)-C_{\ell}(G)\cdot A_{s_{\ell},\ell}\right|\geq\frac{\varepsilon_{\ell-1}}{4}E_{s_{\ell},\ell}\right)\leq 4^{\ell+1}\exp(-\alpha_{\ell}s_{\ell}).$
First, let’s address case (1) of the corollary. Suppose $G$ is $(\varepsilon_{\ell},\alpha_{\ell},s_{\ell})$-exponentially robustly quasirandom with respect to $\ell$-clique counts. By definition of exponential quasirandomness (Section 3.1), all but a $4^{\ell+1}\exp(-\alpha_{\ell}s_{\ell})$ fraction of size-$s$ multisets $S$ satisfy
$C_{\ell}(S)\in\left(1\pm\varepsilon_{\ell-1}\right)E_{s_{\ell},\ell}.$
Let’s now move to case (2) of the corollary. Lemma 3.4 implies that when $G$ is $(\varepsilon_{\ell-1},\alpha_{\ell-1},s_{\ell})$-exponentially robustly quasirandom with respect to $(\ell-1)$-clique counts, with probability at least $1-4^{\ell+1}\exp\left(-\alpha_{\ell}s_{\ell}\right)$,
$C_{\ell}(G)\cdot A_{s_{\ell},\ell}\in C_{\ell}(S)\pm\frac{\varepsilon_{\ell-1}}{4}E_{s_{\ell},\ell}.$
Equivalently, with probability at least $1-4^{\ell+1}\exp\left(-\alpha_{\ell}s_{\ell}\right)$,
$C_{\ell}(G)\in C_{\ell}(S)\cdot\frac{1}{A_{s_{\ell},\ell}}\pm\frac{\varepsilon_{\ell-1}}{4}\binom{n}{\ell}\ell!\cdot p^{\binom{\ell}{2}}$. (10)
Additionally, by Lemma 3.4, under the high-probability event,
$C_{\ell}(S)\in C_{\ell}(G)\cdot A_{s_{\ell},\ell}\pm\frac{\varepsilon_{\ell-1}}{4}E_{s_{\ell},\ell}.$ (11)
Suppose that $C_{\ell}(S)\in(1\pm\varepsilon_{\ell-1})\binom{n}{\ell}\ell!\cdot p^{\binom{\ell}{2}}\cdot\binom{s_{\ell}}{\ell}\frac{\ell!}{n^{\ell}}$. Therefore, $C_{\ell}(S)\cdot\frac{n^{\ell}}{\binom{s_{\ell}}{\ell}}\in(1\pm\varepsilon_{\ell-1})\binom{n}{\ell}\ell!\cdot p^{\binom{\ell}{2}}$. By Equation 10, with probability at least $1-4^{\ell+1}\exp\left(-\alpha_{\ell}s_{\ell}\right)$, $C_{\ell}(G)\in(1\pm\frac{5}{4}\varepsilon_{\ell-1})\binom{n}{\ell}\ell!\cdot p^{\binom{\ell}{2}}$. We now argue that the graph is exponentially robustly quasirandom with high probability. By Equation 11, all but a $4^{\ell+1}\exp(-\alpha_{\ell}s_{\ell})$ fraction of size-$s_{\ell}$ multisets $S^{\prime}$ satisfy
$C_{\ell}(S^{\prime})\in\left(1\pm\frac{3}{2}\varepsilon_{\ell-1}\right)\binom{n}{\ell}\ell!\cdot p^{\binom{\ell}{2}}\cdot\binom{s_{\ell}}{\ell}\frac{\ell!}{n^{\ell}}.$
Since $\varepsilon_{\ell}=\frac{3}{2}\varepsilon_{\ell-1}$, therefore $G$ is $(\varepsilon_{\ell},\alpha_{\ell},s_{\ell})$-exponentially robustly quasirandom with respect to $\ell$-clique counts. ∎
Proof of Lemma 3.2.
Given Section 3.1.2, Lemma 3.2 follows from a simple argument, given below.
Suppose that $G$ is $(\varepsilon_{\ell-1},\alpha_{\ell-1},s_{\ell})$-exponentially robustly quasirandom with respect to $(\ell-1)$-clique counts. Define the tester $T_{\ell}$ that samples $s_{\ell}$ (as specified in Equation 4) random vertices of $G$ (with replacement) and accepts if and only if the labeled $\ell$-clique count in the induced subgraph on the sampled vertices is within $(1\pm\varepsilon_{\ell-1})E_{s_{\ell},\ell}$.
By the corollary (Section 3.1.2) of the concentration lemma (Lemma 3.3), the tester $T_{\ell}$ satisfies the following properties.
Completeness: Suppose that $G$ is $(\varepsilon_{\ell-1},\alpha_{\ell},s_{\ell})$-exponentially robustly quasirandom with respect to $\ell$-cliques. Then, with probability at least $1-4^{\ell+1}\exp(-\alpha_{\ell}s_{\ell})$, $C_{\ell}(S)\in(1\pm\varepsilon_{\ell-1})E_{s_{\ell},\ell}$.
Soundness: If the clique count on the sampled multiset is in $(1\pm\varepsilon_{\ell-1})\binom{n}{\ell}\ell!\cdot p^{\binom{\ell}{2}}\cdot\binom{s_{\ell}}{\ell}\frac{\ell!}{n^{\ell}}$, then with probability at least $1-4^{\ell+1}\exp(-\alpha_{\ell}s_{\ell})$, $G$ is $(\varepsilon_{\ell},\alpha_{\ell},s_{\ell})$-exponentially robustly quasirandom with high probability, so accepted graphs satisfy the quasirandomness condition. ∎
3.1.3 Proof of Theorem 3.1
Finally, we use the reasoning of Lemma 3.2 to argue about the correctness of the algorithm and the proof of Theorem 3.1. Let $s=s_{k}$ as in Equation 5, and observe that $s\geq s_{\ell}$ for all $\ell\in[k]$.
Observation 3.4.
Any input graph $G$ is $(0,\infty,s)$-exponentially robustly quasirandom with respect to the count of vertices (1-cliques), for any $s\in\mathbb{N}$.
Proof.
Since we are counting vertices in the multiset with repetitions, the count of vertices in the multiset sampled is always $s$. ∎
We prove another corollary of Lemma 3.3 which, as opposed to Section 3.1.2, separates the case of $C_{k}(G)\not\in(1\pm\varepsilon)\binom{n}{k}p^{\binom{k}{2}}$ from the typical $C_{k}(G)$ range of $G\sim\mathcal{G}_{n,p}$ (Section 2.3). We remark that Section 3.1.2 was suitable for testing quasirandomness for $\ell\leq k-1$, since the algorithm needed to handle any possible value of $C_{\ell}(S)$ on sample $S$ for $\ell\leq k-1$, and weaker quasirandomness parameters were necessary for the induction. However, once we are finally ready to test the number of $k$-cliques in the graph, given approximate guarantees about the number of $\ell$-cliques in the graph for $\ell\leq k-1$, we are dealing with a promise problem where either $G\sim\mathcal{G}_{n,p}$ or $C_{k}(G)\not\in(1\pm\varepsilon)\binom{n}{k}p^{\binom{k}{2}}$, which is reflected in the following corollary.
Corollary 3.4 (Corollary of Lemma 3.3).
Suppose $G$ is $(\varepsilon,\alpha_{k-1},s)$-exponentially robustly quasirandom with respect to (k1)(k-1)-clique counts, and $C_{k}(G)\in(1\pm\frac{1}{4}\varepsilon)\binom{n}{k}k!\cdot p^{\binom{k}{2}}$. Sample a multiset $S$ by selecting $s$ vertices uniformly at random (with repetition). Then with probability at least $1-4^{k+1}\exp(-\alpha_{k}s)$, $C_{k}(S)\in(1\pm\frac{1}{2}\varepsilon)E_{s,k}$. If $C_{k}(G)\not\in(1\pm\varepsilon)\binom{n}{k}k!\cdot p^{\binom{k}{2}}$, then with probability at least $1-4^{k+1}\exp(-\alpha_{k}s)$, $C_{k}(S)\not\in(1\pm\frac{3}{4}\varepsilon)E_{s,k}$.
Proof.
The proof follows similarly to that of Section 3.1.2. ∎
For the completeness argument of Theorem 3.1, we need to argue that global concentration of clique counts around the mean allows us to conclude about local concentration – that is, exponential quasirandomness. This is given in the following claim.
Claim 3.4.
Suppose that a graph $G$ satisfies the property that for all $\ell\in\{2,3,\dots,k\}$ the number of labeled $\ell$-cliques in $G$ is in the range $(1\pm\frac{\varepsilon}{4(k+1)})\binom{n}{\ell}\ell!\cdot p^{\binom{\ell}{2}}$ (for $p=\Omega\left(\varepsilon^{-2}n^{-2/(k-1)}\right)$). Then, for size $s_{\ell}$ multisets, for all $\ell\in[k-1]$, $G$ is $\left((\ell+1)\cdot\frac{\varepsilon}{4(k+1)},(\ell+1)^{2}\cdot\frac{\varepsilon^{2}}{(k+1)^{2}}\cdot\frac{p^{2\ell-2}}{2048},s_{\ell}\right)$-exponentially robustly quasirandom with respect to $\ell$-clique counts.
Proof.
We prove this statement by induction on $\ell$, using a repeated application of Lemma 3.3.
Base case: Consider the case of $\ell=1$. By Observation 3.1.3, any graph $G$ is $\left(0,\infty,s_{1}\right)$-exponentially robustly quasirandom with respect to the vertex count.
Inductive hypothesis: Suppose that $G$ is $(\gamma,\beta,s_{\ell-1})$-exponentially robustly quasirandom with respect to $(\ell-1)$-clique counts, for $\gamma=\ell\cdot\frac{\varepsilon}{4(k+1)}$ and $\beta=\ell^{2}\cdot\frac{\varepsilon^{2}}{(k+1)^{2}}\cdot\frac{p^{2\ell-2}}{2048}$.
Inductive argument: Assume the inductive hypothesis. By Lemma 3.3, since $G$ is also $(\gamma,\beta,s_{\ell})$-exponentially robustly quasirandom with respect to $(\ell-1)$-clique counts,
$\Pr_{S}\left(\left|C_{\ell}(S)-\left(1\pm\frac{\varepsilon}{4(k+1)}\right)E_{s_{\ell},\ell}\right|\geq\frac{\gamma}{4}E_{s_{\ell},\ell}\right)$
$\leq 2\cdot 4^{\ell}\exp(-\beta s_{\ell})+2\exp\left(-\gamma^{2}s_{\ell}\cdot\frac{p^{2\ell-2}}{2048}\right).$
Since $\beta\geq\frac{\gamma^{2}p^{2\ell-2}}{2048}$, this is
$\leq 4^{\ell+1}\exp\left(-\gamma^{2}s_{\ell}\cdot\frac{p^{2\ell-2}}{2048}\right).$
Therefore, also using that $\frac{\varepsilon}{4(k+1)}+\frac{\gamma}{4}\leq(\ell+1)\cdot\frac{\varepsilon}{4(k+1)}$, we find that $G$ is
$\left((\ell+1)\cdot\frac{\varepsilon}{4(k+1)},(\ell+1)^{2}\cdot\frac{\varepsilon^{2}}{(k+1)^{2}}\cdot\frac{p^{2\ell-2}}{2048},s_{\ell}\right)$-exponentially robustly quasirandom with respect to $\ell$-clique counts. ∎
Proof of Theorem 3.1 – query complexity upper bound.
Completeness. By Section 2.3, with high constant probability, for $G\sim\mathcal{G}_{n,p}$ with $p=\Omega\left(\varepsilon^{-2}n^{-2/(k-1)}\right)$, for all $\ell\leq k$ the number of labeled $\ell$-cliques in $G$ is in the range $(1\pm\frac{\varepsilon}{4(k+1)})\binom{n}{\ell}\ell!\cdot p^{\binom{\ell}{2}}$. In the following, suppose that the high-probability event that the clique counts are all in these typical ranges holds. Then, by Claim 3.1.3, for all $\ell\in[k]$, $G$ is $\left(\delta_{\ell},\nu_{\ell},s_{\ell}\right)$-exponentially robustly quasirandom with respect to $\ell$-clique counts, for
$\delta_{\ell}=(\ell+1)\cdot\frac{\varepsilon}{4(k+1)},~~\nu_{\ell}=(\ell+1)^{2}\cdot\frac{\varepsilon^{2}}{(k+1)^{2}}\cdot\frac{p^{2\ell-2}}{2048}.$
By Section 3.1.2, from this exponential quasirandomness, we can conclude that, with probability at least $1-k\cdot 4^{k}\exp\left(-\nu_{k}s_{*}\right)\geq 9/10$, for all $\ell\in[k]$, the uniformly sampled multiset $S$ satisfies $C_{\ell}(S)\in(1\pm\frac{\varepsilon}{4})E_{s,\ell}$. Therefore, the algorithm accepts $G\sim\mathcal{G}_{n,p}$ with high constant probability.
Soundness. We begin by noting that, by Observation 3.1.3, every input graph $G$ is $(0,\infty,s_{1})$-exponentially robustly quasirandom with respect to the count of vertices. Next, suppose that, for some $\ell\in\{1,2,\dots,k-1\}$, the input graph $G$ is $(\varepsilon_{\ell^{\prime}},\alpha_{\ell^{\prime}},s_{\ell^{\prime}})$-exponentially robustly quasirandom with respect to the count of $\ell^{\prime}$-cliques for all $\ell^{\prime}<\ell$ and is not $(\varepsilon_{\ell},\alpha_{\ell},s_{\ell})$-exponentially robustly quasirandom with respect to the count of $\ell$-cliques. Then, by the inductive lemma (Lemma 3.2), with probability at least $1-4^{\ell+1}\exp(-\alpha_{\ell}s_{\ell})\geq 9/10$, the algorithm rejects on $G$ (on the basis of the count of $\ell$-clique counts in the sampled $S$ being too small or large).
Suppose that the input graph is $(\varepsilon_{k-1},\alpha_{k-1},s_{*})$-exponentially robustly quasirandom with respect to the count of (k1)(k-1)-cliques. Equivalently, since $\varepsilon_{k-1}=\varepsilon$ (the epsilon as input to our quality control problem), we can assume $G$ is $(\varepsilon,\alpha_{k-1},s)$-exponentially robustly quasirandom with respect to the count of (k1)(k-1)-cliques. By Section 3.1.3, with probability at least $1-4^{k+1}\exp(-\alpha_{k}s)$, $C_{k}(S)\not\in(1\pm\frac{3}{4}\varepsilon)E_{s,k}$ and the algorithm outputs reject.
Therefore, the algorithm rejects each $G$ with $C_{k}(G)\not\in(1\pm\varepsilon)\binom{n}{k}p^{\binom{k}{2}}$ with high probability.
Query complexity: The algorithm solely queries all adjacencies between vertices in a $S$ of size $s_{*}$ (as defined in Equation 5). Therefore, the algorithm makes $\binom{s_{*}}{2}=\frac{1}{p^{O(k)}}\cdot\frac{1}{\varepsilon^{4}}$ adjacency matrix queries.
Runtime: The algorithm stated counts $\ell$-cliques, for all $\ell\leq k$, for the induced subgraph on a set $S$ of $s_{*}$ vertices. For worst-case input graphs, this can take up to $O(s_{*}^{k})=\frac{1}{p^{O(k^{2})}}\cdot\frac{1}{\varepsilon^{2k}}$ time. ∎
3.2 Query complexity lower bound
We now prove Theorem 3.1 Part (3), more formally given as follows.
Theorem 3.5 (Theorem 3.1 Part (3)).
Consider parameters $n,p$ such that $p\geq\Omega\left(n^{-2/(k-1)}\right)$. Any $(\mathcal{G}_{n,p},\rho_{k})$-quality control algorithm with query access to the adjacency matrix must make at least $\Omega\left(p^{-(k-1)/2}\right)$ queries.
This theorem is a direct implication of a result we prove in the more general setting of lower bounds for $(\mathcal{G}_{n,p},\rho_{H})$-quality control for any motif $H$. We defer the proof to Section 4.2.
3.3 Algorithmic complexity
We have proven that a query complexity of $\frac{1}{p^{O(k)}}$ suffices for the quality control problem for comparing the $k$-clique count to Erdős–Rényi graphs. In this section, we enhance the algorithm from Section 3.1 by leveraging the notion of graph jumbledness [Tho87]. This allows us to obtain an algorithm with query complexity and runtime of $\frac{1}{p^{O(k)}}$, proving Theorem 3.1 Part (2).
Theorem 3.6 (Theorem for $k$-cliques).
For constant error parameter $\varepsilon$, and
$p=\Omega\left(\log(n)/n\right)^{1/(4k-1)}$, there is an algorithm for $(\mathcal{G}_{n,p},\rho_{k})$-quality control that makes $\frac{1}{p^{O(k)}}$ adjacency matrix queries to the input graph and has a runtime of $\frac{1}{p^{O(k)}}$.
We remark that the above is stated for constant $\varepsilon$ and $k$. The runtime with the explicit dependence on $\varepsilon,k$ is $\exp(\mathrm{poly}(k/\varepsilon))\cdot p^{-O(k)}$. The bound on $p$ with the explicit dependence on $\varepsilon,k$ is $p=\Omega\left(\exp(\mathrm{poly}(k/\varepsilon))\log(n)/n\right)^{1/(4k-1)}$.
We begin by defining parameters $\zeta_{\ell},\eta_{\ell}$, which play the role that $\varepsilon_{\ell},\alpha_{\ell}$ did in Section 3.1. Our efficient algorithm relies on approximate counting of the number of copies of $H$; as we move to approximate counting, some inductive dependencies between the parameters from $\ell$ to $\ell+1$ change, which is why we need to redefine these.
Definition 3.6.
For $2\leq\ell\leq k-1$, define
$\zeta_{\ell}=\varepsilon\cdot\left(\frac{4}{7}\right)^{k-\ell-1}.$
Define $\eta_{2}=\varepsilon^{2}\cdot\left(\frac{4}{7}\right)^{2(k-3)}/128$, and for $\ell\in\{3,4,\dots,k\}$ define
$\eta_{\ell}=\frac{p^{2\ell-2}\cdot\zeta_{\ell-1}^{2}}{4096}.$
We also redefine $s$, which will still be $1/p^{O(k)}$ for a larger hidden constant in the $O(\cdot)$.
Definition 3.6.
Let $C$ be a constant $C=C(\varepsilon,k)=\exp(\mathrm{poly}(k/\varepsilon))$. Define $s_{\ell}$ to be:
$s_{\ell}=\frac{1200C^{2}\cdot\ell^{2}\cdot 8192}{p^{5\ell}\cdot\varepsilon^{2}}\cdot\left(\frac{7}{4}\right)^{2(k-\ell)}\cdot\ln\left(\frac{(6\ell)^{\ell}}{p^{\binom{\ell+2}{2}}\cdot\varepsilon}\right).$
Let $s_{*}:=s_{k}$.
Recall that, as defined in Equation 7, $E_{s,\ell}$ is the expected number of (distinct labeled induced) $\ell$-cliques in a subgraph on a multiset of $s$ vertices sampled with replacement, when the underlying graph is drawn from $\mathcal{G}_{n,p}$.
Algorithm Clique-Quality-Efficient:
On inputs $G,n,p,\mathcal{G}_{n,p},\varepsilon$, and $k$
  1. 1.
    Sample $s_{*}$ vertices, each uniformly at random (with repetition). Let $S$ be the corresponding set of vertices, with induced graph $G_{S}$.
  2. 2.
    Perform adjacency matrix queries on all pairs in $S$ to find all edges.
  3. 3.
    Test that $G_{S}$ is $(p,\beta)$-jumbled for $\beta\leq\frac{p^{k}s_{*}}{C}$ for a constant (with respect to $n$) $C=\exp\left(\text{poly}(k/\varepsilon)\right)$ by checking the following: Let $\delta=O(p^{2k}/C^{2})$. For every $u\in S$, check that the degree of $u$ in $S$ is $(s_{*}-1)(p\pm\delta)$. For every $(u,v)\in S\times S$, check that the number of common neighbors of $(u,v)$ is $(s_{*}-2)(p^{2}\pm\delta)$. If any pair in $S\times S$ does not satisfy this, reject.
  4. 4.
    Else, approximate the number of $\ell$-cliques in $S$ to within a $\zeta_{\ell-1}E_{s_{*},\ell}/8$ additive factor, for all $\ell\leq k$, in time $O(s_{*}^{c})$ for a constant $c$, using the algorithm referenced in Section 2.2 (applied with error parameter $\varepsilon/k^{k}$). If, for any $\ell$, the approximation of the number of $\ell$-cliques in $G_{S}$ is not in $(1\pm\frac{5}{8}\varepsilon)E_{s_{*},\ell}$, reject. Otherwise, accept.
Observe that, by Section 2.2, the runtime of this algorithm (from Step 4) with the explicit dependence on $\varepsilon$ is $\exp(\mathrm{poly}(k/\varepsilon))\cdot p^{-O(k)}$. Also note that this algorithm counts the number of $k$-cliques in the graph on the vertices in $S$, without repetitions for multiple appearances of the same vertex in $S$. We will argue in the proofs below that, since the probability of a repeated vertex is $o(1)$, this does not impact the error probability.
We now prove Theorem 3.6. From Section 3.1, we know that it suffices to reduce the global $k$-clique counting problem to the local problem of counting $k^{\prime}$-cliques, for $k^{\prime}\leq k$, over subgraphs with $s_{*}$ (as in Section 3.3) vertices. We need to handle the inclusion of the test for jumbledness in the analysis now, as well as the shift from exact to approximate clique counting in the subgraph on $s$ vertices.
We begin by modifying Corollaries 3.1.2 and 3.1.3 to account for the latter of these points: the shift from exact to approximate counting.
Corollary 3.6 (Modification of Section 3.1.2).
Consider $2\leq\ell\leq k-1$. If $G$ is $(\zeta_{\ell-1},\eta_{\ell-1},s_{\ell})$-exponentially robustly quasirandom with respect to $(\ell-1)$-clique counts, then the following hold:
  1. 1.
    Suppose that $G$ is additionally $(\zeta_{\ell-1},\eta_{\ell},s_{\ell})$-exponentially robustly quasirandom with respect to $\ell$-clique counts. Sample a multiset $S$ by selecting $s=s_{\ell}$ vertices uniformly at random (with repetition). Consider a $(1\pm\frac{1}{8}\zeta_{\ell-1})$-approximation $\widehat{C}_{\ell}(S)$ of the number of $\ell$-cliques in $S$. Then with probability at least $1-4^{\ell+1}\exp(-\eta_{\ell}s_{\ell})$ (over the choice of $S$), $\widehat{C}_{\ell}(S)\in(1\pm\frac{9}{8}\zeta_{\ell-1})E_{s_{\ell},\ell}$.
  2. 2.
    Sample a multiset $S$ by selecting vertices uniformly at random (with repetition). Consider a $(1\pm\frac{1}{8}\zeta_{\ell-1})$-approximation $\widehat{C}_{\ell}(S)$ of the number of $\ell$-cliques in $S$. If $\widehat{C}_{\ell}(S)\in(1\pm\frac{9}{8}\zeta_{\ell-1})E_{s_{\ell},\ell}$, then with probability at least $1-4^{\ell+1}\exp(-\eta_{\ell}s_{\ell})$ (over the choice of $S$), $G$ is $(\zeta_{\ell},\eta_{\ell},s_{\ell})$-exponentially robustly quasirandom with respect to $\ell$-clique counts.
Proof.
For part (1), if $G$ is $(\zeta_{\ell-1},\eta_{\ell},s_{\ell})$-exponentially robustly quasirandom with respect to $\ell$-clique counts, then by definition, with probability at least $1-4^{\ell+1}\exp(-\eta_{\ell}s_{\ell})$ (over the choice of $S$), $C_{\ell}(S)\in(1\pm\zeta_{\ell-1})E_{s_{\ell},\ell}$. Therefore the $(1\pm\frac{1}{8}\varepsilon)$-approximation $\widehat{C}_{\ell}(S)$ of $C_{\ell}(S)$ must be in $\widehat{C}_{\ell}(S)\in(1\pm\frac{9}{8}\zeta_{\ell-1})E_{s_{\ell},\ell}$.
For part (2), if $\widehat{C}_{\ell}(S)\in(1\pm\frac{9}{8}\zeta_{\ell-1})E_{s_{\ell},\ell}$, then $C_{\ell}(S)\in(1\pm\frac{5}{4}\zeta_{\ell-1})E_{s_{\ell},\ell}$. Using the same reasoning as in Section 3.1.2, and since $\zeta_{\ell}=\frac{7}{4}\zeta_{\ell-1}$, we find that $G$ is $(\zeta_{\ell},\eta_{\ell},s_{\ell})$-exponentially robustly quasirandom with respect to $\ell$-clique counts with probability at least $1-\exp(-\eta_{\ell}s_{\ell})$. ∎
Corollary 3.6 (Implication of Section 3.1.3).
Let $\gamma>0$. Suppose $G$ is $(\varepsilon,\eta_{k-1},s_{*})$-exponentially robustly quasirandom with respect to (k1)(k-1)-clique counts, and $C_{k}(G)\in(1\pm\frac{1}{4}\varepsilon)\binom{n}{k}p^{\binom{k}{2}}$. Sample a multiset $S$ by selecting $s_{*}$ vertices uniformly at random (with repetition). Consider a $(1\pm\frac{1}{8}\varepsilon)$-approximation $\widehat{C}_{k}(S)$ of the number of $k$-cliques in $S$. Then with probability at least $1-\exp(-\alpha_{k}s_{*})$, $\widehat{C}_{k}(S)\in(1\pm\frac{5}{8}\varepsilon)E_{s,k}$. If $C_{k}(G)\not\in(1\pm\varepsilon)\binom{n}{k}p^{\binom{k}{2}}$, then with probability at least $1-\exp(-\alpha_{k}s_{*})$, $\widehat{C}_{k}(S)\not\in(1\pm\frac{5}{8}\varepsilon)E_{s_{*},k}$.
Proof.
This follows from Section 3.1.3 in exactly the same way that Section 3.3 followed from Section 3.1.3. That is, once we incorporate the error terms of the approximate count of the number of $\ell$-cliques, the implication follows immediately. ∎
We are now ready to argue about the completeness and soundness of Algorithm Clique-Quality-Efficient. The analysis will incorporate the new test for jumbledness of subgraphs.
Proof of Theorem 3.6.
Completeness. We first argue that, with high probability, $G\sim\mathcal{G}_{n,p}$ is accepted by Algorithm Clique-Quality-Efficient. Consider $G\sim\mathcal{G}_{n,p}$, and choose a random subgraph of $G$ on $s_{*}$ vertices. By Corollary 2.3, for $p\geq\left(\frac{1200C^{4}\log n}{n}\right)^{1/(4k-1)}$ and $s_{*}\geq 1200C^{4}/p^{5k}$, with probability at least $4/5$, the subgraph on these $s_{*}$ vertices satisfies the conditions checked for in Step 3 with high probability, as proven in Section 2.3. Therefore, Step 3 does not reject $G\sim\mathcal{G}_{n,p}$ with high constant probability.
Next, we argue that, with probability 1o(1)1-o(1) the approximation of the number of $\ell$-cliques, for each $\ell\in[k]$, returned by the algorithm is a $(1\pm\frac{1}{8}\zeta_{\ell-1})$-approximation of $C_{\ell}(S)$. The difference is that $C_{\ell}(S)$ counts $\ell$-cliques multiple times if the vertices appear in $S$ multiple times, but the algorithm does not count multiples. We claim that with probability at least 1o(1)1-o(1), $S$ consists of $s_{*}$ distinct vertices. To see this, the probability that there exists a pair of vertices that are equivalent, by a union bound, is at most $\binom{s}{2}\frac{1}{n}$, which is $o(1)$ by the lower bound on $p$ in terms of $n$ and the setting of $s_{*}$. Therefore, the probability that $S$ consists of $s_{*}$ distinct vertices is 1o(1)1-o(1). Therefore, for each $\ell\in[k]$, the number of $\ell$-cliques in $S$, without repetitions for repeated vertices, equals the number of $\ell$-cliques with repetitions (i.e., $C_{\ell}(S)$). In this case, for each $\ell\in[k]$, the approximation output by the algorithm of the number of $\ell$-cliques in $S$, without repetitions, is a $(1\pm\frac{1}{8}\zeta_{\ell-1})$-approximation of $C_{\ell}(S)$, and we can use the assumptions of the corollaries proven above.
Next, we argued in Section 3.1 that the global problem of verifying $k$-clique counts can be reduced to a local problem. Applying the new parameters $\zeta_{\ell},\eta_{\ell}$ of this section to the concentration lemma (3.3), a parameter-specific version of the concentration lemma analogous to Lemma 3.4 holds for our new parameters, and likewise the inductive lemma (Lemma 3.2) holds for our new parameters. Therefore, we can use the reasoning of the proof of Theorem 3.1 from Section 3.1 to claim that $G\sim\mathcal{G}_{n,p}$ is $(\zeta_{\ell-1},\eta_{\ell},s_{\ell})$-exponentially robustly quasirandom for all $\ell$. Lemma 3.2 (for the new parameters, and since $s_{\ell}$ is larger than in Section 3.1) therefore implies that the algorithm accepts $G\sim\mathcal{G}_{n,p}$ with high probability. (More precisely, the error probability of Algorithm Clique-Quality-Efficient is controlled by a union bound over the probability that the algorithm does not return a good approximation of $C_{\ell}(S)$ and the error probability (from when $C_{\ell}(S)$ is not in a certain range) from the algorithm in Lemma 3.2).
Soundness. Suppose that $C_{k}(G)\not\in(1\pm\varepsilon)\binom{n}{k}p^{\binom{k}{2}}$. First, additionally suppose that the input graph is not $(p,\beta)$-jumbled. Since Section 2.2 proves that Step 3 will reject graphs that are not $(p,\beta)$-jumbled with high probability, such graphs will be rejected.
On the other hand, suppose that the input graph is not rejected by Step 3. As in the completeness case, with probability 1o(1)1-o(1), for all $\ell\in[k]$, the approximate number of $\ell$-cliques returned by the algorithm is a $(1\pm\frac{1}{8}\zeta_{\ell-1})$-approximation of $C_{\ell}(S)$. In this case, as in the proof of Theorem 3.1, since every input graph $G$ is $(0,\infty,s_{1})$-exponentially robustly quasirandom with respect to the vertex count, the inductive lemma allows us to reject graphs that are not $(\zeta_{\ell},\eta_{\ell},s_{\ell})$-quasirandom with respect to $\ell$-clique counts for any $\ell\in[k-1]$ with high probability, by approximately counting $\ell$-cliques.
Thus, if the algorithm hasn’t rejected when it starts counting $k$-cliques, we can assume it is $(\zeta_{k-1}=\varepsilon,\eta_{k-1},s_{k}=s_{*})$-exponentially robustly quasirandom. By Section 3.3, with probability at least $1-4^{k+1}\exp(-\eta_{k}s)$, the estimate $\widehat{C}_{k}(S)$ of the number of $k$-cliques in $S$ satisfies $\widehat{C}_{k}(S)\not\in(1\pm\frac{5}{8}\varepsilon)E_{s_{*},k}$ and the input is rejected by the algorithm. Therefore, the algorithm rejects each $G$ with $C_{k}(G)\not\in(1\pm\varepsilon)\binom{n}{k}p^{\binom{k}{2}}$ with high probability.
Query complexity: The algorithm queries all pairs of vertices in a set $S$ of size $s_{*}$ (Section 3.3), therefore using $1/p^{O(k)}\cdot\varepsilon^{-4}$ adjacency matrix queries.
Runtime: Step 1 of the algorithm takes time $O(s_{*})$. Step 2 takes time $O(s_{*}^{2})$. Step 3 takes time $O(s_{*}^{3})$. Finally, Step 4 takes time $Cs_{*}^{c}$ for a constant $C=\exp(\text{poly}(|H|/\varepsilon))$ and a universal constant $c$ (independent of the value of $k$). Therefore, since $s_{*}=1/p^{O(k)}$, the runtime of Algorithm Clique-Quality-Efficient is $1/p^{O(k)}\cdot\exp(\text{poly}(k/\varepsilon))$. ∎
4 Quality control of general motif counts
Let $C_{H}(G)$ be the number of labeled, induced copies of a motif $H$ in the graph $G$. Let $\rho_{H}(G)=C_{H}(G)/\mathbb{E}_{G^{\prime}\sim\mathcal{G}_{n,p}}\left[C_{H^{\prime}}(G^{\prime})\right]$. Let $m_{H}:=\max_{F\subseteq H}e(F)/v(F)$. Let $m_{H}^{\star}$ be the maximum value of $m_{H^{\prime}}$ over all $H^{\prime}$ with $|H^{\prime}|=|H|$ and edges that are a superset of the edges of $H$. Recall the definition of $\sigma(H)$ from Equation 2.
In this section, we prove the Theorem 1.4, restated below.
Theorem 4.1 (Theorem 1.4).
There exist constants $c_{1},c_{2},c_{3}$ such that for every graph $H$ and
$p=\Omega\left(n^{-1/m_{H}^{\star}}\right)$, the quality control problem $(\mathcal{G}_{n,p},\rho_{H})$ corresponding to estimating the number of copies of $H$ in a graph $G$ supposedly drawn according to $\mathcal{G}_{n,p}$ satisfies the following:
  1. 1.
    The problem is solvable in $O(p^{-c_{1}\Delta(H)})$ adjacency matrix queries to $G$.
  2. 2.
    When $p$ is additionally $\Omega\left(\log(n)/n\right)^{1/(4\sigma(H)-1)}$, the problem is also solvable in $O(p^{-c_{3}\Delta(H)})$ time and queries.
  3. 3.
    Furthermore no algorithm solves this problem with $o(p^{-c_{2}\Delta(H)})$ queries.
To illustrate this bound, we give some implications of this result for specific motifs $H$:
Corollary 4.1.
The following are implied by Theorem 4.1.
  1. 1.
    Consider any cycle $H$ on $k$ vertices. $(\mathcal{G}_{n,p},\rho_{H})$-quality control takes $1/p^{\Theta(1)}$ adjacency matrix queries and time. Observe that this is independent of the size of the cycle and $n$. In comparison, approximating the number of cycles on $k$ vertices in worst-case graphs, even with access to random edge samples, takes $1/p^{O(k)}$ queries [AKK19].
  2. 2.
    Consider a $k$-clique $H$. $(\mathcal{G}_{n,p},\rho_{H})$-quality control takes $1/p^{\Theta(k)}$ adjacency matrix queries and time. In comparison, approximating the number of $k$-cliques in graphs whose $k$-clique count is approximately the expected amount in $\mathcal{G}_{n,p}$ takes $1/p^{\Theta(k^{2})}$ queries to the adjacency lists, adjacency matrix, and degrees [ERS20a].
  3. 3.
    Consider the star graph $H$ on $k$ vertices. $(\mathcal{G}_{n,p},\rho_{H})$-quality control takes $1/p^{\Theta(k)}$ adjacency matrix queries and time. [ABG+16] implies that approximating the number of $k$-stars for worst-case graphs with approximately the same amount of $k$-stars as expected in $\mathcal{G}_{n,p}$ requires $\widetilde{O}(n/p)$ queries, which is incomparable.
4.1 Query complexity upper bound
In this section, we prove Theorem 4.1 Part (1), which states the query complexity of quality control for the count of a motif $H$ compared to $\mathcal{G}_{n,p}$ is $p^{-O(\Delta(H))}\cdot\mathrm{poly}(1/\varepsilon)$, where $\Delta(H)$ is the maximum vertex degree of $H$.
We prove quality control results for comparing labeled induced motif counts to the expected amount from $\mathcal{G}_{n,p}$. By an induced motif, we mean the number of induced subgraphs in the graph that form a copy of the motif. Recall that an induced subgraph is a graph formed from a subset of the vertices of the graph and all of the edges between these vertices.
We will show how our results are implied from a deferred result (given in Section 5) that holds for $(D,\rho_{H})$-quality control for any distribution $D$ over graphs that has good concentration of motifs $H^{\prime}\subseteq H$. This broader result relies on the same principles as the proof for quality control of the $k$-clique count compared to $\mathcal{G}_{n,p}$. That is, a notion of exponential quasirandomness (see Section 5.1) is given for motif counts, and the proof iteratively proves (or, at some point, refutes) that exponential quasirandomness holds for the input graph. A concentration lemma (Lemma 5.3) and inductive lemma (Lemma 5.2) are developed for this context as well.
We now turn to setting up the notation and framework needed to apply Section 5.
For a motif $H$ on $\ell$ vertices, the expected number of induced motif counts in $G\sim\mathcal{G}_{n,p}$ is $\binom{n}{\ell}\ell!\cdot p^{e(H)}(1-p)^{\binom{\ell}{2}-e(H)}$.
Notation.
Recall the following notation from Section 5.1.
For a motif $M$ on $\ell$ vertices, as in Equation 14, let $F_{\mathcal{D}}(M)$ be defined as:
$F_{\mathcal{G}_{n,p}}(M)=\mathbb{E}_{G\sim\mathcal{G}_{n,p}}\left[\#\text{ labeled induced copies of }M\text{ in }G\right]\cdot\frac{1}{\binom{n}{\ell}\ell!}$.
$F_{\mathcal{G}_{n,p}}(H)$ captures the density/frequency of $H$ in $\mathcal{G}_{n,p}$. We abbreviate $F_{\mathcal{G}_{n,p}}(H)$ as $F(H)$ in this section, when it is clear from context that the input graph is being compared to the distribution $\mathcal{G}_{n,p}$ over graphs.
Suppose we sample $s$ vertices uniformly at random with repetition. Let
$S_{\mathcal{G}_{n,p}}(H,s)=E_{\mathcal{G}_{n,p}}(H)\cdot\binom{s}{\ell}\frac{1}{n^{\ell}}$ (12)
be the expected number (over $\mathcal{G}_{n,p}$ and the choice of random vertices) of labeled induced copies of $H$ in the subgraph induced by a multiset of size $s$ consisting of vertices chosen uniformly at random (with replacement). It is proven in Claim 5.1.1 that this is the correct expression for the expectation. We abbreviate $S_{\mathcal{G}_{n,p}}(H,s)$ as $S(H,s)$ in this section, when the underlying distribution over input graphs is clear from context.
For a motif $H$ on $k$ vertices, let $\mathcal{H}_{k-1}$ be the set of all (labeled induced) subgraphs of $H$ on k1k-1 vertices. Define $r_{\ell}(H)$ as in Equation 17:
$r_{\ell}(H)=\min_{H_{1}\in\mathcal{H}_{\ell}}\frac{F(H_{1})^{2}}{\max_{\begin{subarray}{c}H_{2}\subseteq H_{1}\\ |H_{2}|=\ell-1\end{subarray}}F(H_{2})^{2}}.$
The quantity $r_{\ell}(H)$ captures the ratio of the density of subgraphs $H_{1}$ on $\ell$ vertices with their densest subgraphs $H_{2}\subseteq H_{1}$ on $\ell-1$ vertices.
Proving the upper bound.
To apply Theorem 5.1 (quality control over motif counts for a general family of distributions over graphs), we need to understand the values of $r_{\ell}$ for $\ell\in\{2,3,\dots,k\}$. We prove that, for $G\sim\mathcal{G}_{n,p}$, the maximum value of $r_{\ell}(H)$ scales with $np^{\Delta(H)}$, where $\Delta(H)$ is the maximum vertex degree in $H$.
Let us first understand $F_{\mathcal{G}_{n,p}}(M)$ for any motif $M$ in the graph. Suppose $M$ is a motif on $\ell$ vertices. Then
$F_{\mathcal{G}_{n,p}}(M)=p^{e(M)}(1-p)^{\binom{\ell}{2}-e(M)}.$
Observe that, for $p\leq 1/2$, $F(M)\leq F(M^{\prime})$ if $e(M)\geq e(M^{\prime})$.
Let us first understand $r_{\ell}(H)$ for $|H|=k,\ell\in[k]$ when $G\sim\mathcal{G}_{n,p}$.
Lemma 4.2.
Suppose $G\sim\mathcal{G}_{n,p}$, $p\leq 1/2$, and consider a motif $H$ on $k$ vertices. For all $\ell\in[k]$, the following holds:
$r_{\ell}(H)=(1-p)^{2\ell-2}\min_{H_{1}\in\mathcal{H}_{\ell}}\left(\frac{p}{1-p}\right)^{2\Delta(H_{1})}.$
Proof.
Let’s expand the expression for $r_{\ell}(H)$ for $G\sim\mathcal{G}_{n,p}$:
$r_{\ell}(H)=\min_{H_{1}\in\mathcal{H}_{\ell}}\frac{F(H_{1})^{2}}{\max_{\begin{subarray}{c}H_{2}\subseteq H_{1}\\ |H_{2}|=\ell-1\end{subarray}}F(H_{2})^{2}}=\min_{H_{1}\in\mathcal{H}_{\ell}}\frac{p^{2e(H_{1})}(1-p)^{2\binom{\ell}{2}-2e(H_{1})}}{\max_{\begin{subarray}{c}H_{2}\subseteq H_{1}\\ |H_{2}|=\ell-1\end{subarray}}p^{2e(H_{2})}(1-p)^{2\binom{\ell-1}{2}-2e(H_{2})}}$
$(1-p)^{2\ell-2}\min_{H_{1}\in\mathcal{H}_{\ell}}\frac{p^{2e(H_{1})}(1-p)^{-2e(H_{1})}}{\max_{\begin{subarray}{c}H_{2}\subseteq H_{1}\\ |H_{2}|=\ell-1\end{subarray}}p^{2e(H_{2})}(1-p)^{-2e(H_{2})}}.$
Since $p\leq 1/2$, in the denominator, the maximum over $H_{2}\subset H_{1}$ of the expression $p^{2e(H_{2})}(1-p)^{-2e(H_{2})}$ is achieved when $H_{2}$ has as few edges as possible. That is, $H_{2}$ is $H_{1}$ with the highest-degree vertex removed; i.e. the maximum is achieved for $H_{2}$ satisfying $e(H_{2})=e(H_{1})-\Delta(H_{1})$.
Therefore, the expression for $r_{\ell}(H)$ becomes:
$r_{\ell}(H)=(1-p)^{2\ell-2}\min_{H_{1}\in\mathcal{H}_{\ell}}\frac{p^{2e(H_{1})}(1-p)^{-2e(H_{1})}}{p^{2e(H_{1})-2\Delta(H_{1})}(1-p)^{-2e(H_{1})+2\Delta(H_{1})}}$
$(1-p)^{2\ell-2}\min_{H_{1}\in\mathcal{H}_{\ell}}\left(\frac{p}{1-p}\right)^{2\Delta(H_{1})},$
as stated in the lemma. ∎
We verify that the values of $r_{\ell}$ decrease as $\ell$ increases. This corresponds to the statement that the density of subgraphs of $H$ on $\ell$ vertices in the subsets of $\ell$ vertices is less than the density of their respective subgraphs on $\ell-1$ vertices in the subsets of $\ell-1$ vertices. This was a necessary condition for applying Theorem 5.1.
Lemma 4.3.
For all $\ell\in\{2,3,\dots,k\}$, $r_{\ell-1}(H)\geq r_{\ell}(H)$.
Proof.
First, observe that $(1-p)^{2(\ell-1)-2}\geq(1-p)^{2\ell-2}$. Next, observe that
$\min_{H_{2}\in\mathcal{H}_{\ell-1}}\left(\frac{p}{1-p}\right)^{2\Delta(H_{2})}\geq\min_{H_{1}\in\mathcal{H}_{\ell}}\left(\frac{p}{1-p}\right)^{2\Delta(H_{1})}.$ (13)
This holds because $\left(p/(1-p)\right)^{c}$ decreases as $c$ increases, for $c>0$. As we consider $\Delta(H_{1})$ over subsets $H_{1}$ on more vertices, the maximum maximum-degree $\Delta(H_{1})$ can only increase, giving us Equation 13.
Combining the two observed inequalities gives us the lemma. ∎
Next, as noted in Section 5, the query complexity will depend on $r_{k}(H)^{-1}$ (the ratio of densities of subgraphs of $H$). We observe the following.
Observation 4.3.
$r_{k}(H)=p^{2\Delta(H)}(1-p)^{2k-2-2\Delta(H)}$.
Proof.
This follows directly from Lemma 4.2, observing that, by definition, $\mathcal{H}_{k}=\{H\}$ when $H$ is a motif on $k$ vertices. ∎
Observation 4.3.
Define $s$ as in Section 5. Then $s=\widetilde{O}\left(r_{k}(H)^{-1}\varepsilon^{-2}\right)=\widetilde{O}\left(p^{-2\Delta(H)}\varepsilon^{-2}\right).$
We are now ready to prove the upper bound in Theorem 4.1 as an implication of Theorem 5.1.
Proof.
Consider a motif $H$ on $k$ vertices. By Section 2.3, for $p=\Omega\left(n^{-1/m_{H}}\right)$ $\mathcal{G}_{n,p}$ satisfies the properties that all $H^{\prime}\subseteq H$ the number of labeled induced copies of $H^{\prime}$ is in the range $(1\pm\frac{\varepsilon}{100})\binom{n}{i}C_{H^{\prime}}$ for some $C_{H^{\prime}}$ with high constant probability, and that $r_{\ell-1}(H)\geq r_{\ell}(H)$ for all $\ell\in\{2,3,\dots,k\}$. Therefore, Theorem 5.1 implies that quality control for the count of $H$ in an input graph $G$ compared to $\mathcal{G}_{n,p}$ can be accomplished in $O\left(r_{k}(H)^{-2}\varepsilon^{-4}\right)=p^{-O(\Delta(H))}\cdot\mathrm{poly}(1/\varepsilon)$ queries. ∎
We conclude with the following observation about quality control for unlabeled or non-induced motifs. First, recall the following facts regarding the relationship between induced/non-induced and labeled/unlabeled motifs in a graph. The number of non-induced copies of a motif $H$ is the sum of the count of all induced motifs $H^{\prime}$ of size $k$ such that the edges of $H$ are a subset of the edges of $H^{\prime}$. For an unlabeled motif $H$, let $\text{Aut}(H)$ be the set of automorphisms of $H$; that is, the set of all permutations of the vertices of $H$ that preserve the structure/adjacencies of $H$. Let $|\text{Aut}(H)|$ be the number of automorphisms of $H$. Recall that the number of labeled copies of a motif $H$ in a graph $G$ is the number of unlabeled copies times $|\text{Aut}(H)|$.
Observation 4.3.
Quality control for non-induced motifs $H$ (of size $k$) compared to $\mathcal{G}_{n,p}$ can be accomplished via quality control algorithms for labeled counts of motifs $H^{\prime}$ of size $k$ such that the edges of $H$ are a subset of the edges of $H^{\prime}$. Quality control for labeled motif counts can be accomplished directly from quality control for unlabeled motif counts, and vice versa.
4.2 Query complexity lower bound
For graphs $H$ and $G$, let $C_{H}(G)$ count the number of induced labeled copies of $H$ in $G$ (or the number of homomorphisms from $H$ to $G$). For a parameter $p\in[0,1]$ and integer $n$, let $\mu_{H}=\mu_{H}^{n,p}$ denote the quantity $\mu_{H}:=\mathbb{E}_{G\sim\mathcal{G}_{n,p}}[C_{H}(G)]$. Finally, let $\rho_{H}(G)=\rho_{H}^{n,p}(G)$ equal $C_{H}(G)/\mu_{H}$. For a graph $H$, let $\Delta(H)$ denote the maximum vertex degree in $H$.
In this section, we prove Theorem 4.1 Part (3), which states that the $(\mathcal{G}_{n,p},\rho_{H})$-quality control problem, i.e., the quality control problem for $\mathcal{G}_{n,p}$ with respect to $\rho_{H}$, requires $p^{-\Omega(\Delta(H))}$ queries for every graph $H$. Specifically, we have the following theorem.
Theorem 4.4.
For every graph $H$, there exists a constant $c>0$ and polynomial $n_{0}$ such that for every $p\in(0,1/2]$ and every $n\geq n_{0}(1/p)$ , every algorithm that solves the $(\mathcal{G}_{n,p},\rho_{H})$-quality control problem has query complexity at least $c\cdot p^{-\Delta(H)/2}$.
Our results follow the standard paradigm for lower bounds in graph property testing. Specifically, we consider two distributions on graphs: a “Yes” distribution [^2] In our case, since we are performing quality control for $\mathcal{G}_{n,p}$, we are forced to set $D_{Y}=\mathcal{G}_{n,p}$. $D_{Y}=D_{Y}^{n}:=\mathcal{G}_{n,p}$ which is mostly supported on graphs $G$ that have $\rho(G)\approx 1$, and a family of “No” distributions $D_{N}=D_{N}^{n,\ell}$ that aims to be mostly supported on graphs with $|\rho(G)-1|\gg\epsilon$. Our (easy) “indistinguishability” lemma (which is independent of $H$) shows that no deterministic $o(n/\ell)$ query algorithm gets much advantage in distinguishing instances drawn from $D_{Y}^{n}$ from instances drawn from $D_{N}^{n,\ell}$. The (slightly more) involved “parameter-separation” lemma shows that for an appropriate choice of $\ell$, depending on $H$, $D_{N}$ does place most of its mass on instances with $|\rho(G)-1|\gg\epsilon$. Putting the two together yields a proof of Theorem 4.4.
Section 4.2 below defines our distributions $D_{N}^{n,\ell}$. It is followed by Lemma 4.5, which proves the indistinguishability of $D_{Y}^{n}$ and $D_{N}^{n,\ell}$, and then by Lemma 4.6. which proves that $D_{N}$ is mostly supported on instances with $\rho$-value far from $1$. We conclude the section with a proof of Theorem 4.4.
Definition 4.4.
We describe $D_{N}$ by showing how to sample a graph $G^{\prime}$ according to this distribution: $G^{\prime}\sim D_{N}^{n,\ell}$ is sampled as follows: We first sample $G\sim\mathcal{G}_{n,p}$. Then we sample $T\sim\mathcal{G}_{n,\frac{1}{2}}$. Finally we sample $S\subseteq[n]$ with $|S|=\ell$ uniformly among all such sets. Our output graph $G^{\prime}=G^{\prime}(G,S,T)$ is set to be equal to $G$ on $\overline{S}\times\overline{S}$ and equal to $T$ otherwise. In other words, $(u,v)\in E(G^{\prime})$ if $\{u,v\}\cap S=\emptyset$ and $(u,v)\in E(G)$ or if $\{u,v\}\cap S\neq\emptyset$ and $(u,v)\in E(T)$.
Lemma 4.5.
Let $A$ be a deterministic algorithm making $q$ queries to the adjacency matrix of an $n$-vertex graph and outputting $0/1$. Then
$\left|\Pr_{G\in D_{Y}^{n}}[A(G)=1]-\Pr_{G^{\prime}\in D_{N}^{n,\ell}}[A(G^{\prime})=1]\right|\leq\frac{2q\ell}{n-\ell}.$
Proof.
We prove something even stronger, namely that for every graph $G$, with high probability over the choice of $S\subseteq[n]$ with $|S|=\ell$ and for every graph $T$ we have
$\Pr_{S}\left[A(G^{\prime}(G,S,T))\neq A(G)\right]\leq\frac{2q\ell}{n-\ell},$
where the notation $G^{\prime}(G,S,T)$ is as in Section 4.2. The lemma follows by sampling $G\sim\mathcal{G}_{n,p}$ and $T\sim\mathcal{G}_{n,\frac{1}{2}}$.
Note every query of $A$ to its input graph is of the form “Is $(u,v)\in E$?”. Let $(u_{i},v_{i})$, for $i=1$ to $q$ denote the $q$ queries of $A$ on input $G$. (Since $A$ is deterministic and the input graph is fixed, these queries are completely determined.) Let $Q=\{u_{i}~|~i\in[q]\}\cup\{v_{i}~|~i\in[q]\}$. We observe that if $Q\cap S=\emptyset$ then $A(G)=A(G^{\prime}(G,S,T))$ since every query gets the same answer in $G^{\prime}$ as in $G$ and $A$’s queries and output are deterministic functions of previous queries and their answers. The lemma follows from the simple claim that for every set $Q\subseteq[n]$ with $|Q|\leq 2q$ the probability that it intersects a random set $S\subseteq[n]$ of size $\ell$ is at most $2q\ell/(n-\ell)$. ∎
Lemma 4.6.
For every graph $H$ of maximum degree $\Delta$, there exists a polynomially growing function $f$ such that for every $p\in(0,1/2]$ and $n\geq 20(2k)^{2k+1}p^{-\Delta}/\varepsilon^{2}$, and every $\ell\geq f(1/p)$ and $n/2\geq\ell\geq 2^{2\binom{k}{2}+1}np^{\Delta/2}$, the following hold:
$\Pr_{G\sim D_{Y}^{n}}[|\rho_{H}(G)-1|\leq 0.1]\geq 1-1/12,$
 $\mbox{ and }\Pr_{G^{\prime}\sim D_{N}^{n,\ell}}[\rho_{H}(G^{\prime})\geq 1.9]\geq 1-1/12.$
Proof.
We prove the lemma for $f(x)=x^{\Delta}$.
On the one hand we have that the expected number of induced labeled copies of $H$ in $G\sim\mathcal{G}_{n,p}$ is $\mu_{H}=\binom{n}{k}k!\cdot p^{e(H)}(1-p)^{\binom{k}{2}-e(H)}=n^{k}p^{e(H)}(1-p)^{\binom{k}{2}-e(H)}(1\pm o_{n}(1))$, and this is concentrated provided $n\geq p^{-m_{H}}$ by Section 2.3, where $m_{H}=\max_{H^{\prime}\subseteq H}\frac{e(H^{\prime})}{v(H^{\prime})}$. This holds for $n\geq 20(2k)^{2k+1}p^{-\Delta(H)}/\varepsilon^{2}$.
As a consequence, we have
$\Pr_{G\sim D_{Y}^{n}}[|C_{H}(G)-\mu_{H}|\leq 0.1\cdot\mu_{H}]=1-o_{n}(1)\geq 1-1/12.$
We now turn to analyzing $\rho_{H}(G^{\prime})$ for $G^{\prime}\sim D_{N}$. Let the vertices of $H$ be denoted as $h_{1},\ldots,h_{k}$, with the degree of $h_{k}$ being $\Delta$. Let $e(H)$ denote the number of edges of $H$. Let $H^{\prime}$ be the induced subgraph on vertices $h_{1},\ldots,h_{k-1}$. Note that the number of edges of $H^{\prime}$ is exactly $e(H^{\prime})=e(H)-\Delta$. We lower bound the expected number of induced copies of $H$ in $G^{\prime}$ by first bounding $N_{1}$, the number of induced labeled copies of $H^{\prime}$ in $G^{\prime}$, and then lower bounding the expected number of extensions of a given copy of $H^{\prime}$. We claim that $N_{1}\geq\mu_{H^{\prime}}\cdot 2^{-\binom{k-1}{2}}$ with high probability since every edge is present in $G^{\prime}$ with probability at least $p$ and absent with probability at least $1/2$. Fix such a copy $A$ of $H^{\prime}$ in $G^{\prime}$, say on vertices $v_{1},\ldots,v_{k-1}$. Now consider picking a random vertex $v_{k}\in[n]\setminus\{v_{1},\ldots,v_{k-1}\}$. We claim the probability that this extends $A$ into a copy of $H$ is at least $\frac{\ell-k}{n-k}2^{-(k-1)}$, where the probability is over the random choice of $S$ and $T$. In particular $v_{k}$ lands in $S$ with probability at least $(\ell-k)/(n-k)$ and conditioned on this event, each edge from $v_{k}$ to $v_{i}$ is present or absent in $T$ with probability exactly $1/2$. Adding up over the nkn-k choices of $v_{k}$ we conclude that each copy of $H^{\prime}$ in $G^{\prime}$ extends to $(\ell-k)2^{-(k-1)}$ copies of $H$ in expectation. Thus $\mathbb{E}[C_{H}(G^{\prime})]\geq N_{1}\cdot(\ell-k)2^{-(k-1)}\geq\mu_{H^{\prime}}(\ell-k)2^{-\binom{k}{2}}$. To compare this with $\mu_{H}$, note that $\mu_{H}=\mu_{H^{\prime}}\cdot n\cdot p^{\Delta}(1-p)^{k-1-\Delta}(1\pm o_{n}(1))$. Thus
$\mathbb{E}[C_{H}(G^{\prime})]$ $\geq\mu_{H^{\prime}}(\ell-k)2^{-\binom{k}{2}}$
$\geq\mu_{H}n^{-1}p^{-\Delta}(1-p)^{-(k-1-\Delta)}(1\pm o_{n}(1))(\ell-k)2^{-\binom{k}{2}}$
$\geq\mu_{H}\frac{\ell}{2^{\binom{k}{2}+2}np^{\Delta}}$
where the final inequality uses $(1-p)^{-(k-1-\Delta)}\geq 1$, $\ell\geq 2k$ and that $n$ is large enough so that the $o_{n}(1)$ term is $\leq 1/2$. Thus if $\ell\geq 2^{\binom{k}{2}+3}p^{\Delta}n$ we get that $\mathbb{E}_{G^{\prime}}[C_{H}(G^{\prime})]\geq 2\mu_{H}$. By applying Section D.2, we get that
$\Pr_{G^{\prime}\sim D_{N}^{n,\ell}}[C_{H}(G^{\prime})\geq 1.9\mu_{H}]\geq 1-1/12,$
provided $n\geq 20(2k)^{2k+1}p^{-\Delta(H)}/\varepsilon^{2}$, $\ell\geq(1/p)^{\Delta}$, and $n/2\geq\ell\geq 2^{2\binom{k}{2}+1}np^{\Delta/2}$. ∎
Proof of Theorem 4.4..
We prove the theorem for $n_{0}(x)=f(x)\cdot x^{\Delta}$ and $c=2^{2\binom{k}{2}+1}/12$. Let $n\geq n_{0}(1/p)$. Let $\ell=2^{2\binom{k}{2}+1}np^{\Delta}$. In particular our choice of $n_{0}(x)$ will be set to ensure that the $o_{\ell}(1)$ and $o_{n}(1)$ terms in Lemma 4.6 are at most $1/12$.
Let $A$ be an algorithm making $q$ queries that accepts $G\sim\mathcal{G}_{n,p}$ with probability at least $2/3$ while accepting $G$ with $\rho_{H}(G)>1.5$ with probability at most $1/3$. Now note that by our choice of $\ell$ we have, by Lemma 4.6, that $\Pr_{G^{\prime}\sim D_{N}^{n,\ell}}[\rho_{H}(G^{\prime})\leq 1.9]=o_{\ell}(1)\leq 1/12$ (by our choice of $n$). So we get $\Pr_{G\sim\mathcal{G}_{n,p},R}[A(G)=1]-\Pr_{G^{\prime}\sim D_{N}^{n,\ell},R}[A(G^{\prime})=1]\geq 1/6$, where the randomness $R$ denotes the random coin tosses of the algorithm $A$. In particular there exists $R_{0}$ such that $\mathbb{E}_{G\sim\mathcal{G}_{n,p}}[A^{\prime}(G)=1]-\mathbb{E}_{G^{\prime}\sim D_{N}^{n,\ell}}[A^{\prime}(G^{\prime})=1]\geq 1/6$ where $A^{\prime}$ denotes the deterministic version of $A$ obtained by fixing its randomness to $R_{0}$. Applying Lemma 4.5, we get that $2q\ell/(n-\ell)\geq 1/6$, which in turn implies $q\geq(n-\ell)/(12\ell)$. ∎
4.3 Algorithmic complexity
We now move to proving Theorem 4.1 Part (2) regarding the algorithmic complexity of quality control of motifs compared to $\mathcal{G}_{n,p}$. We re-state the result below. Recall the definition of $\sigma(H)$ from Equation 2. Let $m_{H}^{\star}$ be the maximum value of $m_{H^{\prime}}$ over all $H^{\prime}\supseteq H$ with $|H^{\prime}|=|H|$ and edges that are a superset of the edges of $H$.
Theorem 4.7 (Theorem 4.1 Part (2)).
For constant error parameter $\varepsilon$ and
$p=\Omega\left(\max\left\{\left(\log(n)/n\right)^{1/(4\sigma(H)-1)},n^{-1/m_{H}^{\star}}\right\}\right)$, and induced motif $H$ on $k$ vertices, there is an algorithm for $(\mathcal{G}_{n,p},\rho_{H})$-quality control that makes $\frac{1}{p^{O(\Delta(H))}}$ adjacency matrix queries and has a runtime of $\frac{1}{p^{O(\Delta(H))}}$.
While the above is stated for constant $\varepsilon$, the bounds with the explicit dependence on $\varepsilon$ are: $p^{-O(k)}\cdot\mathrm{poly}(1/\varepsilon)$ for the query complexity, and $p^{-O(k)}\cdot\exp(\mathrm{poly}(|H|/\varepsilon))$ for the runtime.
4.3.1 Non-induced motifs
We begin by studying non-induced motifs, moving to induced motifs in Section 4.3.2. Our query complexity upper bounds so far hold for the setting of non-induced motifs, as noted in Observation 4.1. We will extend Section 3.3, using similar ideas of testing graph jumbledness and then applying an algorithm from Section 2.2. We utilize the fact that the algorithm from Section 2.2 already holds for approximately counting motifs in jumbled graphs.
We begin by proving Theorem 4.7 for the quality control of the counts of non-induced motifs, and then use an inclusion-exclusion principle argument to reason about induced motifs in . Considering non-induced motifs allows us to more directly apply the efficient algorithm from Section 2.2 for approximate counting of motifs in sufficiently jumbled graphs.
For motif $H$, let $j=\lceil\sigma(H)\rceil$ and $|H|=k$.
In this section we are considering the case where $p\geq\left(\frac{100C^{4}\log n}{n}\right)^{1/(4j-1)}$, where $C$ is a constant $C=C(\varepsilon,k)=\exp\left(\text{poly}(k/\varepsilon)\right)$. In this setting of $p$, we can use the jumbledness of small induced subgraphs of $\mathcal{G}_{n,p}$.
For a motif $H$, recall the definition of $\sigma(H)$ from Equation 2. Suppose that $j=\lceil\sigma(H)\rceil$. Let $S(H,s)$ be the expected number of copies of $H$ in a multiset on $s$ vertices (with repetition), as in Equation 12.
Consider the following parameters.
Definition 4.7.
For $2\leq\ell\leq k-1$, define
$\zeta_{\ell}=\varepsilon\cdot\left(\frac{4}{7}\right)^{k-\ell-1}.$
Define $\eta_{2}=\varepsilon^{2}r_{2}(H)\cdot\left(\frac{4}{7}\right)^{2(k-3)}/128$, and for $\ell\in\{3,4,\dots,k\}$ define
$\eta_{\ell}=\frac{r_{\ell}(H)\cdot\zeta_{\ell-1}^{2}}{1024\cdot\ell^{2}}.$
We also redefine $s$ as follows.
Definition 4.7.
Let $C$ be a constant $C=C(\varepsilon,k)=\exp(\mathrm{poly}(k/\varepsilon))$. Define $s_{*}$ to be:
$s_{*}=\frac{1200C^{2}\cdot k^{2}\cdot 1024}{p^{5\sigma(H)}\cdot\varepsilon^{2}}\cdot\left(\frac{7}{4}\right)^{2(k-\ell)}\cdot 8\cdot\ln\left(\frac{(6k)^{k}}{p^{\binom{k+2}{2}}\cdot\varepsilon}\right).$
We apply the following algorithm, which is an extension of the algorithm in Section 3.3.
Algorithm NonInduced-Motif-Quality-Efficient:
On inputs $G,n,p,\mathcal{G}_{n,p},\varepsilon$, and non-induced motif $H$
  1. 1.
    Sample $s_{*}$ vertices, each uniformly at random (with repetition). Let $S$ be the corresponding set of vertices, with induced graph $G_{S}$.
  2. 2.
    Perform adjacency matrix queries on all pairs in $S$ to find all edges.
  3. 3.
    Test that $G_{S}$ is $(p,\beta)$-jumbled for $\beta\leq\frac{p^{j}s_{*}}{C}$ for a constant $C=\exp\left(\text{poly}(k/\varepsilon)\right)$ by checking the following: Let $\delta=O(p^{2j}/C^{2})$. For every $u\in S$, check that the degree of $u$ in $S$ is $(s_{*}-1)(p\pm\delta)$. For every $(u,v)\in S\times S$, check that the number of common neighbors of $(u,v)$ is $(s_{*}-2)(p^{2}\pm\delta)$. If any pair in $S\times S$ does not satisfy this, reject.
  4. 4.
    Else, for each $\ell\leq k$ and each $H^{\prime}\subseteq H$ with $|H^{\prime}|=\ell$, approximate the number of labeled non-induced copies of $H^{\prime}$ in $S$ to within a $\zeta_{\ell-1}S(H^{\prime},s_{*})/8$ additive factor, in time $O(s_{*}^{c})$ for a constant $c$, using the algorithm referenced in Section 2.2 (applied with error parameter $\varepsilon/k^{k}$). If, for any $H^{\prime}$, the approximation of the number of $H^{\prime}$ copies in $G_{S}$ is not in $(1\pm\frac{5}{8}\varepsilon)S(H^{\prime},s_{*})$, reject. Otherwise, accept.
Proof of Theorem 4.7.
This proof follows the same steps as the proof of Theorem 3.6 from Section 3.3 (using our new setting of parameters $\gamma_{\ell},\eta_{\ell}$) for this section). We therefore omit the proof.
The two changes are the following. For completeness, now apply Section 2.3 for $\beta\leq p^{\sigma(H)}s_{*}/C$ (that is, replace $k$ by $\sigma(H)$), and let $C=\exp(\text{poly}(k/\varepsilon))$ as before (where $|H|=k$). We can apply Section 2.3 because $s_{*}$ is set so that $s_{*}\geq 1200C^{4}/p^{5\sigma(H)}$.
Additionally, completeness, which relies on concentration of $H$ counts in $\mathcal{G}_{n,p}$, requires $p=\Omega(n^{-1/m_{H}^{\star}})$, as defined before Theorem 4.7. Additionally, instead of relying on the inductive and concentration lemmas from Section 3.1 for $k$-cliques (Lemma 3.2 and Lemma 3.3, respectively), rely on the inductive and concentration lemmas from Section 5.1 (Lemma 5.2 and Lemma 5.3, respectively), which hold for general motifs and a range of distributions over graphs.
We now argue about the query complexity and runtime. Since the algorithm queries all pairs of vertices in a set $S$ of size $s_{*}$ (Section 4.3.1), the algorithm uses $\widetilde{O}\left(p^{-10\lceil\sigma(H)\rceil}\varepsilon^{-4}\right)$ adjacency matrix queries. As noted in [CFZ14], $\sigma(H)\leq\Delta(H)+2$, and so the algorithm uses $1/(p^{O(\Delta(H))}\varepsilon^{4})$ adjacency matrix queries. The runtime is $Cs_{*}^{c}$ for some constants $C=$
$\exp(\text{poly}(|H|/\varepsilon))$ and $c\geq 3$, which is $1/(p^{O(\Delta(H))})\cdot\exp(\text{poly}(|H|/\varepsilon))$. ∎
4.3.2 Moving to induced motifs
We can transfer the results/analysis for non-induced motifs to induced motifs as follows. Observe that we can write the number of induced motifs $H$ in the graph in terms of a sum over (possibly negated) counts of all of the non-induced counts of motifs $H^{\prime}\subseteq H$. Therefore, using the inclusion-exclusion principle allows us to move to induced motifs.
We apply the following algorithm to transfer the results for non-induced motifs to induced motifs. Suppose $H$ is supported on $k$ vertices. Below, let $H\subseteq H^{\prime},|H^{\prime}|=k$ mean that $H^{\prime}$ is a motif on $k$ vertices whose edges are a superset of the edges of $H$.
Algorithm Induced-Motif-Quality-Efficient:
On inputs $G,n,p,\mathcal{G}_{n,p},\varepsilon$, and induced $H$
  1. 1.
    For every $H^{\prime}$ satisfying $H\subseteq H^{\prime}$ and $|H^{\prime}|=k$:
    1. (a)
      Run Algorithm NonInduced-Motif-Quality-Efficient with inputs $G,n,p,\mathcal{G}_{n,p},\frac{\varepsilon}{2^{k(k-1)}}$, and non-induced $H^{\prime}$. If Algorithm NonInduced-Motif-Quality-Efficient rejects, then reject. Else, proceed.
  2. 2.
    Accept.
The proof of correctness of this algorithm will rely on the following inclusion-exclusion formula relating induced counts to non-induced motif counts. For a motif $P$, let $C_{P}(G)$ be the number of labeled induced copies of $P$ in $G$, and let $\widetilde{C}_{P}(G)$ be the number of labeled non-induced copies of $P$ in $G$.
Observation 4.7.
Consider any motif $H$ on $k$ vertices, and any graph $G$. Then
$C_{H}(G)=\sum_{i=0}^{\binom{k}{2}-e(H)}(-1)^{i}\cdot\sum_{\begin{subarray}{c}H^{\prime}\supseteq H\\ |H^{\prime}|=k\\ e(H^{\prime})=e(H)+i\end{subarray}}\widetilde{C}_{H^{\prime}}(G).$
We also observe that Section 2.3 can imply that, for $p=\Omega\left(\max_{H^{\prime}\supseteq H,|H^{\prime}|=k}n^{-1/m_{H^{\prime}}}\cdot\varepsilon^{-2}\right)$, with probability at least $9/10$, for all $H^{\prime}\supseteq H,|H^{\prime}|=k$, the number of copies of $H^{\prime}$ in $G$ is in the range $(1\pm\frac{\varepsilon}{100k^{k}}\binom{n}{k}p^{e(H^{\prime})}(1-p)^{\binom{k}{2}-e(H)}$. This is because, as noted in Section 2 after Section 2.3, the choice of $\varepsilon/100$ was not specific and can be changed with the hidden constants in the lower bound on $p$. The same applies for the error probability.
We now prove Theorem 4.7 for induced motifs.
Proof of Theorem 4.7.
Completeness: Suppose $G\sim\mathcal{G}_{n,p}$. Then, by the observation made above, with probability at least $9/10$, for all $H^{\prime}\supseteq H,|H^{\prime}|=k$, the number of copies of $H^{\prime}$ in $G$ is in the range $(1\pm\frac{\varepsilon}{100k^{k}}\binom{n}{k}p^{e(H^{\prime})}(1-p)^{\binom{k}{2}-e(H)}$. Additionally, with similar (up to constants) parameter choices for the algorithm, the correctness of Algorithm NonInduced-Motif-Quality-Efficient on an input $H$ can be boosted to $1-1/(10\cdot 2^{\binom{k}{2}})$. Therefore, by the correctness of NonInduced-Motif-Quality-Efficient, with probability at least $9/10$, all calls to Algorithm NonInduced-Motif-Quality-Efficient in Step 1 of Algorithm Induced-Motif-Quality-Efficient accept. Therefore, Algorithm Induced-Motif-Quality-Efficient accepts $G\sim\mathcal{G}_{n,p}$ with high probability.
Soundness: Let $\mu_{H}=\mathbb{E}_{G^{\prime}\sim\mathcal{G}_{n,p}}\left[C_{H}(G^{\prime})\right]=\binom{n}{k}p^{e(H)}(1-p)^{\binom{k}{2}-e(H)}$ (induced count) and $\widetilde{\mu}_{H}=\mathbb{E}_{G^{\prime}\sim\mathcal{G}_{n,p}}\left[\widetilde{C}_{H}(G^{\prime})\right]=\binom{n}{k}p^{e(H)}$ (noninduced count). Suppose that $G$ satisfies $C_{H}(G)\not\in(1\pm\varepsilon)\mu_{H}$. By the triangle inequality, we find that:
$\varepsilon\mu_{H}\leq\left|C_{H}(G)-\mu_{H}\right|=\left|\sum_{i=0}^{\binom{k}{2}-e(H)}(-1)^{i}\cdot\sum_{\begin{subarray}{c}H^{\prime}\supseteq H\\ |H^{\prime}|=k\\ e(H^{\prime})=e(H)+i\end{subarray}}\widetilde{C}_{H^{\prime}}(G)-\widetilde{\mu}_{H^{\prime}}\right|$
$\leq\sum_{i=0}^{\binom{k}{2}-e(H)}\sum_{\begin{subarray}{c}H^{\prime}\supseteq H\\ |H^{\prime}|=k\\ e(H^{\prime})=e(H)+i\end{subarray}}\left|\widetilde{C}_{H^{\prime}}(G)-\widetilde{\mu}_{H^{\prime}}\right|.$
Therefore, there exists an $H^{\prime}\supset H,|H^{\prime}|=k$ such that $\left|\widetilde{C}_{H^{\prime}}(G)-\widetilde{\mu}_{H^{\prime}}\right|\geq\varepsilon\mu_{H}/2^{\binom{k}{2}}$. We refer to this $H^{\prime}$ as $H^{\circ}$
Next observe that, for $H^{\prime}\supseteq H,|H^{\prime}|=k$, $\mu_{H^{\prime}}\leq\mu_{H}$. Applying this and $\widetilde{\mu}_{H^{\prime}}\leq\mu_{H^{\prime}}\cdot 2^{\binom{k}{2}}$, we find that:
$\left|\widetilde{C}_{H^{\circ}}(G)-\widetilde{\mu}_{H^{\circ}}\right|\geq\varepsilon\mu_{H^{\circ}}/2^{\binom{k}{2}}\geq\varepsilon\widetilde{\mu}_{H^{\circ}}/2^{k(k-1)}.$
Therefore, by the soundness of Algorithm NonInduced-Motif-Quality-Efficient, when Algorithm NonInduced-Motif-Quality-Efficient is called on input $H^{\circ}$ in Step 1 of Algorithm Induced-Motif-Quality-Efficient, with high constant probability the algorithm will reject.
Therefore, Algorithm Induced-Motif-Quality-Efficient rejects $G$ with $C_{H}(G)\not\in(1\pm\varepsilon)\mu_{H}$ with high probability.
Query complexity and runtime: Since $k$ is constant, and so the number of $H^{\prime}$ considered in Step 1 is constant, the query complexity and runtime are asymptotically the same as those of Algorithm NonInduced-Motif-Quality-Efficient. ∎
5 Quality control for other random graph models
In this section, we generalize our results to the setting of quality control of any motif $H$ as compared to any distribution $\mathcal{D}_{n}$ over graphs with a good concentration of motif counts. Particularly, we will consider distributions $\mathcal{D}_{n}$ over graphs such that, for $G\sim\mathcal{D}_{n}$, for all $P\subseteq H$ the number of labeled induced copies of $P$ is in the range $(1\pm\frac{\varepsilon}{100})E_{\mathcal{D}_{n}}(P)$ for some $E_{\mathcal{D}_{n}}(P)$ with high constant probability.
Before stating the main theorem we prove in this section (Theorem 5.1), we set up some notation and terminology.
Notation and terminology
In what follows, for distribution $\mathcal{D}_{n}$ over graphs and motif $P$ on $\ell$ vertices, define $E_{\mathcal{D}_{n}}(P),F_{\mathcal{D}_{n}}(P)$ as follows:
$E_{\mathcal{D}_{n}}(P)=\mathbb{E}_{G\sim\mathcal{D}_{n}}\left[\#\text{ labeled induced copies of }P\text{ in }G\right]=\binom{n}{\ell}\ell!\cdot F_{\mathcal{D}_{n}}(P).$ (14)
We will refer to $F_{\mathcal{D}_{n}}(P)$ as the density of $P$ in $G\sim\mathcal{D}_{n}$, since it reflects the probability that a randomly chosen set of size $\ell$ is equivalent to a labeled induced copy of $P$.
Sample $s\in\mathbb{N}$ vertices uniformly at random with repetition. Let
$A_{s,\ell}=\binom{s}{\ell}\cdot\ell!\cdot\frac{1}{n^{\ell}}.$ (15)
Then,
$S_{\mathcal{D}_{n}}(P,s)=E_{\mathcal{D}_{n}}(P)\cdot A_{s,\ell}$ (16)
is the expected number of labeled induced copies of $P$ in the subgraph induced by the multiset of vertices (as proven below in Claim 5.1.1).
We begin by defining the notion of the density increments of a motif, which will appear in the query complexity upper bound achieved for $(\mathcal{D}_{n},\rho_{H})$-quality control.
We begin by defining the density increments $r_{\ell}(H)$ of a motif $H$.
Definition 5.0.
Consider a motif $H$ on $k$ vertices. For $1\leq\ell\leq k$, let $\mathcal{H}_{\ell}$ consist of all (labeled, induced) subgraphs of $H$ on $\ell$ vertices. Let
$r_{\ell}(H)=\min_{H_{1}\in\mathcal{H}_{\ell}}\frac{F_{\mathcal{D}_{n}}(H_{1})^{2}}{\max_{\begin{subarray}{c}H_{2}\subseteq H_{1}\\ |H_{2}|=\ell-1\end{subarray}}F_{\mathcal{D}_{n}}(H_{2})^{2}}.$ (17)
This is the minimum ratio between the density of any subgraph $H_{1}$ of $H$ on $\ell$ vertices and the densest subgraph $H_{2}\subset H_{1}$ of $H_{1}$ on $\ell-1$ vertices.
To understand the definition of the density increment, let us return briefly to the case of $\mathcal{G}_{n,p}$ (with $p\leq 1/2$). For any $P$ ($|P|=j$), the maximum $\max_{\begin{subarray}{c}P^{\prime}\subseteq P\\ |P|=j-1\end{subarray}}F_{\mathcal{D}_{n}}(P^{\prime})^{2}$ will be achieved by removing the highest-degree vertex in $P$. This means, for example, that $r_{k}(H)$ (for $H$ with $|H|=k$) satisfies $r_{k}(H)\geq p^{\Theta\left(\Delta(H)\right)}$, where $\Delta(H)$ is the maximum vertex degree in $H$.
Let $C_{H}(G)$ be the number of labeled induced copies of $H$ in the graph $G$. Let $\rho_{H}$ be defined as $\rho_{H}(G):=C_{H}(G)/\mathbb{E}_{G^{\prime}\sim\mathcal{D}_{n}}\left[C_{H}(G^{\prime})\right]$.
Theorem 5.1.
Consider a motif $H$ on $k$ vertices. Consider any distribution $\mathcal{D}_{n}$ over graphs satisfying the following properties:
  1. 1.
    For $G\sim\mathcal{D}_{n}$, for all $\ell\in[k]$ and all $P\subseteq H$ on $\ell$ vertices, the number of labeled induced copies of $P$ is in the range $(1\pm\frac{\varepsilon}{100})E_{\mathcal{D}_{n}}(P)$ with high constant probability.
  2. 2.
    For all $\ell\in\{2,3,\dots,k\}$, $r_{\ell}(H)\leq(4/9)r_{\ell-1}(H)$.
Then, there exists an algorithm for $(\mathcal{D}_{n},\rho_{H})$-quality control problem that uses $\widetilde{O}\left(r_{k}(H)^{-2}\right)$ adjacency matrix queries to the input graph.
The algorithm has a runtime of $\widetilde{O}\left(r_{k}(H)^{-2k}\right)$. Additionally, the bounds with the dependence on $\varepsilon$ are $\widetilde{O}\left(r_{k}(H)^{-1}\varepsilon^{-4}\right)$ for the query complexity and $\widetilde{O}\left(r_{k}(H)^{-2k}\varepsilon^{-4k}\right)$ for the runtime.
Other random graph models:
We now apply Theorem 5.1 to other random graph models. For illustrative purposes, we focus on the case of $k$-cliques. The condition we need out of the random graph models we can study is the concentration of the counts of all $k^{\prime}$-cliques within the model, for $k^{\prime}\leq k$. The bound on the query complexity also requires an understanding of a lower-bound on $\mathbb{E}\left[C_{k}(G)\right]/\mathbb{E}\left[C_{k-1}(G)\right]$, where the expectations are taken with respect to the other models/distributions considered.
Theorem 5.1 implies the following results, when we specify the distribution $\mathcal{D}_{n}$ that we apply the result to.
Corollary 5.1 (Implication of Theorem 5.1).
Consider $(\mathcal{D}_{n},\rho_{k})$-quality control, for $\rho_{k}(G)=C_{k}(G)/\mathbb{E}_{G^{\prime}\sim\mathcal{D}_{n}}\left[C_{k}(G^{\prime})\right]$. Define parameters $n,p$ where $p=\Omega\left(1/n^{2/(k-1)}\right)$ and $p\leq 1/2$. For the following settings of $\mathcal{D}_{n}$, the query complexity of $(\mathcal{D}_{n},\rho_{k})$-quality control is $1/p^{O(k)}$:
  1. 1.
    $\mathcal{D}_{n}=\mathcal{G}_{n,p}$.
  2. 2.
    $\mathcal{D}_{n}$ is a Stochastic Block Model (as defined, e.g. in [HLL83]) with a constant number of communities and minimum in-community/between-community probability at least $p$.
  3. 3.
    $\mathcal{D}_{n}$ is the uniform distribution over $d$-regular graphs on $n$ vertices, where $d=np$.
Proof.
Part (1) is already proven in Section 3. It follows because concentration of $k^{\prime}$-clique counts for $k^{\prime}\leq k$ holds for $p=\Omega\left(1/n^{2/(k-1)}\right)$. Additionally, $r_{k}(H)=p^{2(k-1)}$, so the query complexity is $1/p^{O(k)}$.
For Part (2), concentration of $k^{\prime}$-clique counts for $k^{\prime}\leq k$ can be proven similarly to the proof for a specific stochastic block model in Section D.2, when $p=\Omega\left(1/n^{2/(k-1)}\right)$. We can show that $\mathbb{E}_{G\sim\mathcal{D}_{n}}\left[C_{k}(G)\right]\geq p^{k-1}(n-k+1)\mathbb{E}_{G\sim\mathcal{D}_{n}}\left[C_{k-1}(G)\right]$:
$\mathbb{E}_{G\sim\mathcal{D}_{n}}\left[C_{k}(G)\right]=\mathbb{E}\left[\sum_{S:|S|=k-1}\sum_{i\in[n],i\not\in S}\mathbbm{1}(S\equiv(k-1)\text{-clique})\cdot\mathbbm{1}((i,j)\in e(G)~\forall j\in S)\right].$
Since the multiplied indicator random variables in each term of the sum are independent, this is at least:
$\geq\mathbb{E}_{G\sim\mathcal{D}_{n}}\left[C_{k-1}(G)\right]\cdot p^{k-1}(n-k+1).$
Therefore, $r_{k}(H)\geq p^{O(k)}$, and so the query complexity is $1/p^{O(k)}$.
In Part (3), the concentration and computation of the expected value come from [KSV07]. Particularly, the expected value is around $\binom{n}{k}k!\cdot p^{\binom{k}{2}}$, and so $r_{k}(H)$ is $p^{O(k)}$. Therefore, the query complexity is $1/p^{O(k)}$. ∎
We remark that there are other well-studied random graph models for which our results may apply, such as random geometric graphs and inhomogeneous random graphs (i.e., Chung-Lu random graphs). See [BR18, DGLU11] for references about the concentration and expectation of $k$-clique counts in random geometric graphs, and [FGKS24] for inhomogeneous random graphs.
The algorithm:
We consider the following algorithm, which extends the algorithm for $k$-clique counts compared to $\mathcal{G}_{n,p}$. Let the sample size be $s_{*}=\widetilde{O}(r_{k}(H)^{-1}\varepsilon^{-2})$.
Algorithm Motif-Quality:
On inputs $G$, $n,\varepsilon$, $\mathcal{D}_{n}$, and $H$
Sample $s_{*}$ vertices, each uniformly at random (with repetition). Let $S$ be the corresponding multiset of vertices. and each $H^{\prime}\subseteq H$ on $\ell$ vertices, count the number of labeled induced copies of $H^{\prime}$ from $S$ in $G$ (i.e., $C_{H^{\prime}}(S)$ from Section 5). If, for any $\ell$, there exists an $H^{\prime}$, $|H^{\prime}|=\ell$, such that $C_{H^{\prime}}(S)$ is not in $(1\pm\varepsilon/2)S_{\mathcal{D}_{n}}(H^{\prime},s_{*})$, reject. Otherwise, accept.
Parameters.
We will analyze concentration and inductive lemmas over general parameters, then later set the parameters to the following (which are extensions of $\alpha_{\ell},\varepsilon_{\ell},s_{\ell}$ from previous sections).
Definition 5.1.
For $2\leq\ell\leq k-1$, define
$\varepsilon_{\ell}=\varepsilon\cdot\left(\frac{2}{3}\right)^{k-\ell-1}.$
Define $\alpha_{2}=\varepsilon^{2}r_{2}(H)\cdot\left(\frac{2}{3}\right)^{2(k-3)}/128$, and for $2\leq\ell\leq k$ define
$\alpha_{\ell}=\frac{r_{\ell}(H)\cdot\varepsilon_{\ell-1}^{2}}{4096\cdot\ell^{4}}.$
We will specifically consider the following setting of $s_{\ell}$. We use the same notation as the sample multiset size from $k$-cliques because the $s_{\ell}$ in this setting is a generalization of the $s_{\ell}$ from the $k$-cliques case.
Definition 5.1.
Suppose that $H$ is a motif on $k$ vertices. For $\ell\in[k]$, define
$s_{\ell}:=\frac{\ell^{4}\cdot 4096}{r_{\ell}(H)\cdot\varepsilon^{2}}\cdot\left(\frac{3}{2}\right)^{2(k-\ell)}\cdot 8\cdot\ln\left(\frac{(6\ell)^{\ell}}{F_{\mathcal{D}_{n}}(H)\cdot\varepsilon}\right).$
Definition 5.1 (Sample size ($s_{*}$)).
Suppose that $H$ is a motif on $k$ vertices. Then define $s_{*}:=s_{k}$, i.e.:
$s_{*}:=\frac{k^{4}\cdot 4096}{r_{k}(H)\cdot\varepsilon^{2}}\cdot 8\cdot\ln\left(\frac{(6k)^{k}}{F_{\mathcal{D}_{n}}(H)\cdot\varepsilon}\right).$
Observation 5.1.
Observe that $s_{*}$ from Section 5 satisfies:
$s_{*}\geq\frac{1}{\alpha_{k}}\cdot 8\cdot\ln\left(\frac{(6k)^{k}}{F_{\mathcal{D}_{n}}(H)\cdot\varepsilon}\right).$
Therefore, $4^{k+1}\exp(-\alpha_{k}s_{*})\leq\gamma\cdot F_{\mathcal{D}_{n}}(P)/(16\cdot 2^{k})$.
Definition 5.1.
For a multiset $S$ of size $s$, and a motif $P$, let $C_{P}(S)$ be the number of induced labeled copies of $P$ in $S$. Since $S$ is a multiset, copies of $P$ are counted as follows: if a vertex appears $c$ times in the multiset, the induced labeled copies of $P$ it is a part of are counted $c$ times.
5.1 Technical Lemmas
In this section, we prove general versions of the inductive lemma (Lemma 1.6) and concentration lemma (Lemma 1.7) for general motif counts and general models of random graphs.
We begin by defining exponentially robust quasirandomness for general $\mathcal{D}_{n}$ and $P$. Exponentially robust quasirandomness captures the property that, among multisets of size $s$, the number of copies of $P$ is very well concentrated around its expected value.
Recall the definition of $C_{P}(S)$ from Section 5, for motif $P$ and multiset $S$.
Definition 5.1 (Exponentially robustly quasirandom).
We say that a graph $G=([n],E)$ is $(\varepsilon,\alpha,s_{0})$-exponentially robustly quasirandom with respect to distribution $\mathcal{D}_{n}$ and motif $P$ on $\ell$ vertices if for every $s\geq s_{0}$ we have that
$\Pr_{X_{1},\ldots,X_{s}\sim_{\mathrm{i.i.d.}}[n]}\left[C_{P}(\{X_{1},\ldots,X_{s}\})\not\in(1\pm\varepsilon)S_{\mathcal{D}_{n}}(P,s)\right]\leq 4^{\ell+1}\cdot\exp(-\alpha s).$
We prove the following lemmas. Below, let $\alpha_{\ell},\varepsilon_{\ell},s_{\ell}$ be parameters such from Section 5 and Section 5.
Lemma 5.2 (Inductive Lemma).
Consider a distribution $\mathcal{D}_{n}$ over graphs, and a motif $P$ on $\ell$ vertices. Suppose that for all $j\in[\ell-1]$, $r_{j+1}(P)\leq(4/9)r_{j}(P)$. Let $\mathcal{P}_{\ell-1}$ be the set of (labeled, induced) subgraphs of $P$ on $\ell-1$ vertices.
Suppose that the graph $G$ is $(\varepsilon_{\ell-1},\alpha_{\ell-1},s_{\ell})$-exponentially robustly quasirandom with respect to counts of each $P^{\prime}\in\mathcal{P}_{\ell-1}$ compared to $\mathcal{D}_{n}$.
Define the tester $T_{\ell}$ that samples $s_{\ell}$ (as specified in Section 5) random vertices of $G$ (with replacement) and accepts if and only if the count of $P$ in the induced subgraph on the sampled vertices is in $(1\pm\varepsilon_{\ell-1})S_{\mathcal{D}_{n}}(P,s_{\ell})$. (Recall that, as in Equation 16, $S_{\mathcal{D}_{n}}(P,s_{\ell})$ is the expected number of copies of $P$ in a multiset of $s_{\ell}$ vertices in graphs chosen according to $\mathcal{D}_{n}$.)
$T_{\ell}$ uses $\widetilde{O}\left(r_{\ell}(H)^{-2}\varepsilon^{-4}\right)$ adjacency matrix queries. With probability at least $1-4^{\ell+1}\exp(-\alpha_{\ell}s_{\ell})$, $T_{\ell}$ accepts graphs that are $(\varepsilon_{\ell-1},\alpha_{\ell},s_{\ell})$-exponentially robustly quasirandom with respect to counts of $P$ compared to $\mathcal{D}_{n}$ and rejects graphs that are not $(\varepsilon_{\ell},\alpha_{\ell},s_{\ell})$-exponentially robustly quasirandom with respect to counts of $P$ compared to $\mathcal{D}_{n}$.
The following lemma provides a concentration inequality regarding the probability that the count of $P$ in a random multiset $S$ is far from its expectation in terms of the assumed quasirandomness on smaller sub-motifs and ratio between motif densities.
Lemma 5.3 (Concentration Lemma).
Let $P$ be a motif on $\ell$ vertices and let $s\in\mathbb{N}$. Suppose that graph $G$ is $(\gamma,\beta,s)$-exponentially robustly quasirandom with respect to distribution $\mathcal{D}_{n}$ and all motifs $P^{\prime}\subseteq P$ on $\ell-1$ vertices. Suppose the following three conditions hold:
  1. 1.
    Size of s: $s$ satisfies $\exp(-\beta s)\leq\gamma\cdot F_{\mathcal{D}_{n}}(P)/(16\cdot 9^{\ell})$.
  2. 2.
    Maximum density: $\max_{\begin{subarray}{c}P^{\prime}\subseteq P\\ |P^{\prime}|=\ell-1\end{subarray}}F_{\mathcal{D}_{n}}(P^{\prime})\leq\gamma/32$.
  3. 3.
    Lower bound on density increment: $r_{\ell}(P)\cdot\frac{\gamma^{2}}{4096\cdot\ell^{4}}\geq\kappa$ for some $\kappa>0$.
Then:
$\Pr_{S}\left(\left|C_{P}(S)-C_{P}(G)\cdot A_{s,\ell}\right|\geq\frac{\gamma}{4}S_{\mathcal{D}_{n}}(P,s)\right)\leq 2\cdot 4^{\ell}\exp(-\beta s)+2\exp\left(-\kappa s\right).$
We show that this lemma implies the following corollary, which is applicable for a specific setting of parameters (given by the choice of $P$ and $\mathcal{D}_{n}$), as given in Section 5.
Corollary 5.3.
Let $P$ be a motif on $\ell$ vertices and $\mathcal{D}_{n}$ be a distribution over graphs. Suppose that Conditions 2 and 3 of the Lemma 5.3 hold. Given $P$ and $\mathcal{D}_{n}$, define $s_{\ell}$ as Section 5. For $2\leq\ell\leq k-1$, consider the definitions of $\varepsilon_{\ell}$ and $\alpha_{\ell}$ from Section 5.
Suppose that graph $G$ is $(\varepsilon_{\ell-1},\alpha_{\ell-1},s_{\ell})$-exponentially robustly quasirandom with respect to distribution $\mathcal{D}_{n}$ and all motifs $P^{\prime}\subseteq P$ on $\ell-1$ vertices. Suppose also that for all $j\in\{2,3,\dots,\ell\}$, $r_{j}(P)\leq(4/9)r_{j-1}(P)$.
Then:
$\Pr_{S}\left(\left|C_{P}(S)-C_{P}(G)\cdot A_{s_{\ell},\ell}\right|\geq\frac{\varepsilon_{\ell-1}}{4}S_{\mathcal{D}_{n}}(P,s_{\ell})\right)\leq 4^{\ell+1}\exp(-\alpha_{\ell}s_{\ell}).$
5.1.1 Concentration Lemma
In this section, we prove the general concentration lemma (Lemma 5.3).
Proof of Lemma 5.3
We first prove the following two claims. Let $C_{P}(G)$ be the number of induced labeled copies of $P$ in the graph $G$. For a subgraph $S$ of $G$, define $C_{P}(S)$ similarly.
Claim 5.3.
Let $S$ be a multiset generated by sampling $s$ vertices from the graph, with replacement. Then the expected number of labeled induced copies of $P$ in $S$ is $C_{P}(G)\cdot A_{s,\ell}$.
Proof.
Suppose that $P$ is an $\ell$-vertex motif. Let us first write the number of labeled induced copies of $P$ in $S$ (denoted as $C_{P}(S)$, defined in Section 5) as a sum of indicator random variables. For ordered set $T$ and ordered $P$, let $T\equiv P$ mean that the ordered set of vertices of $T$ is exactly that of $P$. Observe that:
$C_{P}(S)=\sum_{P\in C_{P}(G)}\sum_{T\in S^{\ell}}1(T\equiv P).$
Fix a size $\ell$ subset of $S$ and some $P\in C_{P}(G)$. The probability that $T\equiv P$ is $1/n^{\ell}$, since $T$ is created by sampling $\ell$ vertices, each time uniformly at random (with repetitions).
By linearity of expectation, $\mathbb{E}\left(C_{P}(S)\right)=C_{P}(G)\cdot\binom{s}{\ell}\ell!/n^{\ell}=C_{P}(G)\cdot A_{s,\ell}.$
Claim 5.3.
Suppose that $P$ is a motif on $\ell$ vertices. Consider any two sets $S,S^{\prime}$ such that, for all $P^{\prime}\subseteq P$ on $\ell-1$ vertices, the number of $P^{\prime}$ copies in each of $S$ and $S^{\prime}$ is in the range $C_{P^{\prime}}(S),C_{P^{\prime}}(S^{\prime})\in(1\pm\varepsilon)S_{\mathcal{D}_{n}}(P^{\prime},s)$. Then,
$\left|C_{P}(S)-C_{P}(S^{\prime})\right|\leq|S\triangle S^{\prime}|\cdot(1+\varepsilon)\ell^{2}\cdot\max_{\begin{subarray}{c}P^{\prime}\subseteq P\\ |P^{\prime}|=\ell-1\end{subarray}}S_{\mathcal{D}_{n}}(P^{\prime},s).$
Proof.
Let $C_{P}(v,S)$ be the number of copies of $P$ in $S$ that involve vertex $v\in S$ (and define such a quantity for $S^{\prime}$ similarly). The only copies of $P$ that contribute to the difference $\left|C_{P}(S)-C_{P}(S^{\prime})\right|$ are those involving vertices $u\in S\setminus S^{\prime}$ or vertices $v\in S^{\prime}\setminus S$.
For any vertex $u\in S$ (respectively $S^{\prime}$), the number of unlabeled $P$ copies it contributes to in the graph is upper-bounded by the number of copies of all $P^{\prime}\subseteq P$ on at most $\ell-1$ vertices in $S$ (respectively $S^{\prime}$), which itself is upper-bounded by $\ell$ times the maximum count of any such $P^{\prime}$: $\ell\cdot\max_{\begin{subarray}{c}P^{\prime}\subseteq P\\ |P^{\prime}|=\ell-1\end{subarray}}C_{P^{\prime}}(S)$. Therefore, the number of labeled copies any $u\in S$ contributes to is upper bounded by $\ell^{2}\cdot\max_{\begin{subarray}{c}P^{\prime}\subseteq P\\ |P^{\prime}|=\ell-1\end{subarray}}C_{P^{\prime}}(S)$ (and similarly for $S^{\prime}$).
This yields the following:
$\left|C_{P}(S)-C_{P}(S^{\prime})\right|\leq\ell^{2}\sum_{u\in S\setminus S^{\prime}}\max_{\begin{subarray}{c}P^{\prime}\subseteq P\\ |P^{\prime}|=\ell-1\end{subarray}}C_{P^{\prime}}(u,S)+\ell^{2}\sum_{v\in S^{\prime}\setminus S}\max_{\begin{subarray}{c}P^{\prime}\subseteq P\\ |P^{\prime}|=\ell-1\end{subarray}}C_{P^{\prime}}(v,S^{\prime})$
$\leq|S\setminus S^{\prime}|\cdot\ell^{2}\cdot\max_{\begin{subarray}{c}P^{\prime}\subseteq P\\ |P^{\prime}|=\ell-1\end{subarray}}C_{P^{\prime}}(S)+|S^{\prime}\setminus S|\cdot\ell^{2}\cdot\max_{\begin{subarray}{c}P^{\prime}\subseteq P\\ |P^{\prime}|=\ell-1\end{subarray}}C_{P^{\prime}}(S^{\prime})$
$\leq|S\triangle S^{\prime}|\cdot(1+\varepsilon)\ell^{2}\max_{\begin{subarray}{c}P^{\prime}\subseteq P\\ |P^{\prime}|=\ell-1\end{subarray}}S_{\mathcal{D}_{n}}(P^{\prime},s).$
Proof of Lemma 5.3.
Consider a motif $P$ on $\ell$ vertices with maximum degree $\Delta(P)$. Suppose that $G$ is $(\gamma,\beta,s)$-exponentially robustly quasirandom with respect to distribution $\mathcal{D}_{n}$ and all motifs $P^{\prime}\subseteq P$ on $\ell-1$ vertices. Let $X_{1},\dots,X_{s}$ be the random variables corresponding to which vertices are selected to be in $S$ (i.e., $X_{i}$ is the $i$-th vertex added to $S$, with repetitions).
By Section 5.1 of exponentially robust quasirandomness for $\mathcal{D}_{n}$ and all motifs $P^{\prime}\subseteq P$, $|P^{\prime}|=\ell-1$, for every $s$ and $P^{\prime}\subseteq P$ with $|P^{\prime}|=\ell-1$, we have that
$\Pr\left[C_{P^{\prime}}(X_{1},X_{2},\dots,X_{s})\not\in(1\pm\gamma)S_{\mathcal{D}_{n}}(P,s)\right]\leq 4^{\ell}\exp(-\beta s).$
We will apply McDiarmid’s Inequality when differences are bounded with high probability (Section 2.4) to the function $C_{P}:[n]^{s}\to\mathbb{R}$, which is the number of $P$ copies in the multiset.
Let us set up the relevant definitions for the inequality. First, let the domain be $n]^{s}$, and define $\mathcal{Y}\subseteq[n]^{s}$ to be the multisets of the domain such that, for each $P^{\prime}\subseteq P$ on $\ell-1$ vertices, the number of copies of $P^{\prime}$ in $\mathcal{Y}$ is in $(1\pm\gamma)S_{\mathcal{D}_{n}}(P^{\prime},s)$. Then, observe that exponential-quasirandomness guarantees that $q:=1-\mathbb{P}\left((X_{1},X_{2},\dots,X_{s})\in\mathcal{Y}\right)\leq\ell\cdot 4^{\ell}\exp(-\beta s)$.
We argue about bounded differences of $C_{P}$ on $\mathcal{Y}$. For each $i\in[s]$, define
$c_{i}=(1+\gamma)\ell^{2}\cdot\max_{\begin{subarray}{c}P^{\prime}\subseteq P\\ |P^{\prime}|=\ell-1\end{subarray}}S_{\mathcal{D}_{n}}(P^{\prime},s).$
By Claim 5.1.1, for all $S=(x_{1},x_{2},\dots,x_{s})\in\mathcal{Y}$ and $S^{\prime}=(y_{1},y_{2},\dots,y_{s})\in\mathcal{Y}$,
$\left|C_{P}(S)-C_{P}(S^{\prime})\right|\leq|S\triangle S^{\prime}|\cdot(1+\gamma)\ell^{2}\cdot\max_{\begin{subarray}{c}P^{\prime}\subseteq P\\ |P^{\prime}|=\ell-1\end{subarray}}S_{\mathcal{D}_{n}}(P^{\prime},s)=\sum_{i:x_{i}\neq y_{i}}c_{i},$
so the bounded difference property holds with this setting of $c_{i}$’s.
We want to apply Section 2.4 with error bound $\delta=\gamma S_{\mathcal{D}_{n}}(P,s)/4$. To do so, we need to verify that
$q\leq\min\left\{\frac{\delta}{2\max_{S}(C_{P}(S))},\frac{\delta}{4\mathbb{E}(C_{P}(X_{1},X_{2},\dots,X_{s}))},\frac{1}{2}\right\}.$
Using $\delta=\gamma S_{\mathcal{D}_{n}}(P,s)/4$ and $S_{\mathcal{D}_{n}}(P,s)=F_{\mathcal{D}_{n}}(P)\cdot\binom{n}{\ell}\binom{s}{\ell}\cdot\frac{(\ell!)^{2}}{n^{\ell}}$ from Equation 14 and Equation 16, and since $\max_{S}(C_{P}(S))\leq\binom{s}{\ell}\ell$! and $\mathbb{E}(C_{P}(X_{1},X_{2},\dots,X_{s}))\leq\binom{s}{\ell}\ell$!, it suffices to show that:
$q\leq\min\left\{\frac{\gamma S_{\mathcal{D}_{n}}(P,s)}{8\binom{s}{\ell}\ell!},\frac{\gamma S_{\mathcal{D}_{n}}(P,s)}{16\binom{s}{\ell}\ell!},\frac{1}{2}\right\}=\min\left\{\frac{\gamma}{16}\cdot F_{\mathcal{D}_{n}}(P)\cdot\binom{n}{\ell}\ell!\cdot\frac{1}{n^{\ell}},\frac{1}{2}\right\}.$
Since $q\leq\ell\cdot 4^{\ell}\cdot\exp\left(-\beta s\right)\leq\gamma\cdot F_{\mathcal{D}_{n}}(P)/(16\cdot 2^{\ell})$ by assumption, this inequality holds.
We are ready to apply Section 2.4, which implies that:
$\mathbb{P}\left(\left|f_{P}(X_{1},X_{2},\dots,X_{s})-\mathbb{E}\left(f_{P}(X_{1},X_{2},\dots,X_{s})\right)\right|\geq\gamma S_{\mathcal{D}_{n}}(P,s)/4\right)$
$\leq 2q+2\exp\left(-\frac{2\max\left(0,\gamma S_{\mathcal{D}_{n}}(P,s)/4-q\sum_{i=1}^{s}c_{i}\right)^{2}}{\sum_{i=1}^{s}c_{i}^{2}}\right).$ (18)
Let us analyze the second term. First,
$q\sum_{i=1}^{s}c_{i}\leq\ell\cdot 4^{\ell}\cdot\exp(-\beta s)\cdot s\cdot(1+\gamma)\ell^{2}\cdot\max_{\begin{subarray}{c}P^{\prime}\subseteq P\\ |P^{\prime}|=\ell-1\end{subarray}}S_{\mathcal{D}_{n}}(P^{\prime},s).$
We prove that, given assumptions (2), this is less than $\gamma S_{\mathcal{D}_{n}}(P,s)/8$. Towards showing this, first, since $\max_{\begin{subarray}{c}P^{\prime}\subseteq P\\ |P^{\prime}|=\ell-1\end{subarray}}F_{\mathcal{D}_{n}}(P^{\prime})\leq\gamma/32$, we have:
$\max_{\begin{subarray}{c}P^{\prime}\subseteq P\\ |P^{\prime}|=\ell-1\end{subarray}}F_{\mathcal{D}_{n}}(P^{\prime})\leq\frac{\gamma}{8}\cdot\frac{1}{4}\leq\frac{\gamma}{8}\frac{\binom{n}{\ell}}{\binom{n}{\ell-1}}\cdot\frac{\binom{s}{\ell}}{\binom{s}{\ell-1}}\cdot\frac{(\ell!)^{2}}{((\ell-1)!)^{2}}\cdot\frac{n^{\ell-1}}{n^{\ell}}\cdot\frac{1}{s\cdot\ell^{2}}.$
Equivalently,
$F_{\mathcal{D}_{n}}(P)\cdot s\cdot\ell^{2}\cdot\binom{n}{\ell-1}\binom{s}{\ell-1}\cdot\frac{((\ell-1)!)^{2}}{n^{\ell-1}}\max_{\begin{subarray}{c}P^{\prime}\subseteq P\\ |P^{\prime}|=\ell-1\end{subarray}}F_{\mathcal{D}_{n}}(P^{\prime})\leq\frac{\gamma}{8}\cdot\binom{n}{\ell}\binom{s}{\ell}\cdot\frac{(\ell!)^{2}}{n^{\ell}}F_{\mathcal{D}_{n}}(P).$
Since $\ell\cdot 4^{\ell}\cdot\exp(-\beta s)\leq\gamma F_{\mathcal{D}_{n}}(P)/(16\cdot 9^{\ell})$ and using the definition of $S_{\mathcal{D}_{n}}(P,s)$ and $S_{\mathcal{D}_{n}}(P^{\prime})$, this gives us that:
$\ell\cdot 4^{\ell}\cdot\exp(-\beta s)\cdot\frac{1}{\gamma}\cdot(16\cdot 9^{\ell})\cdot s\cdot\ell^{2}\cdot\max_{\begin{subarray}{c}P^{\prime}\subseteq P\\ |P^{\prime}|=\ell-1\end{subarray}}S_{\mathcal{D}_{n}}(P^{\prime},s)\leq\gamma S_{\mathcal{D}_{n}}(P,s)/8.$
By definition of $q$ and $c_{i}$ for $i\in[s]$, this means that
$q\sum_{i=1}^{s}c_{i}\leq\gamma^{2}S_{\mathcal{D}_{n}}(P,s)/(64\cdot 9^{\ell})\leq\gamma S_{\mathcal{D}_{n}}(P,s)/8.$
Therefore, $2\max\left(0,\gamma S_{\mathcal{D}_{n}}(P,s)/4-q\sum_{i=1}^{s}c_{i}\right)^{2}\geq 2\gamma^{2}S_{\mathcal{D}_{n}}(P,s)^{2}/64=\gamma^{2}S_{\mathcal{D}_{n}}(P,s)^{2}/32$.
This implies that (also rounding $(1+\gamma)S_{\mathcal{D}_{n}}(P,s)$ up to $2S_{\mathcal{D}_{n}}(P,s)$ in the denominator of the expression) Equation 18 is upper-bounded by:
$\leq 2q+2\exp\left(-\frac{\gamma^{2}S_{\mathcal{D}_{n}}(P,s)^{2}}{32\cdot s\cdot(2\ell^{2})^{2}\cdot\max_{\begin{subarray}{c}P^{\prime}\subseteq P\\ |P^{\prime}|=\ell-1\end{subarray}}S_{\mathcal{D}_{n}}(P^{\prime},s)^{2}}\right).$
By definition of $S_{\mathcal{D}_{n}}(P^{\prime},s)$ and $S_{\mathcal{D}_{n}}(P,s)$ (combining Equation 14 and Equation 16), the expression above equals:
$2q+2\exp\left(-\frac{\gamma^{2}}{128\cdot\ell^{4}\cdot s}\cdot\left(\frac{(n-\ell+1)(s-\ell+1)}{n}\cdot\frac{F_{\mathcal{D}_{n}}(P)}{\max_{\begin{subarray}{c}P^{\prime}\subseteq P\\ |P^{\prime}|=\ell-1\end{subarray}}F_{\mathcal{D}_{n}}(P^{\prime},s)}\right)^{2}\right).$
For large enough $n$, and by definition of $s$ (Section 5), we can use that $(n-\ell+1)/n\geq 1/2$ and $(s-\ell+1)\geq s/2$ to obtain that the expression is bounded above by:
$\leq 2q+2\exp\left(-\frac{\gamma^{2}}{128\cdot 16\cdot\ell^{4}}\cdot s\cdot\frac{F_{\mathcal{D}_{n}}(P)^{2}}{\max_{\begin{subarray}{c}P^{\prime}\subseteq P\\ |P^{\prime}|=\ell-1\end{subarray}}F_{\mathcal{D}_{n}}(P^{\prime},s)^{2}}\right).$
By definition of $q$ and $r_{\ell}(P)$ this is:
$\leq 2\cdot 4^{\ell}\exp(-\beta s)+2\exp\left(-\frac{\gamma^{2}}{2048\cdot\ell^{4}}\cdot s\cdot r_{\ell}(P)\right).$
By assumption (3) of the concentration lemma we are proving, this is
$\leq 2\cdot 4^{\ell}\exp(-\beta s)+2\exp\left(-\kappa s\right).$
We now prove Section 5.1 given the concentration lemma proven above.
Proof of Section 5.1.
We apply Lemma 5.3 with $\gamma=\varepsilon_{\ell-1}$, $\beta=\alpha_{\ell-1}$, and $s=_{\ell}$.
First, observe that the assumption that $r_{\ell}(P)\leq(4/9)r_{\ell-1}(P)$ implies that for $\alpha_{\ell}$, $\ell\in[k-1]$ as defined in Section 5, $\alpha_{\ell-1}\geq\alpha_{\ell}$. Next, set $\kappa=r_{\ell}(P)\cdot\frac{\gamma^{2}}{4096\cdot\ell^{4}}$ in Lemma 5.3. Observe that $\kappa\geq\alpha_{\ell}$.
Therefore, we find that:
$2\cdot 4^{\ell}\exp(-\alpha_{\ell-1}s_{\ell})+2\exp(-\kappa s_{\ell})\leq 2\cdot 4^{\ell}\exp(-\alpha_{\ell}s_{\ell})+2\exp(-\alpha_{\ell}s_{\ell})\leq 4^{\ell+1}\exp(-\alpha_{\ell}s_{\ell}).$
Combining this with Lemma 5.3 gives the corollary. ∎
5.1.2 Inductive Lemma
We now prove the general inductive lemma (Lemma 5.2).
Proof of Lemma 5.2 given Lemma 5.3
Towards proving the inductive lemma (Lemma 5.2) from the concentration lemma (Lemma 5.3), we first present some corollaries of Lemma 5.3.
The following corollary proves that if $G$ is $(\varepsilon,\alpha_{\ell-1},s_{\ell})$-exponentially robustly quasirandom with respect to counts of every $P^{\prime}\in\mathcal{P}_{\ell-1}$ compared to $\mathcal{D}_{n}$, then a size-$s_{\ell}$ multiset chosen uniformly at random will, with very high probability, reveal what the global number of $P$ copies in $G$ is and whether $G$ is $(\varepsilon,\alpha_{\ell},s_{\ell})$-exponentially robustly quasirandom with respect to counts of $P$ counts (compared to $\mathcal{D}_{n}$).
Corollary 5.3 (Corollary of Section 5.1).
Consider $2\leq\ell\leq k-1$. Consider a distribution $\mathcal{D}_{n}$ over graphs, and a motif $P$ on $\ell$ vertices, such that for all $j\in[\ell-1]$, $r_{j+1}(P)\leq(4/9)r_{j}(P)$. Let $\mathcal{P}_{\ell-1}$ be the set of (labeled, induced) subgraphs of $P$ on $\ell-1$ vertices.
If a graph $G$ is $(\varepsilon_{\ell-1},\alpha_{\ell-1},s_{\ell})$-exponentially robustly quasirandom with respect to counts of each $P^{\prime}\in\mathcal{P}_{\ell-1}$ compared to $\mathcal{D}_{n}$, then the following hold:
  1. 1.
    Suppose that $(\varepsilon_{\ell-1},\alpha_{\ell},s_{\ell})$-exponentially robustly quasirandom with respect to the count of $P$. Sample a multiset $S$ by selecting $s_{\ell}$ vertices uniformly at random (with repetition). Then with probability at least $1-4^{\ell+1}\exp(-\alpha_{\ell}s_{\ell})$ (over the choice of $S$), $C_{P}(S)\in(1\pm\varepsilon_{\ell-1})S_{\mathcal{D}_{n}}(P,s_{\ell})$.
  2. 2.
    Sample a multiset $S$ by selecting vertices uniformly at random (with repetition). If $C_{P}(S)\in(1\pm\varepsilon_{\ell-1})S_{\mathcal{D}_{n}}(P,s_{\ell})$, then with probability at least $1-4^{\ell+1}\exp(-\alpha_{\ell}s_{\ell})$ (over the choice of $S$), $C_{P}(G)\in(1\pm\frac{5}{4}\varepsilon_{\ell-1})E_{\mathcal{D}_{n}}(P)$ and, furthermore, $G$ is $(\varepsilon_{\ell},\alpha_{\ell},s_{\ell})$-exponentially robustly quasirandom with respect to counts of $P$ compared to $\mathcal{D}_{n}$.
Proof of Section 5.1.2.
Suppose that $G$ is $(\varepsilon_{\ell-1},\alpha_{\ell-1},s_{\ell})$-exponentially robustly quasirandom with respect to counts of each $P^{\prime}\in\mathcal{P}_{\ell-1}$ compared to $\mathcal{D}_{n}$. Then, by Section 5.1 of Lemma 5.3, for each $P^{\prime}\in\mathcal{P}_{\ell-1}$,
$\Pr_{S}\left(\left|C_{P^{\prime}}(S)-C_{P^{\prime}}(G)\cdot A_{s_{\ell},\ell}\right|\geq\frac{\varepsilon_{\ell-1}}{4}S_{\mathcal{D}_{n}}(P,s_{\ell})\right)\leq 4^{\ell+1}\exp(-\alpha_{\ell}s_{\ell}).$
First, let’s address case (1). Suppose $G$ is $(\varepsilon_{\ell-1},\alpha_{\ell},s_{\ell})$-exponentially robustly quasirandom with respect to the count of $P$ compared to $\mathcal{D}_{n}$. By definition of exponentially robust quasirandomness (Section 5.1), all but a $4^{\ell+1}\exp(-\alpha_{\ell}s_{\ell})$ fraction of size-$s$ multisets $S$ satisfy
$C_{P}(S)\in\left(1\pm\varepsilon_{\ell-1}\right)S_{\mathcal{D}_{n}}(P,s).$
Let’s now move to case (2). Section 5.1 implies that when $G$ is $(\varepsilon_{\ell-1},\alpha_{\ell-1},s_{\ell})$-exponentially robustly quasirandom with respect to counts of all $P^{\prime}\in\mathcal{P}_{\ell-1}$ compared to $\mathcal{D}_{n}$, with probability at least $1-4^{\ell+1}\exp\left(-\alpha_{\ell}s_{\ell}\right)$,
$C_{P}(G)\cdot A_{s_{\ell},\ell}\in C_{P}(S)\pm\frac{\varepsilon_{\ell-1}}{4}S_{\mathcal{D}_{n}}(P,s_{\ell}).$
Equivalently, using that $S_{\mathcal{D}_{n}}(P,s_{\ell})=E_{\mathcal{D}_{n}}(P)\cdot\binom{s_{\ell}}{\ell}\frac{1}{n^{\ell}}$ from Equation 16, with probability at least $1-4^{\ell+1}\exp\left(-\alpha_{\ell}s_{\ell}\right)$,
$C_{P}(G)\in C_{P}(S)\cdot\frac{1}{A_{s_{\ell},\ell}}\pm\frac{\varepsilon_{\ell-1}}{4}E_{\mathcal{D}_{n}}(P).$ (19)
Additionally, by Lemma 5.3, under the high-probability event,
$C_{P}(S)\in C_{P}(G)\cdot A_{s_{\ell},\ell}\pm\frac{\varepsilon_{\ell-1}}{4}S_{\mathcal{D}_{n}}(P,s_{\ell}).$ (20)
Suppose that $C_{P}(S)\in(1\pm\varepsilon_{\ell-1})S_{\mathcal{D}_{n}}(P,s_{\ell})$. Therefore,
$C_{P}(S)\cdot\frac{1}{A_{s_{\ell},\ell}}\in(1\pm\varepsilon_{\ell-1})E_{\mathcal{D}_{n}}(P).$
By Equation 19, with probability at least $1-4^{\ell+1}\exp\left(-\alpha_{\ell}s_{\ell}\right)$, $C_{P}(G)\in(1\pm\frac{5}{4}\varepsilon_{\ell-1})E_{\mathcal{D}_{n}}(P)$. We now argue that the graph is exponentially robustly quasirandom with high probability. When $C_{P}(G)\in(1\pm\frac{5}{4}\varepsilon_{\ell-1})E_{\mathcal{D}_{n}}(P)$, by Equation 20, all but a $4^{\ell+1}\exp(-\alpha_{\ell}s_{\ell})$ fraction of size-$s_{\ell}$ multisets $S^{\prime}$ satisfy
$C_{P}(S^{\prime})\in\left(1\pm\frac{3}{2}\varepsilon_{\ell-1}\right)S_{\mathcal{D}_{n}}(P,s_{\ell}).$
Since $\varepsilon_{\ell}=\frac{3}{2}\varepsilon_{\ell-1}$, therefore $G$ is $(\varepsilon_{\ell},\alpha_{\ell},s_{\ell})$-exponentially robustly quasirandom with respect to the count of $P$ compared to $\mathcal{D}_{n}$. ∎
Proof of Lemma 5.2.
Given Section 5.1.2, Lemma 5.2 follows from a simple argument, given below.
Suppose that $G$ is $(\varepsilon_{\ell-1},\alpha_{\ell-1},s_{\ell})$-exponentially robustly quasirandom with respect to the counts of all motifs $P^{\prime}\in\mathcal{P}_{\ell-1}$. Define the tester $T_{\ell}$ that samples $s_{\ell}$ (as specified in Section 5) random vertices of $G$ (with replacement) and accepts if and only if the clique count in the induced subgraph on the sampled vertices is within $(1\pm\varepsilon_{\ell-1})S_{\mathcal{D}_{n}}(P,s_{\ell})$.
By the corollary Section 5.1.2 of the concentration lemma (Lemma 5.3), the tester $T_{\ell}$ satisfies the following properties.
Completeness: Suppose that $G$ is $(\varepsilon_{\ell-1},\alpha_{\ell},s_{\ell})$-exponentially robustly quasirandom with respect to counts of all $P^{\prime}\in\mathcal{P}_{\ell-1}$ compared to $\mathcal{D}_{n}$. Then, with probability at least $1-4^{\ell+1}\exp(-\alpha_{\ell}s_{\ell})$, $C_{P}(S)\in(1\pm\varepsilon_{\ell-1})S_{\mathcal{D}_{n}}(P,s_{\ell})$.
Soundness: If the clique count on the sampled multiset is in $(1\pm\varepsilon_{\ell-1})S_{\mathcal{D}_{n}}(P,s_{\ell})$, then with probability at least $1-4^{\ell+1}\exp(-\alpha_{\ell}s_{\ell})$, $G$ is $(\varepsilon_{\ell},\alpha_{\ell},s_{\ell})$-exponentially robustly quasirandom with high probability, so accepted graphs satisfy the quasirandomness condition. ∎
5.2 Proof of Theorem 5.1
We now prove Theorem 5.1 from the general concentration and inductive lemmas proven above, following the framework of the proof of Theorem 3.1 from the concentration and inductive lemmas for $k$-cliques.
We begin with the following corollary that extends Section 3.1.3.
Corollary 5.3 (Corollary of Lemma 5.3).
Consider $2\leq\ell\leq k-1$. Consider a distribution $\mathcal{D}_{n}$ over graphs, and a motif $H$ on $k$ vertices. Let $\mathcal{H}_{\ell-1}$ be the set of (labeled, induced) subgraphs of $H$ on $\ell-1$ vertices.
Suppose $G$ is $(\varepsilon,\alpha_{k-1},s_{*})$-exponentially robustly quasirandom with respect to counts of each $H^{\prime}\in\mathcal{H}_{k-1}$ compared to $\mathcal{D}_{n}$, and $C_{H}(G)\in(1\pm\frac{1}{4}\varepsilon)E_{\mathcal{D}_{n}}(H)$. Sample a multiset $S$ by selecting $s_{*}$ vertices uniformly at random (with repetition). Then with probability at least $1-4^{k+1}\exp(-\alpha_{k}s_{*})$, $C_{H}(S)\in(1\pm\frac{1}{2}\varepsilon)S_{\mathcal{D}_{n}}(H,s_{*})$. If $C_{H}(G)\not\in(1\pm\varepsilon)E_{\mathcal{D}_{n}}(H)$, then with probability at least $1-4^{k+1}\exp(-\alpha_{k}s_{*})$, $C_{H}(S)\not\in(1\pm\frac{3}{4}\varepsilon)S_{\mathcal{D}_{n}}(H,s_{*})$.
Proof.
The proof follows similarly to that of Section 5.1.2. ∎
Next, we show that Claim 3.1.3 from in Section 3 extends rather directly to the setting of motifs and general graph distributions. Interestingly, the fact that the proof highlights that all we needed about $\mathcal{G}_{n,p}$ previously was the concentration of $\ell$-clique counts for all $\ell\in[k]$.
Claim 5.3.
Consider $2\leq\ell\leq k-1$. Consider a $\mathcal{D}_{n}$ over graphs on $n$ vertices and a motif $H$ on $k$ vertices, such that for all $\ell\in[k-1]$, $r_{\ell+1}(H)\leq(4/9)r_{\ell}(H)$. Let $\mathcal{H}_{k-1}$ be the set of (labeled, induced) subgraphs of $H$ on $\ell-1$ vertices. Define $s_{\ell}$, $\ell\in[k]$, as in Section 5.
Suppose that a graph $G$ satisfies the property that for all $\ell\in\{1,2,\dots,k\}$ the number of induced labeled copies of each $H^{\prime}\in\mathcal{H}_{\ell}$ in $G$ is in the range $(1\pm\frac{\varepsilon}{4(k+1)})E_{\mathcal{D}_{n}}(H^{\prime})$. Suppose also that, for all $\ell\in[k-1]$, $r_{\ell}(H)\geq r_{\ell+1}(H)$.
Then, for size $s_{\ell}$ multisets, for all $\ell\in[k-1]$, $G$ is $\left((\ell+1)\cdot\frac{\varepsilon}{4(k+1)},(\ell+1)^{2}\cdot\frac{\varepsilon^{2}}{16(k+1)^{2}}\cdot\frac{r_{\ell}(H)}{2048\cdot\ell^{2}},s_{\ell}\right)$-exponentially robustly quasirandom with respect to the count of all $H^{\prime}\in\mathcal{H}_{\ell}$.
Proof.
Since, for all $\ell\in[k-1]$, $(4/9)r_{\ell}(H)\geq r_{\ell+1}(H)$, then $s_{\ell+1}\geq s_{\ell}$.
The proof of this is then a direct extension of the proof of Claim 3.1.3, which is a proof by induction on $\ell$, using a repeated application of the concentration lemma. ∎
We are ready to prove Theorem 5.1, along the lines of the proof of Theorem 3.1 in Section 3.1.
Proof of Theorem 5.1.
Completeness. By assumption, with high probability $G\sim\mathcal{D}_{n}$ has concentration of motif counts for all $H^{\prime}\subseteq H$, where $H$ is the motif on $k$ vertices that we are studying the count of. In the high-probability event that $G\sim\mathcal{D}_{n}$ is concentrated, Claim 5.2 implies that for all $\ell\in[k]$ and $H^{\prime}\subseteq H$ on $\ell$ vertices, $G$ is $\left(\delta,\nu,s_{\ell}\right)$-exponentially robustly quasirandom with respect to the count of $H^{\prime}$, for
$\delta_{\ell}=(\ell+1)\cdot\frac{\varepsilon}{4(k+1)},~~\nu_{\ell}=(\ell+1)^{2}\cdot\frac{\varepsilon^{2}}{16(k+1)^{2}}\cdot\frac{r_{\ell}(H)}{2048\cdot\ell^{2}}.$
From this, Section 5.1.2 implies that with probability at least $1-k\cdot 4^{k}\exp\left(-\nu_{k}s_{*}\right)\geq 9/10$, the uniformly sampled multiset $S$ satisfies $C_{H^{\prime}}(S)\in(1\pm\frac{\varepsilon}{4})S_{\mathcal{D}_{n}}(H^{\prime},s_{*})$ for all $H^{\prime}\subseteq H$. Therefore, the algorithm accepts $G\sim\mathcal{D}_{n}$ with high constant probability.
Soundness. If there exists some $\ell\in\{2,3,\dots,k-1\}$ and $H^{\prime}\subseteq H$ on $\ell$ vertices such that the input graph $G$ is $(\varepsilon_{\ell^{\prime}},\alpha_{\ell^{\prime}},s_{\ell^{\prime}})$-exponentially robustly quasirandom with respect to the count of $H^{\prime\prime}\subseteq H,|H^{\prime\prime}|=\ell^{\prime}$ for all $\ell^{\prime}<\ell$ and is not $(\varepsilon_{\ell},\alpha_{\ell},s_{\ell})$-exponentially robustly quasirandom with respect to the count of $H^{\prime}$, then Lemma 5.2 implies that with probability at least $1-4^{\ell+1}\exp(-\alpha_{\ell}s_{\ell})\geq 9/10$, the algorithm rejects on $G$.
If, instead, the input graph is $(\varepsilon_{k-1}=\varepsilon,\alpha_{k-1},s_{k}=s_{*})$-exponentially robustly quasirandom with respect to the count of all $H^{\prime}\subseteq H,|H^{\prime}|=k-1$. In this case, Section 5.2 implies that with probability at least $1-4^{k+1}\exp(-\alpha_{k}s_{*})$, $C_{H}(S)\not\in(1\pm\frac{3}{4}\varepsilon)S_{\mathcal{D}_{n}}(H,s_{*})$, in which case the algorithm rejects.
Thus, the algorithm rejects each $G$ with $C_{H}(G)\not\in(1\pm\varepsilon)E_{\mathcal{D}_{n}}(H)$ with high probability.
Query complexity: The adjacency matrix query complexity is $\binom{s_{*}}{2}=\widetilde{O}\left(r_{k}(H)^{-2}\varepsilon^{-4}\right)$.
Runtime: The algorithm counts the number of labeled induced copies of $H^{\prime}$, for all $H^{\prime}\subseteq H$. For worst-case input graphs and arbitrary distributions $\mathcal{D}_{n}$ over graphs, this can take up to time $O(s_{*}^{k})=\widetilde{O}\left(r_{k}(H)^{-2k}\varepsilon^{-4k}\right)$. ∎
Appendix A Quality control of triangle counts
We prove that, under the adjacency list query access model (as defined in Section 2), quality control of triangle counts as compared to $\mathcal{G}_{n,p}$ requires $\Theta\left(\frac{1}{p}\right)$ queries, up to $\mathrm{polylog}(n)$ and $\mathrm{poly}(1/\varepsilon)$ multiplicative factors.
Recall that, for a graph $G$, $C_{3}(G)$ is the number of triangles in the graph.
Theorem A.1.
Let $\rho_{\Delta}=C_{3}(G)/\mathbb{E}_{G^{\prime}\sim\mathcal{G}_{n,p}}\left[C_{3}(G^{\prime})\right]$ and consider parameters $n,p,\varepsilon$ such that $p=\Omega\left(\varepsilon^{-2}\cdot 1/n\right)$. Solving the $(\mathcal{G}_{n,p},\rho_{\Delta},\varepsilon)$-quality control problem can be accomplished in $O\left(\frac{1}{p}\right)\cdot\mathrm{poly}\left(\log(n),1/\varepsilon\right)$ adjacency matrix, adjacency list, and vertex degree queries and $O\left(\frac{1}{p}\right)\cdot\mathrm{poly}\left(\log(n),1/\varepsilon\right)$ runtime, in expectation (over the random choices of the algorithm). Moreover, $(\mathcal{G}_{n,p},\rho_{\Delta},\varepsilon)$-quality control requires $\Omega\left(\frac{1}{p}\right)$ queries.
Studying triangles is interesting because, beyond being one of the most elementary structures possible, the algorithm we construct will directly rely on results from the literature [ERS20a, EMR22]. Studying triangles will, therefore, set us up for understanding the limitations we would face in applying the approaches from the literature to the context of quality control for larger motifs.
A.1 Query complexity upper bound
In this section, we prove the upper bound in Theorem A.1, restated below.
Lemma A.2 (Upper bound in Theorem A.1).
Consider parameters $n,p,\varepsilon$ such that $p=\Omega\left(\varepsilon^{-2}\cdot 1/n\right)$. Solving the $(\mathcal{G}_{n,p},\rho_{\Delta},\varepsilon)$-quality control problem can be accomplished in $O\left(\frac{1}{p}\right)\cdot\mathrm{poly}\left(\log(n),1/\varepsilon\right)$ adjacency matrix, adjacency list, and vertex degree queries and $O\left(\frac{1}{p}\right)\cdot\mathrm{poly}\left(\log(n),1/\varepsilon\right)$ runtime, in expectation (over the random choices of the algorithm).
To prove this, we rely on and combine algorithms from the literature [EMR22, ERS20a] to obtain a quality control algorithm for triangle counts that uses $O(1/p)\cdot\text{poly}(\log n,\frac{1}{\varepsilon})$ queries. At a high level, the algorithm does the following. The algorithm will first approximate the arboricity; if the approximation is less than $2np$, proceed, and else reject. Next, approximate the number of triangles using the sublinear approximation algorithm of [ERS20a] whose query complexity depends on the arboricity of the input graph. Accept if the approximation is in the range $(1\pm\varepsilon/2)\binom{n}{3}p^{3}$; else, reject. The completeness (accepting $G\sim\mathcal{G}_{n,p}$ with high probability) relies on the concentration of the arboricity and number of triangles for Erdős–Rényi graphs (see Section 2.3 and Section 2.3). The soundness follows because we reject all graphs with $\rho_{\Delta}\not\in(1\pm\varepsilon)\mathbb{E}_{G^{\prime}\sim\mathcal{G}_{n,p}}\left[\rho_{\Delta}(G^{\prime})\right]$ (and possibly other graphs, due to the nature of the quality control framework).
Towards this general goal, we present subroutines from [EMR22, ERS20a] that provide “decision versions” of the problem of estimating the arboricity and triangle count. That is, the following lemmas state that it is possible to detect when the arboricity and triangle count are in a certain range, versus sufficiently far from the range.
Lemma A.3 (Claim 3.23 of [EMR22]).
Consider an input graph $G$ with arboricity $\mathrm{arb}(G)$, and a parameter $\alpha$. There exists an algorithm (call it Algorithm A) that performs $O(n/\alpha)$ adjacency list and degree queries and has runtime $O(n/\alpha)$ with the following guarantees. If $\mathrm{arb}(G)\leq\alpha$, then with probability at least $1-1/n^{3}$, the algorithm accepts. If $\mathrm{arb}(G)>100\log^{2}(n)\cdot\alpha$, then with probability at least $1-20\log(n)/n^{4}$, the algorithm rejects.
We will apply this with $\alpha=\Theta\left(np\right)$, so the query complexity of the algorithm is $O(1/p)$.
We also observe that deciding if the average degree is $O(np)$ can be done efficiently, with high probability.
Lemma A.4 ([Fei06], simplified).
There is an algorithm that uses $O\left(\sqrt{n/d}\right)$ degree queries and time and estimates the average degree of an input graph $G$ up to a multiplicative factor of $3$, with high constant probability, as long as $G$ has average degree at least $d$.
We rely on the following corollary, which turns this into a decision version of the problem.
Corollary A.4.
Consider an input graph $G$ with $e(G)$ edges, and a parameter $\beta$. There exists an algorithm (call it Algorithm B) that performs expected $O(\log(n)/p)$ queries and has expected runtime $O(\log(n)/p)$ with the following guarantees. If $e(G)\in\left[\frac{1}{2}n^{2}p,2n^{2}p\right]$, the algorithm accepts with high constant probability. If $e(G)<\frac{1}{5}n^{2}p$ or $e(G)>10n^{2}p$, the algorithm rejects with high constant probability.
Remark A.4.
Let $\rho_{e}(G)=e(G)/\mathbb{E}_{G^{\prime}\sim\mathcal{G}_{n,p}}\left[e(G^{\prime})\right]$. While the focus in this section is on triangles, a modification of this corollary already implies a $(\mathcal{G}_{n,p},\rho_{e})$-quality control algorithm for comparing the number of edges to $\mathcal{G}_{n,p}$, with $\widetilde{O}(\sqrt{1/p})\cdot\text{poly}(1/\varepsilon)$ adjacency list queries and runtime. Particularly, [GR08] provides stronger guarantees than [Fei06], giving a $\widehat{O}(n^{2}/e(G))^{1/2}\cdot\text{poly}(1/\varepsilon)$-time algorithm that outputs a $(1\pm\varepsilon)$-approximation of the average degree in a graph $G$. We can use this to prove a more refined version of Section A.1. For quality control, the completeness case follows from the concentration of the number of edges for $G^{\prime}\sim\mathcal{G}_{n,p}$. Soundness follows from rejecting graphs whose edge count is out of $(1\pm\varepsilon)\binom{n}{2}p$.
Proof of Section A.1.
Consider the algorithm that does the following: first it samples $s=O(\log(n)/p)$ random vertices’ degrees and rejects if the sum is less than $\frac{1}{5}np$. Then, it runs the procedure from [Fei06] (Lemma A.4).
First, we can test if $e(G)<\frac{1}{10}np$ with the first step of the algorithm. Suppose we sample a multiset of $s=O(\log(n)/p)$ vertices. Let $X_{i}$ be the degree of the $i$th vertex chosen. The expectation of $X_{i}$ is $e(G)/n<\frac{1}{10}np$. Observe that, since $\max(X_{i})\leq n$, $\mathbb{E}[X_{i}^{2}]\leq n\mathbb{E}[X_{i}]<\frac{1}{10}n^{2}p$.
By Bernstein’s inequality, the probability that the sum of the $s$ random vertices’ degrees is at most $\frac{1}{5}nps$ is at most
$\exp\left(-\frac{\frac{1}{2}\left(\frac{1}{5}nps\right)^{2}}{\frac{1}{10}sn^{2}p+\frac{1}{3}\frac{1}{5}n^{2}ps}\right)=\exp\left(-Csp\right),$
for some constant $C$. For $s=O(\log(n)/p)$, this is at most $1/n^{2}$. Therefore, if $e(G)<\frac{1}{5}n^{2}p$, the algorithm rejects with high probability.
If the first step does not reject, we run the second step, i.e. the procedure from [Fei06]. If the input graph has $e(G)\geq\frac{1}{5}n^{2}p$, by Lemma A.4 this step takes $O(\sqrt{1/p})$ time. Otherwise, the algorithm can take $O(n^{2})$ time; this will happen with probability at most $1/n^{2}$. The algorithm has expected $O(\sqrt{1/p})$ queries and has expected runtime $O(\sqrt{1/p})$. This procedure has the guarantee that if $e(G)\in\left[\frac{1}{2}n^{2}p,2n^{2}p\right]$, the algorithm accepts with high constant probability, and if $e(G)>10n^{2}p$, the algorithm rejects with high probability. ∎
For a graph $G$, let $C_{3}(G)$ be the number of triangles in the graph. In [ERS20a], the authors construct an algorithm for estimating the number of $k$-cliques in a graph with low arboricity. Their algorithm repeatedly calls an oracle-based algorithm, which takes in an estimate for the number of $k$-cliques and returns different output values depending on the relationship of the number of $k$-cliques in $G$ to this estimate. In the following two lemmas, we state the oracle versions of the results from this paper, which we will need for our context of quality control.
The following two lemmas are from [ERS20a], presented in a simplified form focused on the case of triangle counting.
Lemma A.5 (Lemma 7.1, [ERS20a]).
Consider an input graph $G$ (with $n$ vertices and $m$ edges) with $C_{3}(G)$ triangles, and parameters $\alpha^{\prime}$, $t$, $\delta$, $\gamma$. There exists an algorithm (call it Algorithm C) with the following guarantees, when $\mathrm{arb}(G)\leq\alpha^{\prime}$:
  1. 1.
    If $C_{3}(G)<t$, then with probability at least $1-\gamma$, the algorithm outputs a number $\widehat{t}$ that is less than $t$.
  2. 2.
    If $t\in[C_{3}(G)/4,C_{3}(G)]$, then with probability at least $1-\gamma$, the algorithm outputs a number $\widehat{t}$ that is in the range $(1\pm\delta)C_{3}(G)$.
The algorithm performs adjacency list, adjacency matrix, and degree queries. The expected query complexity and runtime of the algorithm are:
$O\left(\min\left\{\frac{n\cdot(\alpha^{\prime})^{2}}{t},\frac{n}{C_{3}(G)^{1/3}}\right\}+\frac{m\cdot\alpha^{\prime}}{t}\cdot\frac{C_{3}(G)}{t}\right)\cdot\mathrm{poly}\left(\log(n/\gamma),1/\delta\right).$
Lemma A.6 (Corollary 7.2, [ERS20a]).
Consider an input graph $G$ (with $n$ vertices and $m$ edges) with $C_{3}(G)$ triangles, and parameters $\alpha^{\prime}$ and $t$. There exists an algorithm (call it Algorithm D) with the following guarantee.
When $\mathrm{arb}(G)\leq\alpha^{\prime}$, if $C_{3}(G)\geq t$, then with high constant probability, the algorithm outputs a number $\widehat{t}$ that is in the range $(1\pm\delta)C_{3}(G)$.
The algorithm performs adjacency list, adjacency matrix, and degree queries. The expected query complexity and runtime of the algorithm are:
$O\left(\min\left\{\frac{n\cdot(\alpha^{\prime})^{2}}{t},\frac{n}{C_{3}(G)^{1/3}}\right\}+\frac{m\cdot\alpha^{\prime}}{t}\right)\cdot\mathrm{poly}\left(\log(n),1/\delta\right).$
We describe the algorithm below.
Algorithm Triangle-Quality:
On inputs $G$, $n,p,\varepsilon$, and $\mathcal{G}_{n,p}$
  1. 1.
    Run Algorithm A (from Lemma A.2) with (arboricity bound) parameter $\alpha=2np$. If Algorithm A rejects, then reject.
  2. 2.
    Run Algorithm B (from Section A.1). If Algorithm B rejects, then reject.
  3. 3.
    Run Algorithm C (from Lemma A.5) with parameters $\alpha^{\prime}=200\log^{2}(n)\cdot np$ (arboricity bound), $\beta=10np$ (edge bound), $t=(1-\varepsilon)\binom{n}{3}p^{3}$, and estimation accuracy $\delta=\varepsilon/4$. If the output $\widehat{t}$ is <t<t, then reject. If the algorithm has not terminated after $O\left(\frac{1}{p}\right)\cdot\mathrm{poly}\left(\log(n),1/\varepsilon\right)$ steps, then reject.
  4. 4.
    Run Algorithm D (from Lemma A.6) with parameters $\alpha^{\prime}=200\log^{2}(n)\cdot np$ (arboricity bound), $\beta=10np$ (edge bound), and $t=(1-\varepsilon)\binom{n}{3}p^{3}$, and estimation accuracy $\delta=\varepsilon/4$. If the output $\widehat{t}$ is $(1+\frac{\varepsilon}{2})\binom{n}{3}p^{3}$, then reject. If the algorithm has not terminated after $O\left(\frac{1}{p}\right)\cdot\mathrm{poly}\left(\log(n),1/\varepsilon\right)$ steps, then reject.
  5. 5.
    Otherwise, accept the graph.
We now prove that this algorithm implies Lemma A.2.
Proof.
Completeness: We first reason about completeness, i.e., that Algorithm 1 accepts $G\sim\mathcal{G}_{n,p}$ with high probability (both over the distribution over graphs and the algorithm’s randomness). By Section 2.3 and Section 2.3, with high constant probability, since $p=\Omega\left(\varepsilon^{-2}\cdot 1/n\right)$, $G\sim\mathcal{G}_{n,p}$ satisfies $\mathrm{arb}(G)\leq 2np$, $m\in(1\pm\frac{\varepsilon}{100})\binom{n}{2}p$, and $C_{3}(G)\in(1\pm\frac{\varepsilon}{100})\binom{n}{3}p^{3}$.
We first argue that, when $G\sim\mathcal{G}_{n,p}$, all algorithms called as subroutines are able to terminate successfully (and therefore we can utilize their high-probability guarantees). To see this, first note that, given the high-probability bounds on the edge count, triangle count, and arboricity of $G\sim\mathcal{G}_{n,p}$, all algorithms called as subroutines run in $O\left(\frac{1}{p}\right)\cdot\mathrm{poly}\left(\log(n),1/\varepsilon\right)$ expected time. By Markov’s inequality, with probability at least $1-\frac{1}{100}$ the runtime of the algorithm is at most $100$ times its expected amount, and so with probability at least $1-\frac{1}{100}$, the runtime of the algorithm for $G$ with $\mathrm{arb}(G)\leq 2np$, $m\in(1\pm\frac{\varepsilon}{100})\binom{n}{2}p$, and $C_{3}(G)\in(1\pm\frac{\varepsilon}{100})\binom{n}{3}p^{3}$ is $O\left(\frac{1}{p}\right)\cdot\mathrm{poly}\left(\log(n),1/\varepsilon\right)$. By analyzing conditional probabilities, we see that the probability of each step outputting its guarantees given the runtime is at most this amount is the original probability of success, minus $\frac{1}{100}$.
Thus, we can assume that the algorithmic subroutines called all terminate for $G\sim\mathcal{G}_{n,p}$. With this in mind, we proceed to the correctness of the steps of the algorithm. When the conditions (edge count, triangle count, arboricity) are satisfied, first, for Steps 1 and 2, by the correctness of Algorithms A and B, with high probability Algorithms A and B both accept. Next, in Step 3, since $\frac{1}{4}C_{3}(G)\leq(1-\varepsilon)\binom{n}{3}p^{3}\leq C_{3}(G)$ for $\varepsilon\leq 3/(4-1/100)$, Algorithm C outputs
$\widehat{t}\in\left(1\pm\delta\right)C_{3}(G)\in\left(1\pm\frac{\varepsilon}{4}\right)\left(1\pm\frac{\varepsilon}{100}\right)\binom{n}{3}p^{3}\in(1\pm\varepsilon)\binom{n}{3}p^{3}.$
So, the output of Algorithm C is greater than $t$. Next, in Step 4, Algorithm D outputs a $\widehat{t}$ satisfying:
$\widehat{t}\leq\left(1+\frac{\varepsilon}{4}\right)\left(1+\frac{\varepsilon}{100}\right)\binom{n}{3}p^{3}.$
Since the algorithm does not reject in any of Steps 1 through 3, it accepts the graph in Step 4, with high probability over $G\sim\mathcal{G}_{n,p}$.
Soundness: We now argue soundness, i.e., that Algorithm 1 rejects graphs $G$ such that $C_{3}(G)\not\in(1\pm\varepsilon)\binom{n}{3}p^{3}$. First, if the graph additionally satisfies $\mathrm{arb}(G)>200\log^{2}(n)\cdot np$, Algorithm A rejects with high probability. If the graph satisfies $e(G)<\frac{1}{5}n^{2}p$ or $e(G)>20n^{2}p$, Algorithm B rejects with high probability. Otherwise, if $C_{3}(G)<(1-\varepsilon)\binom{n}{3}p^{3}$, Algorithm C outputs a $\widehat{t}<t$ in expected runtime and query complexity $O\left(\frac{1}{p}\right)\cdot\mathrm{poly}\left(\log(n),1/\varepsilon\right)$ and the algorithm rejects. If $C_{3}(G)>(1+\varepsilon)\binom{n}{3}p^{3}$, then Algorithm D outputs
$\widehat{t}>\left(1-\frac{\varepsilon}{4}\right)\left(1+\varepsilon\right)\binom{n}{3}p^{3}>\left(1-\frac{\varepsilon}{2}\right),$
for $\varepsilon<1/2$, in expected runtime and query complexity $O\left(\frac{1}{p}\right)\cdot\mathrm{poly}\left(\log(n),1/\varepsilon\right)$ and the algorithm rejects. Therefore, the algorithm rejects all $G$ such that $C_{3}(G)\not\in(1\pm\varepsilon)\binom{n}{3}p^{3}$, with high probability over the randomness of the algorithm.
Query complexity and runtime: The query complexity and runtime of Step 1 are $O(1/p)$, from Lemma A.3. Next, by definition of the algorithm, Steps 2 and 3 terminate after $O\left(\frac{1}{p}\right)\cdot\mathrm{poly}\left(\log(n),1/\varepsilon\right)$ each. Therefore, we obtain the stated query complexity and runtime. ∎
Remark A.6.
The algorithm above only used the following properties of $\mathcal{G}_{n,p}$: with high constant probability, the arboricity is bounded, and the number of edges and triangles are each concentrated around their expected values. Therefore, the algorithm and analysis extend to any distribution over graphs with these properties. Indeed, this is a recurring theme in our results: we utilize minimal structural properties about $\mathcal{G}_{n,p}$ and the results extend nicely to other distributions over graphs that satisfy these properties.
A.2 Query complexity lower bound
Recall that $\rho_{\Delta}(G)$ counts the normalized fraction of triangles in $\mathcal{G}_{n,p}$, specifically given by $\rho_{\Delta}(G):=C_{K_{3}}(G)/(n(n-1)(n-2)p^{3}/6)$. In this section we prove that the $(\mathcal{G}_{n,p},\rho_{\Delta})$-quality control problem, i.e., the quality control problem corresponding to counting the number of triangles in $\mathcal{G}_{n,p}$, requires $\Omega(1/p)$ queries when given access to (sorted) adjacency lists of vertices.
Theorem A.7 (Lower bound in Theorem A.1).
There exists a constant $c>0$ and a polynomially growing function $n_{0}$ such that for every $p\in(0,1/2]$ and every $n\geq n_{0}(1/p)$ if an an algorithm solves the $(\mathcal{G}_{n,p},\rho_{\Delta})$-quality control problem with $q(n,p)$ queries given (sorted) adjacency list access to the input graph, then $q(n,p)\geq c/p$.
A comparable result when given access to the adjacency matrix was given in Section 4.2; here, the focus is instead on adjacency list access, to complement the upper bound setting for triangle counts. We remark that the sorted adjacency list access model is a very powerful model for access to a graph. In particular, queries to the adjacency matrix can be answered using $O(\log n)$ queries to the sorted adjacency list. We further remark that our proof even allows “degree” queries for unit cost.
To prove Theorem A.7 we use two distributions — a YES distribution, which is forced to be $\mathcal{G}_{n,p}$ by the quality control problem having $\mathcal{G}_{n,p}$ as the target distribution, and a NO distribution which is simply $\mathcal{G}_{n,p}$ with a planted clique of size $\ell=O(pn)$. The technical results show that (1) the two distributions are indistinguishable to algorithms making $o(n/\ell)=o(1/p)$ queries and (2) the distributions differ significantly in triangle counts. Together these results imply the query lower bound on quality control for triangle counting with respect to $\mathcal{G}_{n,p}$.
For a graph $G$ on vertex set $[n]$ and a set $S\subseteq[n]$, we define $G^{\prime}(G,S)$ to be the graph on vertex set $[n]$ whose edges are the union of edges of $G$ and of the complete graph on vertex set $S$. In what follows we set $D_{Y}=D_{Y}^{n}=\mathcal{G}_{n,p}$ and $D_{N}=D_{N}^{n,\ell}$ to be the distribution on graphs $G^{\prime}(G,S)$ where $G\sim\mathcal{G}_{n,p}$ and $S\subseteq[n]$ with $|S|=\ell$ is chosen uniformly (independent of $G$).
Lemma A.8.
Let $A$ be a deterministic algorithm making $q$ queries to the (sorted) adjacency list of an $n$ vertex graph and outputting $0/1$. Then
$\left|\Pr_{G\in D_{Y}^{n}}[A(G)=1]-\Pr_{G^{\prime}\in D_{N}^{n,\ell}}[A(G^{\prime})=1]\right|\leq\frac{2q\ell}{n-\ell}.$
Proof.
We prove the following stronger statement that immediately implies the lemma. For every graph $G$
$\Pr_{S}[A(G^{\prime}(G,S))\neq A(G)]\leq\frac{2q\ell}{n-\ell}.$
Given $G$, let $\{(u_{j},i_{j})\}_{j=1}^{q}$ denote the queries of $A$ on input $G$ (where the response to the query $(u_{j},i_{j})$ includes the degree of $u_{j}$, the name $v_{j}$ of the $i_{j}$th neighbor of $u_{j}$ and the degree of $v_{j}$). Let $Q=\{u_{j}|j\in[q]\}\cup\{v_{j}|j\in[q]\}$. Then we have $\Pr_{S}[S\cap Q\neq\emptyset]\leq\frac{2q\ell}{n-\ell}$. Conditioned on $S\cap Q=\emptyset$ we have that every query of $A$ to $G^{\prime}$ gets the same answers as the corresponding query in $G$, and so $A(G^{\prime})=A(G)$. This proves the lemma. ∎
Lemma A.9.
For every $p\in(0,1/2]$ and $n\geq 10p^{-1}\varepsilon^{-2}$, and every $\ell\geq 2np$, the following hold:
$\Pr_{G\sim D_{Y}^{n}}[|\rho_{\Delta}(G)-1|\leq 0.1]\geq 1-1/12,$
 $\mbox{ and }\Pr_{G^{\prime}\sim D_{N}^{n,\ell}}[\rho_{\Delta}(G^{\prime})\geq 1.9]\geq 1-1/12.$
Proof.
First, the expected number of triangles in $\mathcal{G}_{n,p}$ is $\mu_{3}=\binom{n}{3}p^{3}$. The number of triangles is concentrated in $\mathcal{G}_{n,p}$ given $n\geq 10p^{-1}\varepsilon^{-2}$, by Section 2.3. This implies $\Pr_{G\sim D_{Y}^{n}}[|C_{3}(G)-\mu_{3}|\leq 0.1\cdot\mu_{3}]=1-o_{n}(1)\geq 1-1/12$.
Second, for $G^{\prime}\sim D_{N}^{n,\ell}$, the number of triangles is at least $2\binom{n}{3}p^{3}$ with high probability. Suppose the planted clique is on $S$. Then, the graph on $n]\setminus S$ is an Erdős–Rényi graph itself, and has at least $0.9\binom{n-\ell}{3}p^{3}$ triangles with probability at least 11/121-1/12. Since $S$ is a clique, it has $\binom{\ell}{3}$ triangles. Since $\ell\geq 2np$, the number of triangles in this graph is at least $2\binom{n}{3}p^{3}$ with high probability. ∎
Proof of Theorem A.7.
Let $A$ be an algorithm that makes $q$ queries, and accepts $G\sim\mathcal{G}_{n,p}$ with proability $\geq 2/3$ and rejects $G$ with $\rho_{H}(G)>1.5$ with probability at most $1/3$. By Lemma A.9, $\Pr_{G^{\prime}\sim D_{N}^{n,\ell}}[\rho_{\Delta}(G)\leq 1.9]\leq 1/12$, and therefore $\Pr_{G\sim\mathcal{G}_{n,p},R}[A(G)=1]-\Pr_{G^{\prime}\sim D_{N}^{n,\ell},R}[A(G^{\prime})=1]\geq 1/6$, for $R$ denoting the random coin tosses made by algorithm $A$. Thus, there exists $R_{0}$ satisfying $\mathbb{E}_{G\sim\mathcal{G}_{n,p}}[A^{\prime}(G)=1]-\mathbb{E}_{G^{\prime}\sim D_{N}^{n,\ell}}[A^{\prime}(G^{\prime})=1]\geq 1/6$, for algorithm $A^{\prime}$ that is the deterministic version of $A$ with the random tape set to be $R_{0}$. Combining this with Lemma A.8 yields $q\geq(n-\ell)/(12\ell)$. ∎
A.3 The challenge of extending to k-cliques
Suppose we wanted to extend the approach above to quality control of $k$-cliques as compared to $\mathcal{G}_{n,p}$. When we apply [EMR22] (arboricity approximation) and [ERS20a] ($k$-clique count approximation in low arboricity graphs), we obtain a query complexity and runtime scaling with $1/p^{O(k^{2})}\cdot\text{poly}(\log(n),1/\varepsilon)$. Considering low arboricity graphs eliminates some lower-order $O(k)$ terms in the exponent, but the ratio between the arboricity and the number of $k$-cliques in $\mathcal{G}_{n,p}$ is not enough to bring the query complexity and runtime to $1/p^{O(k)}$. Indeed, additionally assuming access to uniform edge queries and using corresponding algorithms [AKK19] does not get rid of the $O(k^{2})$ term in the exponent. The issue, at an intuitive level, arises because the number of $k$-cliques in $\mathcal{G}_{n,p}$, for sparse $p$, is so few, so bounds that scale inversely with the number of $k$-cliques seem to encounter a $1/p^{O(k^{2})}$-style bound. It is an interesting future direction to adapt and simplify algorithms such as those in [ERS20a, AKK19] for quality control of $k$-clique counts as compared to $\mathcal{G}_{n,p}$. In the next section, we instead take a different approach, iteratively understanding the number of $\ell$-cliques, for $\ell\leq k$, in the graph and developing a notion of “robust quasirandomness” of clique counts on subgraphs of the input graph.
Appendix B Quality control for Sum-NC0 graph parameters
In this section, we provide an alternative perspective on quality control of motif counts and its implications for the structure of an input graph. As we have noted, quality control on motif counts is a fundamental question for several reasons, including that motif counts can provide evidence that a graph looks quasirandom (e.g. [CGW89, SY10]). Quality control for motif counts can further be seen to be a fundamental problem because quality control algorithms for motif counts can be used to construct quality control algorithms for more general parameters of graphs, which we show in this section.
We consider graph parameters (as defined in Section 2.1) computable by $\sum-NC^{0}$ circuits, defined formally below.
Definition B.0.
A Sum-NC0 ($\sum-NC^{0}$) circuit $C=\sum\circ~C_{n,m}$ consists of a summation gate $\sum_{m}:\{0,1\}^{m}\to\mathbb{N}$ composed with a circuit $C_{n,m}:\{0,1\}^{n}\to\{0,1\}^{m}$ with constant fan-in $k$, constant depth $d$, and $m=\mathrm{poly}(n)$, consisting of AND, OR, and NOT gates.
Recall the definition of a graph parameter from Section 2.1. For a graph parameter $F$, the corresponding quality control algorithm must, with high probability, accept $G\sim\mathcal{G}_{n,p}$ and reject graphs $G^{\prime}$ for which $F(G^{\prime})\not\in(1\pm\varepsilon)\cdot\mathbb{E}_{G\sim\mathcal{G}_{n,p}}\left(F(G)\right)$.
For a graph parameter $F:\mathcal{G}(n)\to\mathbb{N}$, let its circuit representation have input from $\{0,1\}^{\binom{n}{2}}$ representing which edges exist in the graph, and an output of the function value of the graph. Let $\rho_{F}(G):=F(G)/\mathbb{E}_{G^{\prime}\sim\mathcal{G}_{n,p}}\left[F(G^{\prime})\right]$.
Theorem B.1.
Consider any graph parameter $F:\mathcal{G}(n)\to\mathbb{N}$ whose circuit representation can be expressed as a $\sum-NC^{0}$ circuit with gate fan-in at most $k$ and depth $d$, such that $\mathbb{E}_{G^{\prime}\sim\mathcal{G}_{n,p}}\left[F(x^{G^{\prime}})\right]>0$. Let $p=\Omega\left(1/n^{2/(2k^{d}-1)}\right)$. The $(\mathcal{G}_{n,p},\rho_{F})$-quality control problem uses $1/p^{O(k^{d})}$ queries to the adjacency matrix of the input graph, and runtime $1/p^{O(k^{d})}$.
Graph parameters with Sum-NC0 circuits are a generalization and alternative formulation of motif counts. These graph parameters count the number of appearances of local patterns in a graph.
We will, without loss of generality, assume and use that $\mathbb{E}_{G^{\prime}\sim\mathcal{G}_{n,p}}\left[F(x^{G^{\prime}})\right]\geq 1$. If this is not the case, scaling $F$ by some factor to obtain a graph parameter $F^{\prime}$ is possible, and will allow us to study and perform quality control on $F$. $\rho_{F}$ is invariant under scaling.
The main technical work for proving Theorem B.1 arises in reducing a graph parameter $F$ whose circuit representation can be expressed as a Sum-NC0 circuit with gate fan-in at most $k$ and depth $d$ to a sum of counts of motifs of size at most $k^{d}$. Once we have done this, we can use the composition of quality control problems. We will perform $(\mathcal{G}_{n,p},\rho_{H})$-quality control for all motifs $H$ on at most $k^{d}$ vertices. If any of the $(\mathcal{G}_{n,p},\rho_{H})$-quality control algorithms called rejects, then so does the algorithm we construct for $(\mathcal{G}_{n,p},\rho_{F})$-quality control, where $\rho_{F}(G)=F(G)/\mathbb{E}_{G^{\prime}\sim\mathcal{G}_{n,p}}\left[F(G^{\prime})\right]$ for graph parameter $F$. We argue that, due to the representation of $F$ as a sum of counts of small motifs, this suffices for $(\mathcal{G}_{n,p},\rho_{F})$-quality control.
Since quality control for any such Sum-NC0 graph parameter can be accomplished via quality control of all small motif counts, this also demonstrates the robustness of the composability property of quality control problems.
B.1 Reducing graph parameters to sums of motif counts
In this section we prove the following lemma.
Lemma B.2.
Consider any graph parameter $F:\mathcal{G}(n)\to\mathbb{N}$ whose circuit representation can be expressed as a $\sum-NC^{0}$ circuit $C=\sum\circ~C_{\binom{n}{2},m}$ where $C_{\binom{n}{2},m}$ has gate fan-in $k$ and depth $d$. Then, $F$ can be expressed as a (scaled) sum over labeled induced counts of motifs on at most $2k^{d}$ vertices in the input graph.
Proof.
First, any $NC^{0}$ circuit $C_{\binom{n}{2},m}:\{0,1\}^{\binom{n}{2}}\to\{0,1\}^{m}$ with gate fan-in $k$ and depth $d$ can be written as a summation over at most $m\cdot 2^{k^{d}}$ AND gates with fan-in at most $k^{d}$. To see this, observe that such a circuit can only depend on at most $k^{d}$ input variables/coordinates. Therefore, the circuit can be written as a summation over at most $2^{k^{d}}$ AND gates with fan-in $k^{d}$, taken over input coordinates and/or their negations. To see this, represent each bit of the output of the $NC^{0}$ circuit as a truth table. Since the $NC^{0}$ circuit has fan-in $k^{d}$, and coordinates may be negated, the truth table corresponds to an OR taken over at most $2^{k^{d}}$ AND gates with fan-in $k^{d}$. This OR can be turned into a summation over AND gates, because this originated from a truth table in which only one AND gate will output 1 on an input.
For $i\in[m\cdot M(k,d)]$, and $S_{i}\subseteq[\binom{n}{2}$], $|S_{i}|\leq k^{d}$, let $g_{i}:S_{i}\to\{0,1\}$ be the $i$-th AND gate in this new representation of $\Sigma-NC^{0}$ circuit $C$. Let $N_{i}\subseteq S_{i}$ consist of the bits of the input that should be negated before taking the AND gate. Let $h_{i}:S_{i}\to\{0,1\}$ be defined as $h_{i}(x_{j})=1-x_{j}$ if $x_{j}\in N_{i}$ and $h_{i}(x_{j})=x_{j}$ if $x_{j}\not\in N_{i}$.
We can express any $g_{i}$ for $i\in[m\cdot M(k,d)]$ as:
$g_{i}(x)=\prod_{j\in S_{i}}h_{i}(x_{j}).$ (21)
Therefore, equating a graph $G$ with the string $x^{G}\in\{0,1\}^{\binom{n}{2}}$ indicating the edges present, we can rewrite $C:\{0,1\}^{n}\to\{0,1\}^{m}$ equivalently as
$C(x^{G})=\sum_{i\in[2^{k^{d}}]}g_{i}(x^{G})=\sum_{i\in[2^{k^{d}}]}\prod_{j\in S_{i}}h_{i}(x^{G}_{j}).$
Consider permutations $\sigma(x^{G})$ of the vertices of the $n$-vertex input graph $G$. First, since $F$ is invariant under graph isomorphism by definition of a graph parameter:
$F(x^{G})=\frac{1}{n!}\sum_{\sigma}F(\sigma(x^{G}))=\frac{1}{n!}\sum_{\sigma}\sum_{i\in[m\cdot M(k,d)]}g_{i}(\sigma(x^{G}))=\frac{1}{n!}\sum_{i\in[m\cdot M(k,d)]}\sum_{\sigma}g_{i}(\sigma(x^{G})).$
We now argue that $\sum_{\sigma}g_{i}(\sigma(x^{G}))$ corresponds to counting some induced motifs on at most $2k^{d}$ vertices in the graph. First, the set $\{\sigma(x^{G})\}_{\sigma}$ includes all ordered sets of vertices in $G$. Therefore, using the formula Equation 21 for $g_{i}$, $\sum_{\sigma}g_{i}(\sigma(x^{G}))$ counts the number of appearances of a specific pattern $P$ of edges and non-edges among all labeled subsets of $|S_{i}|$ edges in the graph. This count can be computed by summing over all induced motifs on up to $2|S_{i}|$ vertices whose edge/non-edge pattern includes $P$. Therefore, $\sum_{\sigma}g_{i}(\sigma(x^{G}))$ is equivalent to a sum over a number of labeled induced counts of some motifs on at most $2|S_{i}|\leq 2k^{d}$ vertices in the input graph. ∎
B.2 Quality control algorithm and analysis
We now present the quality control algorithm for graph parameters computable by Sum-NC0 circuits, as we have developed the technical lemmas needed to prove its correctness as a quality control algorithm.
Algorithm Graph-Parameter-Quality:
On inputs $G$ and $n,p,\varepsilon$
  1. 1.
    For all motifs $H$ on at most $2k^{d}$ vertices, run Algorithm Induced-Motif-Quality-Efficient (on inputs $G,n,p,\mathcal{G}_{n,p},\varepsilon$, and $H$) to perform quality control on the count of labeled induced copies of $H$ in $G$.
  2. 2.
    If the quality control algorithm rejects on any motif $H$ on at most $2k^{d}$ vertices, then output reject. Else, output accept.
Observe that the quality control algorithm is the same for all graph parameter computable by a Sum-NC0 circuit with fan-in $k$.
Towards proving Theorem B.1, we begin with the following lemma.
Lemma B.3.
Let $F$ be a graph parameter $F:\mathcal{G}(n)\to\mathbb{N}$ whose circuit representation is a Sum-NC0 circuit $C=\sum\circ~C_{\binom{n}{2},m}$, which satisfies $\mathbb{E}_{G^{\prime}\sim\mathcal{G}_{n,p}}\left[F(x^{G^{\prime}})\right]\geq 1$. Consider a graph $G\in\mathcal{G}(n)$ and the corresponding string with edges/non-edges $x^{G}\in\{0,1\}^{\binom{n}{2}}$. If
$\left|F(x^{G})-\mathbb{E}_{G^{\prime}\sim\mathcal{G}_{n,p}}\left[F(x^{G^{\prime}})\right]\right|\geq\varepsilon\cdot\mathbb{E}_{G^{\prime}\sim\mathcal{G}_{n,p}}\left[F(x^{G^{\prime}})\right],$
then there exists a motif $H$ on at most $2k^{d}$ vertices whose labeled induced count $C_{H}(G)$ satisfies:
$\left|C_{H}(G)-\mathbb{E}_{G^{\prime}\sim\mathcal{G}_{n,p}}\left[C_{H}(G^{\prime})\right]\right|\geq\varepsilon\cdot\mathbb{E}_{G^{\prime}\sim\mathcal{G}_{n,p}}\left[C_{H}(G^{\prime})\right].$ (22)
Proof.
Let $\mu^{(F)}:=\mathbb{E}_{G^{\prime}\sim\mathcal{G}_{n,p}}\left[F(x^{G^{\prime}})\right]$. Let $\mu_{H}:=\mathbb{E}_{G^{\prime}\sim\mathcal{G}_{n,p}}\left[C_{H}(G^{\prime})\right]$. Let $M(k,d)=\sum_{j=1}^{2k^{d}}2^{\binom{j}{2}}$, which is the number of induced subgraphs on at most $2k^{d}$ vertices.
By the triangle inequality, we see:
$\varepsilon\cdot\mu^{(F)}\leq\left|F(x^{G})-\mu^{(F)}\right|\leq\frac{1}{n!}\sum_{i\in[m\cdot M(k,d)]}\left|\sum_{\sigma}g_{i}(\sigma(x^{G}))-\mathbb{E}_{G^{\prime}\sim\mathcal{G}_{n,p}}\left[\sum_{\sigma}g_{i}(\sigma(x^{G^{\prime}}))\right]\right|.$
Therefore, there exists an $i\in[m\cdot M(k,d)]$ such that $\left|\sum_{\sigma}g_{i}(\sigma(x^{G}))-\mathbb{E}_{G^{\prime}\sim\mathcal{G}_{n,p}}\left[\sum_{\sigma}g_{i}(\sigma(x^{G^{\prime}}))\right]\right|$ is at least $\varepsilon n!/(m\cdot M(k,d))$. Since $\sum_{\sigma}g_{i}(\sigma(x^{G}))$ corresponds to counting some induced motif on at most $k^{d}$ vertices in the graph (as reasoned about in the proof of Lemma B.2), we find that:
$\left|C_{H}(G)-\mathbb{E}_{G^{\prime}\sim\mathcal{G}_{n,p}}\left[C_{H}(G^{\prime})\right]\right|\geq\frac{\varepsilon\cdot\mathbb{E}_{G^{\prime}\sim\mathcal{G}_{n,p}}\left[F(x^{G^{\prime}})\right]\cdot n!}{m\cdot M(k,d)}\geq\varepsilon\cdot\mu_{H},$
where the last inequality holds because $m=\mathrm{poly}(n)$, $\mathbb{E}_{G^{\prime}\sim\mathcal{G}_{n,p}}\left[F(x^{G^{\prime}})\right]\geq 1$ and $\mu_{H}\leq\binom{n}{2k^{d}}$. Thus Equation 22 holds. ∎
We now prove Theorem B.1 for Algorithm Graph-Parameter-Quality.
Proof of Theorem B.1.
Completeness: Suppose the input graph $G$ is drawn from $\mathcal{G}_{n,p}$. Then, by the completeness of Algorithm Induced-Motif-Quality-Efficient proven in Section 4.1, each call of Algorithm Induced-Motif-Quality-Efficient in Step 1 accepts $G$ with probability at least $1-\frac{1}{3\cdot M(kd)}$. (Observe that Algorithms 2 and 3 were analyzed with a different constant success probability, but since $M(k,d)$ is constant with respect to $n,p$, we can replace it with $1-\frac{1}{3\cdot M(k,d)}$ in the analysis.) By a union bound over the $M(k,d)$ possible induced subgraphs over at most $2k^{d}$ vertices, Step 2 then accepts $G\sim\mathcal{G}_{n,p}$ with probability at least $2/3$.
Soundness: Suppose the input $G$ satisfies
$\left|F(x^{G})-\mathbb{E}_{G^{\prime}\sim\mathcal{G}_{n,p}}\left[F(x^{G^{\prime}})\right]\right|\geq\varepsilon\cdot\mathbb{E}_{G^{\prime}\sim\mathcal{G}_{n,p}}\left[F(x^{G^{\prime}})\right].$
By Lemma B.3, there exists a motif $H$ on at most $k^{d}$ vertices whose labeled induced count $C_{H}(G)$ satisfies:
$\left|C_{H}(G)-\mathbb{E}_{G^{\prime}\sim\mathcal{G}_{n,p}}\left[C_{H}(G^{\prime})\right]\right|\geq\varepsilon\cdot\mathbb{E}_{G^{\prime}\sim\mathcal{G}_{n,p}}\left[C_{H}(G^{\prime})\right].$
By the soundness of Algorithm Induced-Motif-Quality-Efficient proven in Section 4.1, therefore the algorithm rejects input graph $G$ due to the count of motif $H$ with high probability.
Query complexity and runtime: These are inherited from the calls of Algorithm Induced-Motif-Quality-Efficient on motifs on at most $k^{d}$ vertices. Therefore, the query complexity and runtime are $1/p^{O(k^{d})}$. ∎
Appendix C Connection to average-case complexity
C.1 From average-case complexity to quality control and back
Quality control has a close connection to the average-case complexity of algorithmic / computational problems, which we further explore in this section. Average-case complexity typically asks for the following: Given a distribution $\mathcal{D}$ over inputs, we want to solve the problem on every input, such that the expected runtime taken over $\mathcal{D}$ is low. We discuss how to view average-case problems through the perspective of quality control, and vice versa.
Average-case problems to quality control problems.
Works on average-case algorithms typically define a “good” property of inputs with the following guarantees. With probability at least $1-1/t_{2}$, an input drawn from $\mathcal{D}$ possesses the good property, and thus there exists an algorithm $A$ with worst-case runtime $t_{1}$ for this input. This “good” property is possible to certify/verify. On the other hand, for worst-case inputs, the algorithm $A$ has worst-case runtime $t_{2}$. In such a set-up, the average-case complexity of the algorithm $A$ is:
$(1-1/t_{2})\cdot t_{1}+(1/t_{2})\cdot t_{2}=O(t_{1}).$
See [DF89, MPR+25] for examples of papers that define and work with such a good property.
When an average-case problem has such a “good” property, we can define a quality control problem as follows. We consider $(\mathcal{D},\rho)$-quality control, with the same distribution $\mathcal{D}$ over inputs, and $\rho$ defined as:
$\rho(x)=\begin{cases}1~~\text{ if $A(x)\leq t_{1}$}\\ 0~~\text{ otherwise.}\end{cases}$
Consider the algorithm $B$ that, on input $x$, accepts if $A$ halts within $\leq t_{1}$ steps on input $x$. If algorithm $A$ has average-case complexity $t_{1}$, and $1/t_{2}\leq 1/3$, then algorithm $B$ is an algorithm for $(\mathcal{D},\rho)$-quality control. For completeness, $x\sim\mathcal{D}$ is good with high probability and thus $\rho(x)=1$. For soundness, if the runtime is longer than $O(t_{1})$, this is detectable and $\rho(x)=0$.
Quality control problems to average-case problems
Quality control problems serve as an intermediate between the problems of “Constructing an algorithm that is correct with high probability over $\mathcal{D}$” and average-case complexity, i.e., “Constructing an algorithm that is correct on every input, and has average-case time $T$ over $\mathcal{D}$.”
$(\mathcal{D},\rho)$-quality control serves as an intermediate because, first, it checks which inputs the algorithm is correct with (by soundness) as opposed to the former, but does not give worst-case algorithmic guarantees for computing $\rho(x)$ when $\rho(x)\not\approx\mathbb{E}_{y\sim\mathcal{D}}[\rho(y)]$, which average-case algorithms must.
In general, we have the following.
Proposition C.0.
Let $\mathcal{D}$ be a distribution supported over $\mathcal{X}$, and let $\rho:\mathcal{X}\to\mathbb{R}_{\geq 0}$ be a parameter satisfying
$\mathbb{P}_{x\sim\mathcal{D}}\left[\left|\rho(x)-\mathbb{E}_{y\sim\mathcal{D}}[\rho(y)]\right|\geq\varepsilon\cdot\mathbb{E}_{y\sim\mathcal{D}}[\rho(y)]\right]\leq 1/T,$
for some $T\geq 0$.
Suppose $C$ is a $(\mathcal{D},\rho)$-quality control algorithm with query complexity and runtime $t_{3}$. Suppose $E$ is an algorithm that computes a $(1\pm\varepsilon)$-approximation of $\rho(x)$ for any $x\in\mathcal{X}$ and has query complexity and runtime $t_{4}\leq T/2$. Then there exists an algorithm that approximates $\rho(x)$ to a $(1\pm\varepsilon)$ multiplicative factor with high constant probability (over the algorithm’s randomness) on any input $x\in\mathcal{X}$ and has average-case query complexity and runtime $O(t_{3})$ over $\mathcal{D}$.
The algorithm for approximating $\rho(x)$ is given as follows. As in Section C.1, suppose $C$ is a $(\mathcal{D},\rho)$-quality control algorithm with query complexity and runtime $t_{3}$. Suppose $E$ is an algorithm that computes a $(1\pm\varepsilon)$-approximation for any $x\in\mathcal{X}$ and has query complexity and runtime $t_{4}$.
Average-case-from-quality-control:
On input $x$:
  1. 1.
    If $C(x)=$ accept, then output $\mathbb{E}_{y\sim\mathcal{D}}[\rho(y)]$.
  2. 2.
    Else, output $E(x)$.
Proof of Section C.1.
First, by the correctness of algorithms $C$ and $E$ for quality control and $\rho$-approximation, respectively, the algorithm always outputs a $(1\pm\varepsilon)$-approximation of $\rho(x)$.
We now argue about the runtime. First, with probability at least 11/T1-1/T, $x\sim\mathcal{D}$ satisfies $\rho(x)\in(1\pm\varepsilon)$, and so with probability at least $(1-1/T)\cdot 2/3\geq 1-1/t_{4}$ (over the input distribution), $x\sim\mathcal{D}$ is accepted by Algorithm $C$. In this case, the algorithm takes time $t_{3}$. If $x$ is not accepted by Algorithm $C$, then Algorithm $E$ is called, and will take time $t_{4}$. Therefore, the average-case runtime of the algorithm is:
$t_{3}\cdot(1-1/t_{4})+(1/t_{4})\cdot t_{4}=O(t_{3}).$
C.2 Approximate motif counting
We now instantiate the framework of moving from quality control problems to average-case problems for the case of approximately counting the number of copies of a motif in a graph. Such an algorithm must be correct with high constant probability (over the algorithm’s randomness) on all inputs.
To put the various types of average-case questions, including quality control, in comparison for motif counting, recall that $G\sim\mathcal{G}_{n,p}$ satisfies the property that the number of copies of a motif $H$ is approximately equal to its expected amount under $\mathcal{G}_{n,p}$, with high probability. However, outputting this expectation for all inputs $G$ is not meaningful. Quality control gives us a way of efficiently verifying, with high probability, whether the number of copies of $H$ is indeed close to its expectation under $\mathcal{G}_{n,p}$, telling us whether this is indeed a good estimate for the count of the motif $H$ in the specific graph $G$. Suppose we want an algorithm that outputs a good estimate for all $G$, with good average-case query complexity and runtime over $\mathcal{G}_{n,p}$. As described more generally in Section C.1, we can do so by combining quality control and a worst-case algorithm that is called very infrequently.
Specifically, from Section C.1, we obtain the following.
Theorem C.1.
Let $H$ be a motif on $k$ vertices. Consider parameters $n,p$ such that $n\geq p^{-c\cdot m_{H}}$ for some constant $c$. The average-case complexity over $\mathcal{G}_{n,p}$ of approximately counting motifs in a graph with high probability is $1/p^{O(\Delta(H))}$.
More generally, our results imply an average-case complexity guarantee when any distribution over graphs is taken that satisfies the concentration guarantees of Theorem 5.1.
Let $C_{H}(G)$ be the number of labeled induced copies of motif $H$ in $G$. We utilize the following lemma, whose proof follows almost exactly as that of Section 2.3.
Lemma C.2.
Let $H$ be a motif on $k$ vertices. There exists a $c$ such that for parameters $n,p$ such that $n/\log(n)\geq p^{-c\cdot m_{H}}$, $G\sim\mathcal{G}_{n,p}$ satisfies $C_{H}(G)\in(1\pm\varepsilon/100)\binom{n}{k}p^{e(H)}(1-p)^{\binom{k}{2}-e(H)}$ with probability at least $1-1/n^{O(k)}$.
For our worst-case algorithm (which returns an approximation for the number of $H$ copies for any graph), we use the trivial algorithm that enumerates over all $|H|$-size subgraphs of the input graph and checks if this subgraph is a copy of $H$. (There may be more efficient algorithms, but we want something that works independent of the number of $H$ copies in the graph or other properties, and this suffices.)
Let $\rho_{H}(G)=C_{H}(G)/\mathbb{E}_{G^{\prime}\sim\mathcal{G}_{n,p}}\left[C_{H}(G^{\prime})\right]$, for $C_{H}(G)$.
Proof of Theorem C.1.
First, Algorithm Induced-Motif-Quality-Efficient from Section 4.3.2 is a $(\mathcal{G}_{n,p},\rho_{H})$-quality control algorithm with query complexity and runtime $1/p^{O(\Delta(H))}$. Combining this with Lemma C.2 and the worst-case algorithm above using $n^{O(k)}$ queries and runtime in the worst case over graphs, by applying Section C.1, we find that the average-case complexity of approximate counting is $1/p^{O(\Delta(H))}$. ∎
Appendix D Concentration of subgraph counts
For the soundness arguments of the paper – i.e., the statement that a $(\mathcal{G}_{n,p},\rho)$-quality control algorithm accepts $G\sim\mathcal{G}_{n,p}$ with high probability – we will utilize that, in $\mathcal{G}_{n,p}$, the counts of small motifs are well-concentrated around their expectations. Additionally, for lower-bounding the query complexity of quality control for motif counts, we will consider an alternative random graph distribution known as the stochastic block model and need concentration of counts of small motifs. In this section, we provide proofs of concentration of small motif counts for these two models, relying only on the second-moment method.
D.1 Erdős–Rényi graphs
In this section, we prove Section 2.3 using the second moment method. For convenience, we re-state Section 2.3.
Proposition D.0 (Section 2.3, restated).
Let $H$ be an induced motif on $k$ vertices. Consider parameters $n,p$ with $n\geq\frac{10}{\varepsilon^{2}}\cdot p^{-m_{H}}$, where $m_{H}:=\max_{F\subseteq H}e(F)/v(F)$.
For $G\sim\mathcal{G}_{n,p}$, with probability at least $9/10$, the number of copies of $H$ in $G$ is in the range $(1\pm\frac{\varepsilon}{100})\binom{n}{v(H)}v(H)!\cdot p^{e(H)}(1-p)^{\binom{v(H)}{2}-e(H)}$.
Proof.
Let $X_{G}$ be the number of labeled induced copies of $H$ in a graph $G$. First, observe that $\mathbb{E}\left[X_{G}\right]=\binom{n}{v(H)}v(H)!\cdot p^{e(H)}(1-p)^{\binom{v(H)}{2}-e(H)}$.
Let’s now analyze the variance of $X_{G}$. For an ordered set $S$, let $S\equiv H$ mean that $S$ forms a labeled induced copy of $H$. We can write $X_{G}$ as:
$X_{G}=\sum_{S\in[n]^{k}}\mathbbm{1}(S\equiv H).$
Therefore, the variance is:
$\mathrm{Var}\left[X_{G}\right]=\mathbb{E}\left[X_{G}^{2}\right]-\mathbb{E}\left[X_{G}\right]^{2}$
$\sum_{S\in[n]^{k}}\sum_{T\in[n]^{k}}\mathbb{E}\left[\mathbbm{1}(S\equiv H)\mathbbm{1}(T\equiv H)\right]-\mathbb{E}\left[X_{G}\right]^{2}$
$\mathbb{E}\left[X_{G}\right]^{2}+\sum_{i=1}^{k-1}\sum_{\begin{subarray}{c}S,T\in[n]^{k}\\ |S\cap T|=i\end{subarray}}\mathbb{E}\left[\mathbbm{1}(S\equiv H)\mathbbm{1}(T\equiv H)\right]+\mathbb{E}\left[X_{G}\right]-\mathbb{E}\left[X_{G}\right]^{2}$
$\mathbb{E}\left[X_{G}\right]+\sum_{i=1}^{k-1}\sum_{\begin{subarray}{c}H^{\prime}\subseteq H\\ |H^{\prime}|=i\end{subarray}}\sum_{\begin{subarray}{c}S,T\in[n]^{k}\\ |S\cap T|=i\end{subarray}}\mathbb{P}((S\cap T)\equiv H^{\prime},S\equiv H,T\equiv H)$
$\mathbb{E}\left[X_{G}\right]+\sum_{i=1}^{k-1}\sum_{\begin{subarray}{c}H^{\prime}\subseteq H\\ |H^{\prime}|=i\end{subarray}}\binom{n}{k}k!\cdot\binom{n-k}{k-i}(k-i)!\cdot k\cdot p^{2e(H)-e(H^{\prime})}(1-p)^{2\binom{k}{2}-2e(H)-\binom{i}{2}+e(H^{\prime})}$
$\leq\mathbb{E}\left[X_{G}\right]+\sum_{i=1}^{k-1}\sum_{\begin{subarray}{c}H^{\prime}\subseteq H\\ |H^{\prime}|=i\end{subarray}}\binom{n}{k}^{2}(k!)^{2}p^{2e(H)}\cdot\binom{n}{i}^{-1}\cdot p^{-e(H^{\prime})}$
$\mathbb{E}\left[X_{G}\right]+\mathbb{E}\left[X_{G}\right]^{2}\cdot\sum_{i=1}^{k-1}\sum_{\begin{subarray}{c}H^{\prime}\subseteq H\\ |H^{\prime}|=i\end{subarray}}\binom{n}{i}^{-1}\cdot p^{-e(H^{\prime})}.$
Since $n\geq\frac{10}{\varepsilon^{2}}\cdot p^{-m_{H}}$, this gives us:
$\text{Var}\left[X_{G}\right]\leq\mathbb{E}\left[X_{G}\right]^{2}\cdot\frac{\varepsilon^{2}}{10}.$
Finally, by Chebyshev’s inequality,
$\mathbb{P}\left[\left|X_{G}-\mathbb{E}\left[X_{G}\right]\right|\geq\varepsilon\mathbb{E}\left[X_{G}\right]\right]\leq\frac{\text{Var}\left[X_{G}\right]}{\varepsilon^{2}\mathbb{E}\left[X_{G}\right]^{2}}\leq\frac{1}{10}.$
D.2 Stochastic Block Models
We now prove that motifs are concentrated in stochastic block models with two communities by using the second moment method.
Consider a two-community stochastic block model. Community 1 has $n_{1}$ vertices (with $n_{1}\geq n/2$), and in-community edge probability $p$. Community 2 has $n_{2}$ vertices, and in-community edge probability $1/2$. The between-community edge probability is $1/2$. Let $\mathcal{G}_{\mathrm{SBM},n,p}$ be the distribution over graphs given by this model. Without loss of generality, suppose Community 1 is on the first $n_{1}$ vertices of the graph, according to some arbitrary ordering.
Proposition D.0.
Consider parameters $n,n_{1},n_{2},p$ satisfying $n_{1}+n_{2}=n$, $n\geq 20(2k)^{2k+1}p^{-\Delta(H)}/\varepsilon^{2}$, and $2^{2\binom{k}{2}+1}np^{\Delta(H)/2}\leq n_{2}\leq n/2$, where $\Delta(H)$ is the maximum vertex degree of $H$.
Consider any motif $H$ on $k$ vertices. With probability at least $\frac{9}{10}$, the number of labeled induced copies of $H$ in $G\sim\mathcal{G}_{\mathrm{SBM},n,p}$ is within a $(1\pm\varepsilon)$ multiplicative factor of its expectation.
We prove this by proving the following stronger statement.
Lemma D.1.
Consider parameters $n,n_{1},n_{2},p$ satisfying $n_{1}+n_{2}=n$, $n\geq 20(2k)^{2k+1}p^{-\Delta(H)}/\varepsilon^{2}$, and $2^{2\binom{k}{2}+1}np^{\Delta(H)/2}\leq n_{2}\leq n/2$, where $\Delta(H)$ is the maximum vertex degree of $H$.
Consider any motif $H$ on $k$ vertices, and any $H^{\prime}\subseteq H$ on $k^{\prime}$ vertices. For a graph $G$ with $n$ vertices, let $N_{H^{\prime},H}(G)$ be the number of copies of $H$ in $G$, such that $H^{\prime}$ is placed within the first $n_{1}$ vertices of $G$, and $H\setminus H^{\prime}$ is placed within the other $n_{2}$ vertices.
With probability at least $1-\frac{1}{10}\cdot\frac{1}{2^{k}}$ over $G\sim\mathcal{G}_{\mathrm{SBM},n,p}$,
$N_{H^{\prime},H}(G)\in(1\pm\varepsilon)\binom{n_{1}}{k^{\prime}}\binom{n_{2}}{k-k^{\prime}}k!\cdot p^{e(H^{\prime})}(1-p)^{\binom{k^{\prime}}{2}-e(H^{\prime})}(1/2)^{\binom{k}{2}-\binom{k^{\prime}}{2}}$. (23)
Proof.
We apply Chebyshev’s inequality; therefore, we must analyze the expectation and variance of $N_{H^{\prime},H}(G)$. For ease of notation, we refer to $N_{H^{\prime},H}(G)$ as $X_{G}$
First, we see
$\mathbb{E}_{G\sim\mathcal{G}_{\mathrm{SBM},n,p}}\left[X_{G}\right]=\binom{n_{1}}{k^{\prime}}\binom{n_{2}}{k-k^{\prime}}k!\cdot p^{e(H^{\prime})}(1-p)^{\binom{k^{\prime}}{2}-e(H^{\prime})}(1/2)^{\binom{k}{2}-\binom{k^{\prime}}{2}}$.
Next, we analyze the variance. For an ordered set $S$, let $S\equiv H$ mean that $S$ forms a labeled induced copy of $H$. First, observe that we can write $X_{G}$ as:
$X_{G}=\sum_{\begin{subarray}{c}S_{1}\in[n_{1}]^{k^{\prime}}\\ S_{2}\in[n_{2}]^{k-k^{\prime}}\end{subarray}}\mathbbm{1}(S_{1}\equiv H^{\prime})\mathbbm{1}(S_{1}\cup S_{2}\equiv H).$
Therefore,
$\mathrm{Var}\left[X_{G}\right]=\mathbb{E}\left[X_{G}^{2}\right]-\mathbb{E}\left[X_{G}\right]^{2}$
$\sum_{\begin{subarray}{c}S_{1}\in[n_{1}]^{k^{\prime}}\\ S_{2}\in[n_{2}]^{k-k^{\prime}}\end{subarray}}\sum_{\begin{subarray}{c}T_{1}\in[n_{1}]^{k^{\prime}}\\ T_{2}\in[n_{2}]^{k-k^{\prime}}\end{subarray}}\mathbb{E}\left[\mathbbm{1}(S_{1}\equiv H^{\prime})\mathbbm{1}(S_{1}\cup S_{2}\equiv H)\mathbbm{1}(T_{1}\equiv H^{\prime})\mathbbm{1}(T_{1}\cup T_{2}\equiv H)\right]-\mathbb{E}\left[X_{G}\right]^{2}.$
We break up the summation into various characterizations of $S_{1},S_{2},T_{1},T_{2}$ in terms of the various amounts of intersection between the sets.
Consider the summation over all $S_{1},S_{2},T_{1},T_{2}$ satisfying the following properties: $S_{1}$ and $T_{1}$ intersect on $i$ vertices, and $S_{2}$ and $T_{2}$ intersect on $j$ vertices.
First, suppose that $i\geq 1$. Let $F$ be the subgraph of $H^{\prime}$ with the fewest edges. Then, for $p\leq 1/2$, we find that the summation taken over the expectation of these indicators for these particular $S_{1},S_{2},T_{1},T_{2}$ is at most:
$\leq\binom{n_{1}}{2k^{\prime}-i}\binom{n_{2}}{2k-2k^{\prime}-j}(2k^{\prime}-i)!(2k-2k^{\prime}-j)!\cdot p^{2e(H^{\prime})-e(F)}(1-p)^{2\binom{k^{\prime}}{2}-2e(H^{\prime})+e(F)}(1/2)^{2\binom{k}{2}-2\binom{k^{\prime}}{2}-\binom{j}{2}+\binom{i}{2}}$. (24)
Since $\binom{n_{1}}{2k^{\prime}-i}\leq\binom{n_{1}}{k^{\prime}}^{2}/\binom{n_{1}}{i}$, $\binom{n_{2}}{2k-2k^{\prime}-j}\leq\binom{n_{2}}{2k-2k^{\prime}}/\binom{n_{2}}{j}$, and $(2k^{\prime}-i)!(2k-2k^{\prime}-j)!\leq(k!)^{2}\cdot(2k)^{2k}$, the above is:
$\leq\mathbb{E}\left[X_{G}\right]^{2}\cdot\binom{n_{1}}{i}^{-1}\binom{n_{2}}{j}^{-1}\cdot p^{-e(F)}(1-p)^{e(F)}\cdot(2k)^{2k}.$ (25)
First, when $i\geq 1$, $n_{1}\geq n/2\geq Cp^{-\Delta(H)}\geq Cp^{-e(F)/i}$, with some sufficiently large constant $C$; e.g., $C\geq 10(2k)^{2k+1}/\varepsilon^{2}$ suffices.
Therefore, $\binom{n_{1}}{i}^{-1}\binom{n_{2}}{j}^{-1}\cdot p^{-e(F)}\leq\frac{1}{10}\cdot\frac{1}{(2k)^{2k}}\cdot\frac{1}{k^{2}}\cdot\frac{1}{2^{k}}\cdot\varepsilon^{2}$. Therefore, Equation 25 is less than:
$\leq\mathbb{E}\left[X_{G}\right]^{2}\cdot\frac{1}{10}\cdot\frac{1}{k^{2}}\cdot\frac{\varepsilon^{2}}{2^{k}}.$
Next, consider upper-bounding Equation 25 when $i=0$, $j\geq 1$. Then Equation 24 is upper-bounded by:
$\leq\binom{n_{1}}{2k^{\prime}}\binom{n_{2}}{2k-2k^{\prime}-j}(2k^{\prime})!(2k-2k^{\prime}-j)!\cdot p^{2e(H^{\prime})}(1-p)^{2\binom{k^{\prime}}{2}-2e(H^{\prime})}(1/2)^{2\binom{k}{2}-2\binom{k^{\prime}}{2}-\binom{j}{2}}$.
$\leq\mathbb{E}\left[X_{G}\right]^{2}\cdot\binom{n_{2}}{j}^{-1}2^{\binom{j}{2}}\cdot(2k)^{2k}.$
Since $n_{2}\geq 2^{2\binom{k}{2}+1}np^{\Delta(H)/2}$ and $n\geq Cp^{-\Delta(H)}$, this is
$\leq\mathbb{E}\left[X_{G}\right]^{2}\cdot\frac{1}{10}\cdot\frac{1}{k^{2}}\cdot\frac{\varepsilon^{2}}{2^{k}}.$
Finally, when $i=j=0$, the summation over the expectation of the indicators for the corresponding $S_{1},S_{2},T_{1},T_{2}$ is $\leq\mathbb{E}\left[X_{G}\right]^{2}$ due to the independence of the $S_{1},S_{2}$ indicators with the $T_{1},T_{2}$ indicators.
Therefore, combining each of the summations (with different $i,j$ values), we find that
$\text{Var}\left[X_{G}\right]\leq\mathbb{E}\left[X_{G}\right]^{2}\cdot\frac{1}{10}\cdot\frac{\varepsilon^{2}}{2^{k}}.$
By Chebyshev’s inequality,
$\mathbb{P}\left[\left|X_{G}-\mathbb{E}\left[X_{G}\right]\right|\geq\varepsilon\mathbb{E}\left[X_{G}\right]\right]\leq\frac{\text{Var}\left[X_{G}\right]}{\varepsilon^{2}\mathbb{E}\left[X_{G}\right]^{2}}\leq\frac{1}{10}\cdot\frac{\varepsilon^{2}}{2^{k}}\cdot\frac{1}{\varepsilon^{2}}=\frac{1}{10\cdot 2^{k}}.$
Proof of Section D.2 from Lemma D.1.
Apply Lemma D.1 over every $H^{\prime}\subseteq H$. There are at most $2^{k}$ such $H^{\prime}$. By a union bound, the probability (over $G\sim\mathcal{G}_{\mathrm{SBM},n,p}$ that there exists an $N_{H^{\prime},H}(G)$ not in the range given in Equation 23 is at most $\frac{1}{10}$. Since the number of labeled induced copies of $H$ in $G$ is the sum over $H^{\prime}$ of the number of labeled induced copies of $H$ with $H^{\prime}$ in Community 1, we find that, with probability at least $\frac{9}{10}$, the number of labeled induced copies of $H$ in $G\sim\mathcal{G}_{\mathrm{SBM},n,p}$ is in a $(1\pm\varepsilon)$ multiplicative factor of its expectation. ∎

### v2

@@
 Definition 1.0 (Quality control algorithms).
-For function $\rho:\mathcal{X}\to\mathbb{R}_{\geq 0}$, small constant $\varepsilon>0$, and a distribution $\mathcal{D}:\mathcal{X}\to\mathbb{R}_{\geq 0}$ such that $\mathbb{P}_{x\sim\mathcal{D}}\left[\left|\rho(x)-1\right|\leq\varepsilon\right]\geq 9/10$, a $(\mathcal{D},\rho)$-quality control algorithm $A$ must satisfy the following:
+For function $\rho:\mathcal{X}\to\mathbb{R}_{\geq 0}$, small constant $\varepsilon>0$, and a distribution $\mathcal{D}:\mathcal{X}\to\mathbb{R}_{\geq 0}$ such that $\mathbb{P}_{x\sim\mathcal{D}}\left[\left|\rho(x)-1\right|\leq\varepsilon\right]\geq 1-o(1)$, a $(\mathcal{D},\rho)$-quality control algorithm $A$ must satisfy the following:
   1. 1.
-    $\mathbb{P}_{x\sim\mathcal{D};R}\left[A(x)=\textsc{Accept}\right]\geq 2/3$, where the probability is taken over distribution $\mathcal{D}$ and the randomness $R$ used by algorithm $A$.
+    $\mathbb{P}_{x\sim\mathcal{D};R}\left[A(x)=\textsc{Accept}\right]\geq 1-o(1)$, where the probability is taken over distribution $\mathcal{D}$ and the randomness $R$ used by algorithm $A$.
   2. 2.
     For all $x$ satisfying $\left|\rho(x)-1\right|>\varepsilon$, $\mathbb{P}_{R}\left[A(x)=\textsc{Reject}\right]\geq 2/3$.
+Above, we use $o(1)$ to mean that a quantity approaches 0 as the runtime of the algorithm tends to infinity.
+(Formally, the quality control algorithm takes a error parameter $\tau$ as input, and for every $\tau>0$ there exists a constant $c$ such that the quality control algorithm with error parameter $\tau$ accepts $x\sim\mathcal{D}$ with probability at least $1-\tau$ and the runtime of the quality control algorithm with parameter $\tau$ is at most $c$ times larger than the runtime with error parameter set to some fixed constant, say, $1/3$.)
 Let us first compare the related problems of testing if an input $x$ was drawn from $\mathcal{D}$ and approximating $\rho(x)$.
@@
 Let $H$ be an induced motif on $k$ vertices. Consider parameters $n,p$ with $p=\omega\left(n^{-1/m_{H}}\cdot\varepsilon^{-2}\right)$, where $m_{H}:=\max_{F\subseteq H}e(F)/v(F)$.
-For $G\sim\mathcal{G}_{n,p}$, with probability at least $9/10$, the number of labeled induced copies of $H$ in $G$ is in the range $(1\pm\frac{\varepsilon}{100})\binom{n}{v(H)}v(H)!\cdot p^{e(H)}(1-p)^{\binom{v(H)}{2}-e(H)}$.
+For $G\sim\mathcal{G}_{n,p}$, with probability $1-o_{1/p}(1)$, the number of labeled induced copies of $H$ in $G$ is in the range $(1\pm\frac{\varepsilon}{100})\binom{n}{v(H)}v(H)!\cdot p^{e(H)}(1-p)^{\binom{v(H)}{2}-e(H)}$.
 [Jan90, JOR04] have proven such a result previously – with more refined guarantees – for the case of non-induced subgraphs.
@@
 Proposition 2.0.
-For $G\sim\mathcal{G}_{n,p}$ and $p=\Omega\left(n^{-2/(k-1)}\cdot\varepsilon^{-2}\right)$, the number of $k$-cliques in $G$ is in the range $(1\pm\frac{\varepsilon}{100})\binom{n}{k}k!\cdot p^{\binom{k}{2}}$ with high constant probability (say, $9/10$).
+For $G\sim\mathcal{G}_{n,p}$ and $p=\omega\left(n^{-2/(k-1)}\cdot\varepsilon^{-2}\right)$, the number of $k$-cliques in $G$ is in the range $(1\pm\frac{\varepsilon}{100})\binom{n}{k}k!\cdot p^{\binom{k}{2}}$ with probability $1-o_{1/p}(1)$.
 Next, [GPS18] proves the following regarding the arboricity $\alpha(G)$ of $G\sim\mathcal{G}_{n,p}$ (which indeed is a precise statement that holds asymptotically almost surely, while we provide a weaker version of their statement below for simplicity).
 Proposition 2.0 ([GPS18]).
-For all $p=\Omega\left(\frac{1}{n}\right)$, there exists an $N$ such that for all $n\geq N$, $\mathbb{P}_{G\sim\mathcal{G}_{n,p}}\left(\alpha(G)\leq 2\frac{e(G)}{n-1}+1\right)\geq 9/10$.
+For all $p=\Omega\left(\frac{1}{n}\right)$, there exists an $N$ such that for all $n\geq N$, $\mathbb{P}_{G\sim\mathcal{G}_{n,p}}\left(\alpha(G)\leq 2\frac{e(G)}{n-1}+1\right)=1-o_{1/p}(1)$.
 Combining this with Section 2.3 (for the concentration of the number of edges in $\mathcal{G}_{n,p}$), we obtain the following corollary.
@@
 For $p=\Omega\left(\varepsilon^{-2}\cdot 1/n\right)$, there exists an $N$ such that for all $n\geq N$,
-$\mathbb{P}_{G\sim\mathcal{G}_{n,p}}\left(\alpha(G)\leq 2np\right)\geq 4/5.$
+$\mathbb{P}_{G\sim\mathcal{G}_{n,p}}\left(\alpha(G)\leq 2np\right)=1-o_{1/p}(1).$
 When we focus on the algorithmic complexity of quality control for $k$-cliques matching $\mathcal{G}_{n,p}$ in Section 3.3, we will utilize the notion of graph jumbledness (Section 2.2) to obtain an efficient runtime.
@@
 Let $\delta=\frac{p^{2k}}{12C^{2}}$.
-With probability at least $9/10$, for $G\sim\mathcal{G}_{n,p}$, at least a $9/10$-fraction of subgraphs $S$ of size $s$ satisfy the following:
+With probability $1-o_{1/p}(1)$, for $G\sim\mathcal{G}_{n,p}$, at least a $(1-o_{1/p}(1))$-fraction of subgraphs $S$ of size $s$ satisfy the following:
   1. 1.
@@
 We find that, with probability at most $1/s$, $G_{S}$ is not good.
-Therefore, the expected number of $S\subseteq[n],|S|=s$ such that $G_{S}$ is not good is at most $\frac{1}{s}\cdot\binom{n}{s}\leq\frac{1}{100}\binom{n}{s}$ for $s\geq 100$.
-For a graph $G$, let $t(G)$ be the number of $S\subseteq[n],|S|=s$ such that $G_{S}$ is not good. By Markov’s inequality,
-$\mathbb{P}\left(t(G)\geq\frac{1}{10}\cdot\binom{n}{s}\right)\leq\frac{\mathbb{E}(t(G))}{\frac{1}{10}\cdot\binom{n}{s}}\leq\frac{\frac{1}{100}\cdot\binom{n}{s}}{\frac{1}{10}\cdot\binom{n}{s}}\leq\frac{1}{10}.$
-Therefore, at least a $9/10$-fraction of subgraphs of size $s$ are good. ∎
+Therefore, the expected number of $S\subseteq[n],|S|=s$ such that $G_{S}$ is not good is at most $\frac{1}{s}\cdot\binom{n}{s}=o\left(\binom{n}{s}\right)$.
+Therefore, by Markov’s inequality, with probability at least $1-o_{1/p}(1)$, at least a $(1-o_{1/p}(1))$-fraction of subgraphs of size $s$ are good.
+∎
 We note the following corollary of this proposition, for ease of use.
@@
 Consider $G\sim\mathcal{G}_{n,p}$, and choose a random subgraph $G_{S}$ of $G$ on $s$ vertices.
-Then, with probability at least $4/5$, the subgraph on these $s$ vertices satisfies the following:
+Then, with probability $1-o_{1/p}(1)$, the subgraph on these $s$ vertices satisfies the following:
   1. 1.
@@
 Proof.
-This corollary follows by taking a union bound over the events in Section 2.3 that $G\sim\mathcal{G}_{n,p}$ does not satisfy the property that at least a $9/10$-fraction of size-$s$ subgraphs are good and that a non-good subgraph is chosen at random.
+This corollary follows by taking a union bound over the events in Section 2.3 that $G\sim\mathcal{G}_{n,p}$ does not satisfy the property that a $(1-o_{1/p}(1))$-fraction of size-$s$ subgraphs are good and that a non-good subgraph is chosen at random.
 ∎
@@
   2. 2.
-    When $p=\Omega(\log(n)/n)^{1/(4k-1)}$, the problem is additionally solvable in $O(p^{-c_{3}k})$ time.
+    When $p$ is additionally $\Omega(\log(n)/n)^{1/(4k-1)}$, the problem is solvable in $O(p^{-c_{3}k})$ time.
   3. 3.
@@
 Claim 3.4.
-Suppose that a graph $G$ satisfies the property that for all $\ell\in\{2,3,\dots,k\}$ the number of labeled $\ell$-cliques in $G$ is in the range $(1\pm\frac{\varepsilon}{4(k+1)})\binom{n}{\ell}\ell!\cdot p^{\binom{\ell}{2}}$ (for $p=\Omega\left(\varepsilon^{-2}n^{-2/(k-1)}\right)$).
+Suppose that a graph $G$ satisfies the property that for all $\ell\in\{2,3,\dots,k\}$ the number of labeled $\ell$-cliques in $G$ is in the range $(1\pm\frac{\varepsilon}{4(k+1)})\binom{n}{\ell}\ell!\cdot p^{\binom{\ell}{2}}$.
 Then, for size $s_{\ell}$ multisets, for all $\ell\in[k-1]$, $G$ is $\left((\ell+1)\cdot\frac{\varepsilon}{4(k+1)},(\ell+1)^{2}\cdot\frac{\varepsilon^{2}}{(k+1)^{2}}\cdot\frac{p^{2\ell-2}}{2048},s_{\ell}\right)$-exponentially robustly quasirandom with respect to $\ell$-clique counts.
@@
 Completeness.
-By Section 2.3, with high constant probability, for $G\sim\mathcal{G}_{n,p}$ with $p=\Omega\left(\varepsilon^{-2}n^{-2/(k-1)}\right)$, for all $\ell\leq k$ the number of labeled $\ell$-cliques in $G$ is in the range $(1\pm\frac{\varepsilon}{4(k+1)})\binom{n}{\ell}\ell!\cdot p^{\binom{\ell}{2}}$.
+By Section 2.3, with high probability, for $G\sim\mathcal{G}_{n,p}$ with $p=\omega\left(\varepsilon^{-2}n^{-2/(k-1)}\right)$, for all $\ell\leq k$ the number of labeled $\ell$-cliques in $G$ is in the range $(1\pm\frac{\varepsilon}{4(k+1)})\binom{n}{\ell}\ell!\cdot p^{\binom{\ell}{2}}$.
 In the following, suppose that the high-probability event that the clique counts are all in these typical ranges holds.
@@
 $\delta_{\ell}=(\ell+1)\cdot\frac{\varepsilon}{4(k+1)},~~\nu_{\ell}=(\ell+1)^{2}\cdot\frac{\varepsilon^{2}}{(k+1)^{2}}\cdot\frac{p^{2\ell-2}}{2048}.$
-By Section 3.1.2, from this exponential quasirandomness, we can conclude that, with probability at least $1-k\cdot 4^{k}\exp\left(-\nu_{k}s_{*}\right)\geq 9/10$, for all $\ell\in[k]$, the uniformly sampled multiset $S$ satisfies $C_{\ell}(S)\in(1\pm\frac{\varepsilon}{4})E_{s,\ell}$.
-Therefore, the algorithm accepts $G\sim\mathcal{G}_{n,p}$ with high constant probability.
+By Section 3.1.2, from this exponential quasirandomness, we can conclude that, with probability at least $1-k\cdot 4^{k}\exp\left(-\nu_{k}s_{*}\right)=1-o(1)$, for all $\ell\in[k]$, the uniformly sampled multiset $S$ satisfies $C_{\ell}(S)\in(1\pm\frac{\varepsilon}{4})E_{s,\ell}$.
+Therefore, the algorithm accepts $G\sim\mathcal{G}_{n,p}$ with high probability.
 Soundness.
@@
 Consider $G\sim\mathcal{G}_{n,p}$, and choose a random subgraph of $G$ on $s_{*}$ vertices.
-By Corollary 2.3, for $p\geq\left(\frac{1200C^{4}\log n}{n}\right)^{1/(4k-1)}$ and $s_{*}\geq 1200C^{4}/p^{5k}$, with probability at least $4/5$, the subgraph on these $s_{*}$ vertices satisfies the conditions checked for in Step 3 with high probability, as proven in Section 2.3.
-Therefore, Step 3 does not reject $G\sim\mathcal{G}_{n,p}$ with high constant probability.
+By Corollary 2.3, for $p\geq\left(\frac{1200C^{4}\log n}{n}\right)^{1/(4k-1)}$ and $s_{*}\geq 1200C^{4}/p^{5k}$, with probability 1o(1)1-o(1), the subgraph on these $s_{*}$ vertices satisfies the conditions checked for in Step 3 with high probability, as proven in Section 2.3.
+Therefore, Step 3 does not reject $G\sim\mathcal{G}_{n,p}$ with high probability.
 Next, we argue that, with probability 1o(1)1-o(1) the approximation of the number of $\ell$-cliques, for each $\ell\in[k]$, returned by the algorithm is a $(1\pm\frac{1}{8}\zeta_{\ell-1})$-approximation of $C_{\ell}(S)$.
@@
 Observation 4.3.
-Define $s$ as in Section 5. Then $s=\widetilde{O}\left(r_{k}(H)^{-1}\varepsilon^{-2}\right)=\widetilde{O}\left(p^{-2\Delta(H)}\varepsilon^{-2}\right).$
+Define $s_{*}$ as in Section 5. Then $s_{*}=\widetilde{O}\left(r_{k}(H)^{-1}\varepsilon^{-2}\right)=\widetilde{O}\left(p^{-2\Delta(H)}\varepsilon^{-2}\right).$
 We are now ready to prove the upper bound in Theorem 4.1 as an implication of Theorem 5.1.
@@
 Consider a motif $H$ on $k$ vertices.
-By Section 2.3, for $p=\Omega\left(n^{-1/m_{H}}\right)$ $\mathcal{G}_{n,p}$ satisfies the properties that all $H^{\prime}\subseteq H$ the number of labeled induced copies of $H^{\prime}$ is in the range $(1\pm\frac{\varepsilon}{100})\binom{n}{i}C_{H^{\prime}}$ for some $C_{H^{\prime}}$ with high constant probability, and that $r_{\ell-1}(H)\geq r_{\ell}(H)$ for all $\ell\in\{2,3,\dots,k\}$.
+By Section 2.3, for $p=\omega\left(n^{-1/m_{H}}\right)$ $\mathcal{G}_{n,p}$ satisfies the properties that all $H^{\prime}\subseteq H$ the number of labeled induced copies of $H^{\prime}$ is in the range $(1\pm\frac{\varepsilon}{100})\binom{n}{i}C_{H^{\prime}}$ for some $C_{H^{\prime}}$ with high probability, and that $r_{\ell-1}(H)\geq r_{\ell}(H)$ for all $\ell\in\{2,3,\dots,k\}$.
 Therefore, Theorem 5.1 implies that quality control for the count of $H$ in an input graph $G$ compared to $\mathcal{G}_{n,p}$ can be accomplished in $O\left(r_{k}(H)^{-2}\varepsilon^{-4}\right)=p^{-O(\Delta(H))}\cdot\mathrm{poly}(1/\varepsilon)$ queries.
@@
 For constant error parameter $\varepsilon$ and
-$p=\Omega\left(\max\left\{\left(\log(n)/n\right)^{1/(4\sigma(H)-1)},n^{-1/m_{H}^{\star}}\right\}\right)$, and induced motif $H$ on $k$ vertices, there is an algorithm for $(\mathcal{G}_{n,p},\rho_{H})$-quality control that makes $\frac{1}{p^{O(\Delta(H))}}$ adjacency matrix queries and has a runtime of $\frac{1}{p^{O(\Delta(H))}}$.
+$p$ in both $\Omega\left(\left(\log(n)/n\right)^{1/(4\sigma(H)-1)}\right)$ and $\omega\left(n^{-1/m_{H}^{\star}}\right)$, and induced motif $H$ on $k$ vertices, there is an algorithm for $(\mathcal{G}_{n,p},\rho_{H})$-quality control that makes $\frac{1}{p^{O(\Delta(H))}}$ adjacency matrix queries and has a runtime of $\frac{1}{p^{O(\Delta(H))}}$.
 While the above is stated for constant $\varepsilon$, the bounds with the explicit dependence on $\varepsilon$ are: $p^{-O(k)}\cdot\mathrm{poly}(1/\varepsilon)$ for the query complexity, and $p^{-O(k)}\cdot\exp(\mathrm{poly}(|H|/\varepsilon))$ for the runtime.
@@
 $C_{H}(G)=\sum_{i=0}^{\binom{k}{2}-e(H)}(-1)^{i}\cdot\sum_{\begin{subarray}{c}H^{\prime}\supseteq H\\ |H^{\prime}|=k\\ e(H^{\prime})=e(H)+i\end{subarray}}\widetilde{C}_{H^{\prime}}(G).$
-We also observe that Section 2.3 can imply that, for $p=\Omega\left(\max_{H^{\prime}\supseteq H,|H^{\prime}|=k}n^{-1/m_{H^{\prime}}}\cdot\varepsilon^{-2}\right)$, with probability at least $9/10$, for all $H^{\prime}\supseteq H,|H^{\prime}|=k$, the number of copies of $H^{\prime}$ in $G$ is in the range $(1\pm\frac{\varepsilon}{100k^{k}}\binom{n}{k}p^{e(H^{\prime})}(1-p)^{\binom{k}{2}-e(H)}$.
+We also observe that Section 2.3 can imply that, for $p=\omega\left(\max_{H^{\prime}\supseteq H,|H^{\prime}|=k}n^{-1/m_{H^{\prime}}}\cdot\varepsilon^{-2}\right)$, with probability at least 1o(1)1-o(1), for all $H^{\prime}\supseteq H,|H^{\prime}|=k$, the number of copies of $H^{\prime}$ in $G\sim\mathcal{G}_{n,p}$ is in the range $(1\pm\frac{\varepsilon}{100k^{k}})\binom{n}{k}p^{e(H^{\prime})}(1-p)^{\binom{k}{2}-e(H)}$.
 This is because, as noted in Section 2 after Section 2.3, the choice of $\varepsilon/100$ was not specific and can be changed with the hidden constants in the lower bound on $p$.
-The same applies for the error probability.
 We now prove Theorem 4.7 for induced motifs.
@@
 Completeness: Suppose $G\sim\mathcal{G}_{n,p}$.
-Then, by the observation made above, with probability at least $9/10$, for all $H^{\prime}\supseteq H,|H^{\prime}|=k$, the number of copies of $H^{\prime}$ in $G$ is in the range $(1\pm\frac{\varepsilon}{100k^{k}}\binom{n}{k}p^{e(H^{\prime})}(1-p)^{\binom{k}{2}-e(H)}$.
-Additionally, with similar (up to constants) parameter choices for the algorithm, the correctness of Algorithm NonInduced-Motif-Quality-Efficient on an input $H$ can be boosted to $1-1/(10\cdot 2^{\binom{k}{2}})$.
-Therefore, by the correctness of NonInduced-Motif-Quality-Efficient, with probability at least $9/10$, all calls to Algorithm NonInduced-Motif-Quality-Efficient in Step 1 of Algorithm Induced-Motif-Quality-Efficient accept.
+Then, by the observation made above, with probability 1o(1)1-o(1), for all $H^{\prime}\supseteq H,|H^{\prime}|=k$, the number of copies of $H^{\prime}$ in $G$ is in the range $(1\pm\frac{\varepsilon}{100k^{k}})\binom{n}{k}p^{e(H^{\prime})}(1-p)^{\binom{k}{2}-e(H)}$.
+Therefore, by the correctness of NonInduced-Motif-Quality-Efficient, with probability 1o(1)1-o(1), all calls to Algorithm NonInduced-Motif-Quality-Efficient in Step 1 of Algorithm Induced-Motif-Quality-Efficient accept.
 Therefore, Algorithm Induced-Motif-Quality-Efficient accepts $G\sim\mathcal{G}_{n,p}$ with high probability.
@@
 Therefore, there exists an $H^{\prime}\supset H,|H^{\prime}|=k$ such that $\left|\widetilde{C}_{H^{\prime}}(G)-\widetilde{\mu}_{H^{\prime}}\right|\geq\varepsilon\mu_{H}/2^{\binom{k}{2}}$.
-We refer to this $H^{\prime}$ as $H^{\circ}$
+We refer to this $H^{\prime}$ as $H^{\circ}$.
 Next observe that, for $H^{\prime}\supseteq H,|H^{\prime}|=k$, $\mu_{H^{\prime}}\leq\mu_{H}$. Applying this and $\widetilde{\mu}_{H^{\prime}}\leq\mu_{H^{\prime}}\cdot 2^{\binom{k}{2}}$, we find that:
@@
 In this section, we generalize our results to the setting of quality control of any motif $H$ as compared to any distribution $\mathcal{D}_{n}$ over graphs with a good concentration of motif counts.
-Particularly, we will consider distributions $\mathcal{D}_{n}$ over graphs such that, for $G\sim\mathcal{D}_{n}$, for all $P\subseteq H$ the number of labeled induced copies of $P$ is in the range $(1\pm\frac{\varepsilon}{100})E_{\mathcal{D}_{n}}(P)$ for some $E_{\mathcal{D}_{n}}(P)$ with high constant probability.
+Particularly, we will consider distributions $\mathcal{D}_{n}$ over graphs such that, for $G\sim\mathcal{D}_{n}$, for all $P\subseteq H$ the number of labeled induced copies of $P$ is in the range $(1\pm\frac{\varepsilon}{100})E_{\mathcal{D}_{n}}(P)$ for some $E_{\mathcal{D}_{n}}(P)$ with high probability.
 Before stating the main theorem we prove in this section (Theorem 5.1), we set up some notation and terminology.
@@
   1. 1.
-    For $G\sim\mathcal{D}_{n}$, for all $\ell\in[k]$ and all $P\subseteq H$ on $\ell$ vertices, the number of labeled induced copies of $P$ is in the range $(1\pm\frac{\varepsilon}{100})E_{\mathcal{D}_{n}}(P)$ with high constant probability.
+    For $G\sim\mathcal{D}_{n}$, for all $\ell\in[k]$ and all $P\subseteq H$ on $\ell$ vertices, the number of labeled induced copies of $P$ is in the range $(1\pm\frac{\varepsilon}{100})E_{\mathcal{D}_{n}}(P)$ with probability 1o(1)1-o(1).
   2. 2.
@@
 $\delta_{\ell}=(\ell+1)\cdot\frac{\varepsilon}{4(k+1)},~~\nu_{\ell}=(\ell+1)^{2}\cdot\frac{\varepsilon^{2}}{16(k+1)^{2}}\cdot\frac{r_{\ell}(H)}{2048\cdot\ell^{2}}.$
-From this, Section 5.1.2 implies that with probability at least $1-k\cdot 4^{k}\exp\left(-\nu_{k}s_{*}\right)\geq 9/10$, the uniformly sampled multiset $S$ satisfies $C_{H^{\prime}}(S)\in(1\pm\frac{\varepsilon}{4})S_{\mathcal{D}_{n}}(H^{\prime},s_{*})$ for all $H^{\prime}\subseteq H$.
-Therefore, the algorithm accepts $G\sim\mathcal{D}_{n}$ with high constant probability.
+From this, Section 5.1.2 implies that with probability at least $1-k\cdot 4^{k}\exp\left(-\nu_{k}s_{*}\right)=1-o(1)$, the uniformly sampled multiset $S$ satisfies $C_{H^{\prime}}(S)\in(1\pm\frac{\varepsilon}{4})S_{\mathcal{D}_{n}}(H^{\prime},s_{*})$ for all $H^{\prime}\subseteq H$.
+Therefore, the algorithm accepts $G\sim\mathcal{D}_{n}$ with high probability.
 Soundness.
@@
 Lemma A.4 ([Fei06], simplified).
-There is an algorithm that uses $O\left(\sqrt{n/d}\right)$ degree queries and time and estimates the average degree of an input graph $G$ up to a multiplicative factor of $3$, with high constant probability, as long as $G$ has average degree at least $d$.
+There is an algorithm that uses $\widetilde{O}\left(\sqrt{n/d}\right)$ degree queries and time and estimates the average degree of an input graph $G$ up to a multiplicative factor of $3$, with probability 1o(1)1-o(1), as long as $G$ has average degree at least $d$.
 We rely on the following corollary, which turns this into a decision version of the problem.
@@
 There exists an algorithm (call it Algorithm B) that performs expected $O(\log(n)/p)$ queries and has expected runtime $O(\log(n)/p)$ with the following guarantees.
-If $e(G)\in\left[\frac{1}{2}n^{2}p,2n^{2}p\right]$, the algorithm accepts with high constant probability.
-If $e(G)<\frac{1}{5}n^{2}p$ or $e(G)>10n^{2}p$, the algorithm rejects with high constant probability.
+If $e(G)\in\left[\frac{1}{2}n^{2}p,2n^{2}p\right]$, the algorithm accepts with high probability.
+If $e(G)<\frac{1}{5}n^{2}p$ or $e(G)>10n^{2}p$, the algorithm rejects with high probability.
 Remark A.4.
@@
 While the focus in this section is on triangles, a modification of this corollary already implies a $(\mathcal{G}_{n,p},\rho_{e})$-quality control algorithm for comparing the number of edges to $\mathcal{G}_{n,p}$, with $\widetilde{O}(\sqrt{1/p})\cdot\text{poly}(1/\varepsilon)$ adjacency list queries and runtime.
-Particularly, [GR08] provides stronger guarantees than [Fei06], giving a $\widehat{O}(n^{2}/e(G))^{1/2}\cdot\text{poly}(1/\varepsilon)$-time algorithm that outputs a $(1\pm\varepsilon)$-approximation of the average degree in a graph $G$.
+Particularly, [GR08] provides stronger guarantees than [Fei06], giving a $\widetilde{O}(n^{2}/e(G))^{1/2}\cdot\text{poly}(1/\varepsilon)$-time algorithm that outputs a $(1\pm\varepsilon)$-approximation of the average degree in a graph $G$.
 We can use this to prove a more refined version of Section A.1.
@@
 The algorithm has expected $O(\sqrt{1/p})$ queries and has expected runtime $O(\sqrt{1/p})$.
-This procedure has the guarantee that if $e(G)\in\left[\frac{1}{2}n^{2}p,2n^{2}p\right]$, the algorithm accepts with high constant probability, and if $e(G)>10n^{2}p$, the algorithm rejects with high probability.
+This procedure has the guarantee that if $e(G)\in\left[\frac{1}{2}n^{2}p,2n^{2}p\right]$, the algorithm accepts with high probability, and if $e(G)>10n^{2}p$, the algorithm rejects with high probability.
 ∎
@@
 There exists an algorithm (call it Algorithm D) with the following guarantee.
-When $\mathrm{arb}(G)\leq\alpha^{\prime}$, if $C_{3}(G)\geq t$, then with high constant probability, the algorithm outputs a number $\widehat{t}$ that is in the range $(1\pm\delta)C_{3}(G)$.
+When $\mathrm{arb}(G)\leq\alpha^{\prime}$, if $C_{3}(G)\geq t$, then with high probability, the algorithm outputs a number $\widehat{t}$ that is in the range $(1\pm\delta)C_{3}(G)$.
 The algorithm performs adjacency list, adjacency matrix, and degree queries. The expected query complexity and runtime of the algorithm are:
@@
 Completeness: We first reason about completeness, i.e., that Algorithm 1 accepts $G\sim\mathcal{G}_{n,p}$ with high probability (both over the distribution over graphs and the algorithm’s randomness).
-By Section 2.3 and Section 2.3, with high constant probability, since $p=\Omega\left(\varepsilon^{-2}\cdot 1/n\right)$, $G\sim\mathcal{G}_{n,p}$ satisfies $\mathrm{arb}(G)\leq 2np$, $m\in(1\pm\frac{\varepsilon}{100})\binom{n}{2}p$, and $C_{3}(G)\in(1\pm\frac{\varepsilon}{100})\binom{n}{3}p^{3}$.
+By Section 2.3 and Section 2.3, with high probability, since $p=\omega\left(\varepsilon^{-2}\cdot 1/n\right)$, $G\sim\mathcal{G}_{n,p}$ satisfies $\mathrm{arb}(G)\leq 2np$, $m\in(1\pm\frac{\varepsilon}{100})\binom{n}{2}p$, and $C_{3}(G)\in(1\pm\frac{\varepsilon}{100})\binom{n}{3}p^{3}$.
 We first argue that, when $G\sim\mathcal{G}_{n,p}$, all algorithms called as subroutines are able to terminate successfully (and therefore we can utilize their high-probability guarantees).
@@
 Remark A.6.
-The algorithm above only used the following properties of $\mathcal{G}_{n,p}$: with high constant probability, the arboricity is bounded, and the number of edges and triangles are each concentrated around their expected values.
+The algorithm above only used the following properties of $\mathcal{G}_{n,p}$: with high probability, the arboricity is bounded, and the number of edges and triangles are each concentrated around their expected values.
 Therefore, the algorithm and analysis extend to any distribution over graphs with these properties.
@@
 Completeness: Suppose the input graph $G$ is drawn from $\mathcal{G}_{n,p}$.
-Then, by the completeness of Algorithm Induced-Motif-Quality-Efficient proven in Section 4.1, each call of Algorithm Induced-Motif-Quality-Efficient in Step 1 accepts $G$ with probability at least $1-\frac{1}{3\cdot M(kd)}$.
-(Observe that Algorithms 2 and 3 were analyzed with a different constant success probability, but since $M(k,d)$ is constant with respect to $n,p$, we can replace it with $1-\frac{1}{3\cdot M(k,d)}$ in the analysis.) By a union bound over the $M(k,d)$ possible induced subgraphs over at most $2k^{d}$ vertices, Step 2 then accepts $G\sim\mathcal{G}_{n,p}$ with probability at least $2/3$.
+Then, by the completeness of Algorithm Induced-Motif-Quality-Efficient proven in Section 4.1, each call of Algorithm Induced-Motif-Quality-Efficient in Step 1 accepts $G$ with probability at least 1o(1)1-o(1).
+By a union bound over the $M(k,d)$ (a constant) possible induced subgraphs over at most $2k^{d}$ vertices, Step 2 then accepts $G\sim\mathcal{G}_{n,p}$ with probability 1o(1)1-o(1).
 Soundness: Suppose the input $G$ satisfies
@@
 Works on average-case algorithms typically define a “good” property of inputs with the following guarantees.
-With probability at least $1-1/t_{2}$, an input drawn from $\mathcal{D}$ possesses the good property, and thus there exists an algorithm $A$ with worst-case runtime $t_{1}$ for this input.
+With probability at least $1-1/t_{2}$, for $t_{2}=\omega(1)$, an input drawn from $\mathcal{D}$ possesses the good property, and thus there exists an algorithm $A$ with worst-case runtime $t_{1}$ for this input.
 This “good” property is possible to certify/verify.
@@
 We now argue about the runtime.
-First, with probability at least 11/T1-1/T, $x\sim\mathcal{D}$ satisfies $\rho(x)\in(1\pm\varepsilon)$, and so with probability at least $(1-1/T)\cdot 2/3\geq 1-1/t_{4}$ (over the input distribution), $x\sim\mathcal{D}$ is accepted by Algorithm $C$.
+First, with probability at least 11/T1-1/T, $x\sim\mathcal{D}$ satisfies $\rho(x)\in(1\pm\varepsilon)$, and so with probability at least $(1-1/T)\cdot(1-o(1))\geq 1-1/t_{4}$ (over the input distribution), $x\sim\mathcal{D}$ is accepted by Algorithm $C$.
 In this case, the algorithm takes time $t_{3}$.
@@
 Proposition D.0 (Section 2.3, restated).
-Let $H$ be an induced motif on $k$ vertices. Consider parameters $n,p$ with $n\geq\frac{10}{\varepsilon^{2}}\cdot p^{-m_{H}}$, where $m_{H}:=\max_{F\subseteq H}e(F)/v(F)$.
-For $G\sim\mathcal{G}_{n,p}$, with probability at least $9/10$, the number of copies of $H$ in $G$ is in the range $(1\pm\frac{\varepsilon}{100})\binom{n}{v(H)}v(H)!\cdot p^{e(H)}(1-p)^{\binom{v(H)}{2}-e(H)}$.
+Let $H$ be an induced motif on $k$ vertices. Consider parameters $n,p$ with $n\geq\frac{\omega_{1/p}(1)}{\varepsilon^{2}}\cdot p^{-m_{H}}$, where $m_{H}:=\max_{F\subseteq H}e(F)/v(F)$.
+For $G\sim\mathcal{G}_{n,p}$, with probability at least $1-o_{1/p}(1)$, the number of copies of $H$ in $G$ is in the range $(1\pm\frac{\varepsilon}{100})\binom{n}{v(H)}v(H)!\cdot p^{e(H)}(1-p)^{\binom{v(H)}{2}-e(H)}$.
 Proof.
@@
 $\mathbb{E}\left[X_{G}\right]+\mathbb{E}\left[X_{G}\right]^{2}\cdot\sum_{i=1}^{k-1}\sum_{\begin{subarray}{c}H^{\prime}\subseteq H\\ |H^{\prime}|=i\end{subarray}}\binom{n}{i}^{-1}\cdot p^{-e(H^{\prime})}.$
-Since $n\geq\frac{10}{\varepsilon^{2}}\cdot p^{-m_{H}}$, this gives us:
-$\text{Var}\left[X_{G}\right]\leq\mathbb{E}\left[X_{G}\right]^{2}\cdot\frac{\varepsilon^{2}}{10}.$
+Since $n\geq\frac{\omega_{1/p}(1)}{\varepsilon^{2}}\cdot p^{-m_{H}}$, this gives us:
+$\text{Var}\left[X_{G}\right]\leq\mathbb{E}\left[X_{G}\right]^{2}\cdot\frac{\varepsilon^{2}}{\omega(1)}=\mathbb{E}\left[X_{G}\right]^{2}\cdot\varepsilon^{2}\cdot o_{1/p}(1).$
 Finally, by Chebyshev’s inequality,
-$\mathbb{P}\left[\left|X_{G}-\mathbb{E}\left[X_{G}\right]\right|\geq\varepsilon\mathbb{E}\left[X_{G}\right]\right]\leq\frac{\text{Var}\left[X_{G}\right]}{\varepsilon^{2}\mathbb{E}\left[X_{G}\right]^{2}}\leq\frac{1}{10}.$
+$\mathbb{P}\left[\left|X_{G}-\mathbb{E}\left[X_{G}\right]\right|\geq\varepsilon\mathbb{E}\left[X_{G}\right]\right]\leq\frac{\text{Var}\left[X_{G}\right]}{\varepsilon^{2}\mathbb{E}\left[X_{G}\right]^{2}}=o_{1/p}(1).$
 D.2 Stochastic Block Models